### v1

# OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows

Abstract
Common methods for studying the kinematics and geometry of outflowing gas rely on modeling emission and absorption lines in integrated spectra using methods that are not physically motivated, including empirical quantiles or fitting multiple Gaussian or Voigt profiles. Such methods are not always consistent with the interpretation of these features and, as a result, miss key underlying physics and can even lead to inaccurate interpretations of observations. To address this problem, we present the publicly available python code OutLines, which provides astrophysical models of spectral emission and absorption line profiles produced by outflows in a variety of environments. The OutLines code accounts for differences in parameterization of the velocity field and density profile while allowing for different outflow geometries, making OutLines versatile and useful for a wide variety of astrophysical phenomena. We demonstrate the wide applicability of OutLines by using the code to model line profiles in an H ii region knot, super star clusters, a starburst galaxy, and an AGN. In each of these contexts, we illustrate how OutLines can illuminate key underlying physics in ways that improve our scientific understanding and address important open questions in astronomy, including the key mechanisms in the baryon cycle, the evolution of H ii regions and galaxies, and even Lyman continuum escape. OutLines will be a critical resource as massively multiplexed spectroscopic surveys like WEAVE-LOFAR and 4MOST/WAVES come online, providing the means to probe feedback kinematics with deeper, higher resolution spectroscopy for unprecedented large samples of galaxies.
keywords:
ISM: jets and outflows – ISM: bubbles – ISM: kinematics and dynamics – galaxies: kinematics and dynamics
pubyear: 2025pagerange: OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows28
1 Introduction
Outflows, winds, and bubbles are ubiquitous astrophysical phenomena occurring on a dynamic range of spatial scales and in a variety of contexts, from young stellar objects (YSOs) to nebulae, active galactic nuclei (AGN), galaxies, and even galaxy clusters. This feedback has the capacity to shape many key astrophysical systems, from YSO winds affecting the formation of exoplanets (e.g., Suzuki2009; Manara2023) to massive stars redistributing the leftover material from their birth clouds (e.g., Dopita2005; Menon2024) to AGN and starbursts driving the baryon cycle of galaxies (e.g., 2005ApJ...618..569M; Tumlinson2017). Many physical mechanisms can accelerate, or drive, winds and bubbles, from cosmic rays (e.g., Ipavich1975) to hot expanding gas (e.g., 1985Natur.317...44C) to radiation pressure on dust (e.g., Poynting1904; LyndenBell1971) and ions (so-called line driving, e.g., Lucy1970), each with direct implications for how feedback shapes the surrounding environment (see, e.g., 2020A&ARv..28....2V; Thompson2024, and references therein). A common method of detecting and studying feedback is through broad, often asymmetric spectroscopic features (e.g., Castor1970; 1975ApJ...195..157C; 1974ApJ...193..651N; 1980A&A....81..172P; 1981MNRAS.195..787P; 1981ApJ...247..403H; Chu1986), perhaps most famously with P Cygni wind lines in stellar atmospheres (e.g., 1986A&A...164...86P; 1987ApJ...314..726L; 2001A&A...369..574V). Physically motivated models for these line profiles is thus of critical importance to understanding how energetic phenomena shape their environment.
While the physics and formalism for spectroscopic features arising from winds, bubbles, and outflows is not new (e.g., Beals1931; 1960mes..book.....S; Castor1970; 1975ApJ...195..157C; 1986A&A...164...86P; 1987ApJ...314..726L), their application is relatively sparse outside the phenomena of stellar atmospheres (e.g., 2001A&A...369..574V; 2021MNRAS.504.2051V; Sander2017; Puls2020). Instead, fitting multiple Gaussian or Voigt profiles to broad, asymmetric lines has become the standard (e.g., 1980A&A....81..172P; 1981MNRAS.195..787P) for a wide variety of objects, including H ii regions (e.g., Bresolin2020; Yarovova2025), AGN (e.g., Liu2020), and galactic winds (e.g., Hogarth2020; Peng2023; Peng2025), despite early evidence that physics-based profiles are not only more appropriate but also demonstrably different in shape while still providing good descriptions of observations (e.g., Zheng1990; TenorioTagle1996). Some studies have skirted the issue by adopting a more empirical approach using velocity quantiles (e.g., 1981ApJ...247..403H; 1985MNRAS.213....1W; 1991ApJS...75..383V) or by fitting the line profile wings with a log-linear (e.g., 2021ApJ...920L..46K; Komarova2025) or exponential (e.g., Rusakov2025) functions to characterize the steepness of the non-Gaussian flux distribution. The benefit of such methods is that they require no physical model while still providing insight into kinematics. However, bridging the gap between these approaches and the underlying physics is not always straightforward.
Some effort towards a physically motivated approach has emerged in recent years, including detailed treatment of radiation pressure and conservation effects (e.g., 2017MNRAS.471.4061K), turbulent mixing layers (e.g., 2009A&A...500..817B), and relativistic accretion disk winds (e.g., DISKWIND1; DISKWIND2; XRADE) along with more generalized models (e.g., 2010ApJ...717..289S; 2011ApJ...734...24P; Martin2013; 2015ApJ...801...43S; 2016MNRAS.463..541C; Luminari2018; Flury2023). Despite these exciting advancements, wide-spread adoption of these methods has stalled. A key to furthering the use of physically-motivated line profiles is versatile, well-documented, and user-friendly software tools which make such models readily accessible to the community to generate spectral line profiles linking observations to outflow physics.
In this paper, we present OutLines, a python-based code for generating model emission, absorption, and P Cygni line profiles arising from astrophysical winds, bubbles, and outflows. The input geometry and wind physics can be readily customized, and functionality for modeling is provided. We outline the formalism and physics for radiative transfer, velocity fields, gas density profiles, and geometry in §2. Then, in §3 we present the implementation of these components in the OutLines code111available at github.com/sflury/OutLines and via pip install.. We demonstrate the versatile applicability and scientific usefulness of OutLines in §4 by modeling broad spectral lines in an H ii region knot, a super star cluster (SSC), a starbursting galaxy, and an active galactic nucleus (AGN), with profiles generated by OutLines. Finally, we summarize OutLines and its potential for high scientific impact in §5, providing worked examples in Appendix D and E.
2 Physical Model
Within the framework of the Sobolev approximation (1944SvA.....21..143S; 1957SvA.....1..678S; 1960mes..book.....S), small scale Doppler motions of gas within clouds can be ignored in the context of the large-scale motion of the clouds themselves. The Sobolev approach has a long-standing precedent in the study of winds (e.g., Castor1970; 1974ApJ...193..651N; 1975ApJ...202..465M; 1987ApJ...314..726L), with successful application to galactic winds and outflows over the past 15 years (e.g., 2010ApJ...717..289S; 2011ApJ...734...24P; Martin2013; 2015ApJ...801...43S; 2016MNRAS.463..541C; Flury2023). Under this assumption, the key characteristics of winds, outflows, and bubbles which dictate the line profile are the velocity field, density profile, and geometry of the clouds (e.g., Zheng1990). Below, we elaborate how the observed projected velocity $u$ in a spectral line, given by
$u=\frac{\lambda-\lambda_{0}}{\lambda_{0}}\frac{c}{v_{\infty}}$ (1)
for a line with rest wavelength $\lambda_{0}$ and normalized to the wind terminal velocity $v_{\infty}$, relates to the intrinsic wind velocity $w$ given by some radially-symmetric velocity field
$w=\frac{v_{\rm wind}[x,\beta]}{v_{\infty}}$ (2)
with respect to some radial distance $x=r/R_{0}$ from the source of the outflow, with radius $R_{0}$, and defined by some exponent $\beta$. In addition to the velocity field, one must also know the radial gas density distribution, given by
$n^{\prime}[r]=n_{0}n[x]$ (3)
for a dimensionless profile $n$ and maximum density $n_{0}$, and the azimuthal geometry of the wind. With the velocity field $w$, density profile $n$, and outflow geometry in hand, radiative processes can translate outflows from the frame of the source to the spectral line in the frame of the observer.
2.1 Radiative Processes
Below, we outline the formalism for absorption and emission line profiles arising from astrophysical winds, with implicit assumptions about the velocity field $w[x]$, the radial gas distribution $n[x]$, and the gas geometry.
2.1.1 Line Absorption
In general, line absorption optical depth $\tau_{\lambda}$ at any particular wavelength $\lambda$ in a feature is given by
$\tau_{\lambda}=\frac{N}{v}f\lambda_{0}\sigma_{a}\phi_{\lambda}$ (4)
[…]
Under the Sobolev approximation, instead one must evaluate the relative column density at each velocity $v$ in the wind or bubble that is projected onto the line of sight velocity $u$. The optical depth profile for the outflow component is given by the Sobolev approximation such that
$\phi_{\lambda}=K^{-1}\int n[x]\frac{{\rm d}x[w]}{{\rm d}w}{\rm d}\Omega$ (5)
over some differential solid angle ${\rm d}\Omega$ given by the geometry (see Rybicki1983, their Equation 2.14) and where $K$ is a normalization constant ensuring $\int_{\lambda}\phi_{\lambda}{\rm d}\lambda=1$ (Rybicki1979, their Equation 1.65). The differential term $\frac{{\rm d}x}{{\rm d}w}$ arises from integrating over all column densities associated with the observed velocity $u$ (Castor1970, their Equation 6, 2016MNRAS.463..541C, their Equation 4, assuming no stimulated emission). The corresponding integral limits are given by constraints from the geometry (§2.4).
The emergent flux, including a static component with a Voigt profile, is therefore given by
$\frac{F_{\lambda}}{F_{\lambda 0}}=\exp\left[\tau_{\rm s}\phi_{\lambda,\rm Voigt}+\tau_{\rm o}\phi_{\lambda,\rm out}\right]$ (6)
for intrinsic (backlight) flux density $F_{\lambda 0}$ and emergent flux density $F_{\lambda}$ and where
$\tau_{\rm s}=f\lambda_{0}\sigma_{a}b^{-1}N_{\rm s}$ (7)
for static gas column density $N_{s}$ and
$\tau_{\rm o}=f\lambda_{0}\sigma_{a}v_{\infty}^{-1}N_{\rm o}$ (8)
for outflowing gas column density $N_{\rm o}$.
2.1.2 Nebular Line Emission
For nebular emission lines, the emergent line flux density $F_{\lambda}$ is simply the profile $\phi_{\lambda}$ scaled by the integrated line flux $F$ such that
$F_{\lambda}=F\phi_{\lambda}$ (9)
for any given wavelength. “Static” gas undergoes Doppler broadening from thermal, orbital, and turbulent motion which produces the traditional Gaussian form of $\phi_{\lambda}$. In the case of an outflow, we must first consider how the line emission depends on density and velocity. The line emission coefficient $j_{\lambda}$ is typically defined as proportional to the density of the populated upper state producing the transition (Rybicki1979; source:osterbrock2006). Since nebular lines emerge from upper electronic states populated by recombination or collisional excitation, $j\propto n[x(w)]^{2}$ (e.g., source:osterbrock2006; Peimbert_2017). Thus, we can simply express the line profile as
$\phi_{\lambda}=K^{-1}\int n\left[x[w]\right]^{2}{\rm d}\Omega$ (10)
where $K$ again is a normalization constant ensuring $\int_{\lambda}\phi_{\lambda}{\rm d}\lambda=1$ (see Flury2023). As with absorption, limits on the integral are given by the geometry in §2.4.
The emergent flux, including a static component with a Gaussian profile, is then given by
$F_{\lambda}=F_{\rm s}\phi_{\lambda,\rm Gauss}+F_{\rm o}\phi_{\lambda,\rm out}$ (11)
for nebular emission.
2.1.3 Resonant and Fluorescent Line Emission
Unlike nebular line emission, which depends on collisional and recombination processes, the upper state population of resonant and fluorescent lines depends primarily on the absorption of photons which are subsequently re-emitted in a random direction. Therefore, the line emission coefficient $j_{\lambda}$ is proportional to nndx/x/dww, meaning that the line profile is determined using Equation 5 but with integral limits for line emission given by integrating over the entire volume (see §2.4). However, photons emitted in this fashion can be re-absorbed and subsequently re-emitted in an arbitrary direction, effectively a scattering process, which will alter the line profile. Resonant and fluorescent lines re-emit light which has previously been absorbed isotropically, meaning that the line flux is determined by the source function $S_{\lambda}=1-\exp[-\tau_{\lambda}]$ (see, e.g., Castor1970, their §6, Rybicki1983, their §3) such that
$\frac{F_{\lambda}}{F_{\lambda 0}}=\epsilon f_{\rm esc,\lambda}\left(1-\exp\left[-\tau_{\lambda}\right]\right)$ (12)
where $\tau_{\lambda}$ is given by line absorption in Equations 4 and 5 with integral limits consistent with line emission in Equation 10, $\epsilon$ is the filling factor to account for clumpiness because individual clouds do not sample every possible sightline even when they are isotropically distributed, and $f_{\rm esc,\lambda}$ is the fraction of photons which escape radiative trapping at a given wavelength $\lambda$. The escape fraction of photons depends on the fraction $p_{i}=A_{i}/\sum A$ of photons which are produced by a transition $i$ given by the transition rate $A_{i}$ divided by the total radiative depopulation rate $\sum A$, where each $i$ corresponds to a resonant ($p_{res}$) or fluorescent ($p_{fluor}$) channel. Following formalism from Castor1970, the probability that a resonant photon can escape from radiative “trapping” through resonant scattering is given by the ratio of emission to absorption $(1-\exp[-\tau_{\rm res,\lambda}])/\tau_{\rm res,\lambda}$. Trapped photons either scatter through emission and re-absorption in the resonant channel or re-emit in the fluorescent channel. After multiple scatterings, the resonant photons are increasingly trapped relative to the fluorescent photons, meaning that the escape fraction is
$f_{\rm esc,\lambda}^{\rm res}=p_{\rm res}\frac{1-\exp[-\tau_{\rm res,\lambda}]}{\tau_{\rm res,\lambda}}\left(1-p_{\rm res}\frac{1-\exp[-\tau_{\rm res,\lambda}]}{\tau_{\rm res,\lambda}}\right)^{-1}$ (13)
for resonant photons and
$f_{\rm esc,\lambda}^{\rm fluor}=p_{\rm fluor}\left(1-p_{\rm res}\frac{1-\exp[-\tau_{\rm res,\lambda}]}{\tau_{\rm res,\lambda}}\right)^{-1}$ (14)
for fluorescent photons under the infinite scattering limit given by 2015ApJ...801...43S.
2.2 Solutions to Wind Equations of Motion
Velocity profiles for winds can take on a variety of shapes arising from different treatments of the wind equations of motion (e.g., CAK theory 1975ApJ...195..157C; 1976ApJS...32..715L; 1977PhDT.........3A; 1977ApJ...213..737B; Castor1979). The simple $\beta$ law approximation
$w[x]=\left(1-x^{-1}\right)^{\beta}{\rm,\hskip 30.00005ptBetaCAK}$ (15)
first appears in early treatment of Wolf-Rayet P Cygni lines for $\beta=0.5$ (Equation 23 Castor1970) and arises from a generalization of CAK theory solutions (see, e.g., Equation 47 of 1975ApJ...195..157C) without loss of physical insight (see, e.g., 1977ApJ...213..737B; Castor1979; 1986A&A...164...86P).The CAK $\beta$ law approximation is ubiquitous in the study of stellar winds (e.g., 2001A&A...369..574V; 2021MNRAS.504.2051V; Telford2024) and has even seen application to galactic outflows (e.g., 2016MNRAS.463..541C; Flury2023). Given its widespread adoption and underlying physical motivation, the CAK $\beta$ law is our default and fiducial velocity field.
In addition to the $\beta$ law approximation to CAK theory, OutLines includes two alternate solutions to the wind equations proposed in the literature. The first, from 2010ApJ...717..289S, assumes a power law approximation to the radial acceleration field to obtain
$w[x]=\left(1-x^{1-\beta}\right)^{1/2}{\rm,\hskip 30.00005ptAccPlaw}$ (16)
where the normalization terms are implicit in $w$ and $x$ and $r_{min}=1$. While those authors do not provide a corresponding physical model, their acceleration power law has a velocity field qualitatively similar to those of Ipavich1975, which assumed cosmic ray driving, and 1985Natur.317...44C, which assumed constant momentum and energy deposition into a steady-state wind from a fluid in which it is entrained (see also, e.g., Owocki2025), and formally akin to radiative acceleration of expanding clouds (see Equations 16-17 Mathews1974). Note that the acceleration power law field with $\beta=2$ is identical to that of the CAK $\beta$ law when $\beta=0.5$.
The second velocity field alternative, from various sources (e.g., FaucherGiguere2012), constitutes a power law approximation to the radial velocity field such that
$w[x]=A(x-1)^{\beta}{\rm,\hskip 30.00005ptVelPlaw}$ (17)
where the wind is assumed to accelerate from rest at $R_{0}$ and $A$ scales the field. We set $A=0.5$ to allow clouds to reach 50% terminal velocity in roughly the same range of radii as the CAK theory $\beta$ law and generally overlap with the same region in $w$-$x$ space as both the CAK $\beta$ law and acceleration power law fields. Single-shell models of stellar wind bubbles can give rise to power law velocity fields from the shared time dependence of $r\propto t^{3/5}$ and $v\propto t^{-2/5}$ ($v\propto r^{-2/3}$, e.g., Weaver1977; Dyson1977). Building on this framework, FaucherGiguere2012 arrive at a velocity power law solution assuming continuous production of “shells” by an AGN to produce a wind.
Figure 1 depicts the different velocity fields for a range of $\beta$. From this comparison, as suggested by the formalism, the power law velocity field approaches $v_{\infty}$ more rapidly than the power law acceleration field or the $\beta$-law approximation to CAK theory, reaching $v_{\infty}$ at distances far closer to the launch radius. Moreover, the power law velocity field does not approach $v_{\infty}$ asymptotically as do the other velocity fields, meaning that the stopping condition at $v_{\infty}$ is imposed solely by integral limits rather than the formalism for the velocity field. We note that physical wind models almost always exhibit an acceleration which gradually decreases as the outflow asymptotically approaches some terminal velocity (e.g., 1975ApJ...195..157C; 1977ApJ...213..737B; 1985Natur.317...44C; Thompson2015; 2017MNRAS.471.4061K; 2018MNRAS.476..512I; Quataert2022a), indicating that the underlying wind physics strongly favors a CAK $\beta$ law or acceleration power law field over a velocity power law field. While individual wind bubble models indicate a velocity power law field (e.g., Weaver1977; Dyson1977), in ensembles of shells, the distribution of shell ages will dominate the net velocity field (e.g., TenorioTagle1996) and is likely determined by the physical driving mechanism (e.g., Chu1994), suggesting that the CAK theory or acceleration power law velocity fields provide a more physically appropriate description for wind scenarios. That being said, the velocity power law field prescription can be consistent with turbulent streaming including the classic Kolmogorov cascade, with the case of $\beta=1$ indicative of Couette flow (e.g., 2009A&A...500..817B), although such scenarios may not capture the underlying driving mechanism.
Figure 1: Radial velocity fields included in OutLines and given by Equations 15-17. Acceleration from rest at the base $R_{0}$ of the wind ($x=1$) proceeds until the wind reachs the terminal velocity $v_{\infty}$ ($w[x]=1$). Velocity fields include the 1986A&A...164...86P $\beta$ law approximation to CAK theory (1975ApJ...195..157C) (blue, Equation 15), the 2010ApJ...717..289S power law acceleration (green, Equation 16), and the power law velocity field (purple, Equation 17). Dash length increases with increasing $\beta$.
We illustrate the effect of varying the velocity field for fixed $\beta$ and the default OutLines geometry and density profiles in Figure 2. As suggested by the velocity fields in Figure 1, the acceleration power law causes $w\to 1$ more rapidly than the other two scenarios, resulting in more emergent flux at higher projected velocities and thus the most pronounced intermediate wings. The $\beta$ CAK law field more rapidly populates moderate velocities than the velocity power law field, resulting in a slightly brighter core. However, the acceleration of the velocity power law field overtakes that of the $\beta$ CAK at high velocities, producing more prominent broad wings in the former than the latter.
Figure 2: Nebular emission (top) and resonant absorption (bottom) line profiles for each of the velocity fields included in OutLines assuming $\beta=1.5$ (CAK theory), $\beta=1$ (velocity power law), and $\beta=2$ (acceleration power law) for the OutLines defaults.
2.3 Radial Density Profiles
The gas / clouds within a wind likely vary with projected radial distance from the launching source due to a variety of factors, including conservation of cloud surface area (e.g., 2017MNRAS.471.4061K), fountains (e.g., 2015ApJ...814...83L), or an isothermal mass-conserving gas (e.g., 2005ApJS..160..115R; 2011ApJ...734...24P).
For simplicity, gas number density $n$ can also be assumed to follow a power law such that
$n[x]=x^{-\alpha}{\rm,\hskip 30.00005ptPowerLaw}$ (18)
a reasonable approximation for stellar and galactic winds (e.g., Castor1979; 1985Natur.317...44C, respectively) and a common choice in the literature (e.g., FaucherGiguere2012; 2015ApJ...801...43S; 2016MNRAS.463..541C; Flury2023; Luminari2024A&A) arising from power law solutions to differential equations (as also the case for the velocity profile) and can be indicative of momentum or energy conservation ($\beta=(\alpha-2)/2$ and $\beta=(\alpha-2)/3$, respectively FaucherGiguere2012, for the power law velocity field $\beta$). As an alternative, we can adopt an exponential profile where
$n[x]=\exp\left[-\gamma(x-1)\right]{\rm,\hskip 30.00005ptExponential}$ (19)
which is also readily invoked from solutions to differential equations and may correspond to the relaxation or dissipation of wind material as it projects out from the launch point. For cases where a continuous flow may flatten or precipitously decline beyond some radius, we also implement a broken power law
$n[x]=\begin{cases}x^{-a_{1}},&x<x_{1}\\ x_{1}^{a_{2}-a_{1}}x^{-a_{2}},&x>=x_{1}\end{cases}{\rm,\hskip 30.00005ptPowerLaw2}$ (20)
with a break-point or “knee” at radius $x_{1}$, inner slope $a_{1}$ and outer slope $a_{2}$. Examples of a broken power law include the shock-driven outflow launched by the interaction of winds from massive stars in proximity to Sag A ($a_{1}<a_{2}$, drop off in density profile, e.g., Quataert2004) and the relativistic outflow triggered by the waking/activating supermassive black hole in Swift J164449.3+57345 ($a_{1}>a_{2}$, flattening in density profile, e.g., Berger2012).
For a variety of physical scenarios, including swept up shells and blast waves, we also allow in Outlines the possibility of pulse-like radial density profiles consistent with a shell or bubble with a scaled radius $x_{1}=R_{\rm shell}/R_{0}$ and characteristic width $\sigma_{x}=\sigma_{r}/R_{0}$, including a shell with gas in a normal distribution
$n[x]=\left(2\pi\sigma_{x}^{2}\right)^{-1/2}\exp\left[-\frac{(x-x_{1})^{2}}{2\sigma_{x}^{2}}\right]{\rm,\hskip 10.00002ptNormal,}$ (21)
with a log normal distribution
$n[x]=\left(2\pi(\log\sigma_{x})^{2}\right)^{-1/2}\exp\left[-\frac{(\log x-\log x_{1})^{2}}{2(\log\sigma_{x})^{2}}\right]{\rm,\hskip 10.00002ptLogNormal,}$ (22)
a fast-rise exponential decay (or FRED, Equation 23, see Norris2005) distribution
$\begin{split}n[x]=&\exp[2\mu]\exp\left[-\frac{\tau_{1}}{x+\mu-x_{1}}-\frac{x+\mu-x_{1}}{\tau_{2}}\right],\\ &\mu=\sqrt{\tau_{1}\tau_{2}}\end{split}{\rm\hskip 10.00002ptFRED,}$ (23)
and a uniform distribution
$n[x]=\begin{cases}\frac{1}{2\sigma_{x}},&|x-x_{1}|<\sigma_{x}\\ 0,\text{otherwise}\end{cases}{\rm,\hskip 10.00002ptUniformShell}.$ (24)
These pulse-like density profiles are consistent with simple single expanding shells or bubbles which are symmetric (normal or uniform), leading (FRED), or trailing (log normal) material with respect to the outward direction of the gas. We illustrate the continuous and pulse-like density profiles available in OutLines in Figure 3.
We show the effects of changing the density distribution for a fixed density ($n_{0}$), total column density ($N$), and velocity field ($\beta$, $v_{\infty}$) in Figure 4 assuming a spherical geometry. The emission line profile wings for the shell or pulse-like density distributions are distinct from the continuous distributions; however, distinctions among the various subsets are less clear, likely owing to the geometric dilution of the density profile by integrating over the entire volume of the line-emitting gas. The effect of pulse profiles on the emission line profile is largely consistent with early fixed-velocity single expanding shell models for Wolf-Rayet features and stellar wind bubbles wherein no line core emerges (e.g., Beals1931; TenorioTagle1996). Such marked differences in radial gas distribution may explain the line profiles of ionized gas around pre- and post-SNe stellar populations as in the different clusters in the Sunburst Arc (e.g., Mainali2022).
The absorption line profiles are more distinct for each individual density profile, likely owing to the direct tracing of the line-of-sight gas distribution without any geometric dilution. Deviations from spherical geometry may affect this result, particularly in the case of a narrow bicone normal to the plane of the sky and aligned with the observer, which directly traces the density distribution in emission and absorption.
Figure 3: Radial density profiles included in OutLines and given by Equations 18-23. Pulse-like density profiles, including the normal (‘Normal’, green, Eqn 21), log normal (‘LogNormal’, yellow, Eqn 22), and fast-rise exponential decay (‘FRED’, red, Eqn 23) distributions, represent physical scenarios such as blast waves, expanding shells, and single burst events. For visualization, the pulse-like profiles are arbitrarily set to maximize density at $r=3R_{0}$. Continuous-like density profiles, including power law (‘PowerLaw’, purple, Eqn 18), exponential (‘Exponential’, orange, Eqn 19), and double power law (‘DoublePowerLaw’, blue, Eqn 20) distributions, represent an integral over many pulses and will always maximize density at the base of the outflow where $r=R_{0}$.
Figure 4: Nebular emission (top) and absorption (bottom) line profiles predicted by OutLines for fixed density, column density, and velocity field for different density distributions assuming a spherical geometry.
[…]
$n[x]=\left(2\pi\sigma_{x}^{2}\right)^{-1/2}\sum_{i=0}^{N_{p}}\exp\left[-\frac{(x-x_{1}-ix_{k})^{2}}{2\sigma_{x}^{2}}\right]{\rm,\hskip 10.00002ptPulses}$ (25)
where $\sigma_{x}$ is the width of each pulse, $x_{1}$ is the radius of the smallest pulse (akin to a phase shift) and $x_{k}$ is the interval spacing between pulses. We assume $N_{p}$ such that pulses span the observationally relevant velocity field from $x=1$ to $x[0.95]$ for any given $x_{1}$ and $x_{k}$. To allow pulses to relax or dissipate over time, we also include a damped pulse series
$n[x]=\frac{\exp\left[-\gamma x\right]}{\sqrt{2\pi\sigma_{x}^{2}}}\sum_{i=0}^{N_{p}}\exp\left[-\frac{(x-x_{1}-ix_{k})^{2}}{2\sigma_{x}^{2}}\right]{\rm,\hskip 10.00002ptDampedPulses}$ (26)
which is likely a more physical representation of outflows in formation. We show the effects of a series of pulse-like events or damped pulses on the emergent line profiles in Figure 5 for shells resolved and unresolved in velocity. One might invoke the central limit theorem to argue that a system of many shells or bubbles will result in a Gaussian line profile; however, we find that neither the damped nor undamped pulses behave in this manner. Indeed, TenorioTagle1996 find that only a Gaussian velocity field will produce an intrinsically Gaussian line profile for a sum of many shells, a scenario which does not appear to be consistent with the bulk motion of gas predicted by theory. They further demonstrate that profiles might only appear Gaussian due to smearing by a line spread function. These results are consistent with comparisons of integrated and spatially resolved observations of the ensemble of expanding shells in 30 Doradus where the bulk velocity field produces a line profile with a Gaussian-like core but distinctly non-Gaussian wings (Chu1994).
Figure 5: Nebular emission (top) and absorption (bottom) line profiles predicted by OutLines for a series of expanding bubbles, shells, or other pulse-like episodes in cases of unresolved (orange) or resolved (blue) shells with (dashed) and without (solid) damping to mimic the effects of dissipation. In all cases, density, column density, velocity field, and terminal velocity are fixed, and a spherical geometry is assumed.
In all cases, simple algebra gives the density in the wind as a function of velocity. As the density term $n_{0}$ simply scales the total line profiles in our formalism below, the $n_{0}$ is absorbed into column density $N_{o}$ and flux $F_{0}$ terms, respectively.
2.4 Geometry
Following Flury2023, line emission or absorption is distributed evenly across the polar coordinate for each deprojected velocity (see also, e.g., Beals1931; Zheng1990; TenorioTagle1996; 2015ApJ...801...43S; Xrism2025_PDS456) allowing us to rewrite the differential solid angle as
${\rm d}\Omega={2\pi}{\rm d}w$ (27)
[…]
To illustrate the effects of assumed velocity field and related cosine projection onto the line of sight, we illustrate contours of constant observed velocity (isocontours) in Figure 6 (see also Castor1970, their Figure 1, 1987ApJ...314..726L, their Figure 2). Sometimes referred to as velocity surfaces (as in Castor1970), these isocontours represent the velocities which are cosine projected to a single observed velocity, assuming a particular velocity field and varying the power law index.
Figure 6: Velocity isocontours for the different velocity fields available in OutLines: the $\beta$ law approximation to CAK theory (top), an acceleration power law (middle), and a velocity power law (bottom). Colors indicate fixed observed velocities of $v_{obs}/v_{\infty}=u=0.1$ (yellow/green), 0.3 (orange/blue), and 0.5 (red/purple). Line styles indicate low (dotted), intermediate (solid), and high (dash-dotted) values of $\beta$ for each velocity field. Arrow indicates the direction towards the observer, with cool colors representing blue Doppler shifts and warm colors representing red Doppler shift.
2.4.1 Velocity Limits
The equal contributions of each deprojected velocity to the absorption or emission allows us to reduce the geometry to the velocity limits $w_{\ell}$ and $w_{u}$ of the integrals in Equations 5 and 10. Boundary conditions for velocity in the case of a uniform spherical outflow (illustrated in Figure 7) are given by some minimum and maximum projected velocities $w_{\ell},\ w_{u}$, respectively. The lower limit $w_{\ell}$ is simply the observed velocity $|u|$ for a uniform sphere if no obstruction occurs. From Figure 7, the line-emitting gas posterior to the source (redshifted in the case of an outflow) cannot be seen by the observer, meaning that $w_{\ell}$ must account for obstruction by the launching source at $x=1$. Solutions for the source obstruction velocity limit are given in Flury2023, which we summarize here. From the Pythagorean theorem, the minimum velocity $w_{\ell}$ must satisfy
$w_{0}=|u|\frac{x[w_{0}]}{\sqrt{x^{2}[w_{0}]-1}}$ (28)
giving a lower velocity limit of
[…]
where the ‘source effect’ applies only for the line-emitting gas posterior to the launch site relative to the observer (left half of the isocontours in Figure 6, of the sphere in Figure 7, and of the cartoon in Figure 10).
In principle, the upper velocity limit is simply the terminal velocity, i.e., $w_{u}=1$. However, the spectroscopic aperture might not capture the full extent of the outflow. This aperture clipping effect is most likely to affect emission line profiles as the core arises from extended material. OutLines incorporates aperture clipping into the geometry as some maximum velocity $v_{a}$ captured by the spectroscopic aperture on the transverse axis (plane of the sky) assuming that (i) the aperture is circular, as is the case for fiber spectrographs, and (ii) the outflow source is perfectly centered within the aperture. This effect can be imposed via the integral limits relative to the terminal velocity as $w_{a}=v_{a}/v_{\infty}$. The corresponding limiting velocity can be written as
$w_{1}=|u|\frac{x[w_{1}]}{\sqrt{x^{2}[w_{1}]-x^{2}[w_{a}]}}$ (30)
following the same approach as with $w_{0}$ in Equation 28. We show the change in $w_{1}$ for different $w_{a}$ in Figure 9 and illustrate the effect of aperture clipping for different $w_{a}$ in Figure 8.
The limit of $w_{a}$ most dramatically affects the projected velocities close to the transverse axis. As a result, aperture clipping will predominantly truncate the core of the profile close to the central wavelength (see also 2015ApJ...801...43S); i.e., the broadest extent of the wings remains relatively unaffected. Moreover, absorption lines arise from gas anterior to the launch site that is backlit by the source, resulting in a “pencil beam” effect where the maximum velocity contributing to the profile is $w_{0}$ (hatches in Figure 7). Combining $w_{1}$ and $w_{0}$ gives a final upper velocity limit of
$w_{u}=\begin{cases}w_{0}&\textrm{absorption}\\ w_{1}&w_{1}<1\\ 1&\textrm{otherwise}\\ \end{cases}.$ (31)
We visualize the integral limits $w_{\ell}$ and $w_{u}$ in Figure 9.
Figure 7: Spherically symmetric outflow presented in the text with a constant-velocity surface with a radius of $u$. Black arrows indicate the motion of the outflow. Central circle indicates the launching source of the outflow. Filled grey region indicates line-emitting material while the hatched region indicated line-absorbing material. Colored arrows indicate the direction of light propagation towards the observer. Colored circles indicate different deprojected velocities $w$ which can contribute to a single observed velocity surface $u$ with values of $u/w=0.15,0.5,0.85$ corresponding to yellow/green, orange/blue, and red/purple, respectively.
Figure 8: Effect of aperture clipping on the observed line profile for nebular emission (left panel) and for a P Cygni line consisting of absorption and resonant emission with unity filling factor and $p_{res}=0.5$ (right panel). For visualization, lines are normalized to the case of no aperture clipping ($w_{a}=1$). Note that the P Cygni line simplifies to pure absorption in the case where $w_{a}=0$ ($w_{1}=w_{0}$) as in 2015ApJ...801...43S and that the effect of radiative trapping on the line core decreases as $w_{a}$ decreases.
Figure 9: Integral velocity limits for the emission line profile $\phi_{\lambda}$. Lower limit $w_{\ell}$ is shown in grey with (dashed, $w_{0}$, Equation 28) and without (solid, $|u|$) the outflow source. Upper limit $w_{u}$ is shown in black with (dashed, $w_{1}$, Equation 30) and without (solid, unity) aperture clipping.
2.4.2 Directed Outflows
Many outflows are not spherically uniform but rather are observed as bi-directional cones on a wide variety of scales (e.g., Fukui1986; Clegg1987; Heckman1990; Colina1991; Hirano1992; Marlow1995; Cameron2021). We consider a bi-directional conical geometry with a half opening angle $\theta_{o}$ and an inclination angle $i$ with respect to the observer, which we depict in Figure 10 (see also, e.g., Figure 5 in Hjelm1996). OutLines includes a tool PlotGeometry for producing such figures as a visual aid.
Figure 10: Outflow geometry accounting for obstruction by the outflow source (white circle with dark grey shaded region) and by an optional disk (rectangle with light grey shaded region). Red arc indicates the inclination angle $i$ of the cone axis with respect to the observer. Green double arc indicates the opening angle $\theta_{o}$ of the cone between the axis and the edge. Orange triple arc indicates the optional cavity angle $\theta_{c}$ between the cone axis and the inner edge of the outflow.
The major difference with conical geometry is that the polar coordinate of the differential solid angle ${\rm d}\Omega$ now depends on $i$ and $\theta$; we cannot simply assume each deprojected velocity contributes equally to the integral due partial coverage. Therefore, we must account for azimuthal angle of the deprojected velocity which falls within the region of the cone projected onto the plane of the sky. The azimuthal angle inscribed by the projected cone simply scales the total solid angle by a factor $\ell_{\rm w}$$(u,w,i,\theta)\in[0,1]$ to give the relative contribution of a de-projected velocity $w$ to the observed velocity $u$. OutLines computes the conical projection from $i,\theta_{o}$ and then calculates $\ell_{\rm w}$ by determining the angular length(s) of each deprojected velocity inscribed by the projected cone. For details on this method, we refer the reader to Appendix A (see, e.g., Zheng1990; Carr2018; Luminari2018, for earlier approaches to solving this problem). We illustrate the effects of changing the opening angle $\theta_{o}$ and inclination angle $i$ in Figure 11.
Figure 11: Effects of bi-directional conical geometry on the emergent line profiles of emission (top) and absorption (bottom) profiles for fixed density (power law with $\alpha=2$) and velocity ($\beta=1$) fields. To illustrate the effects of geometry, we either vary the inclination with respect to the observer (left) from face-on ($i=0^{\circ}$) to edge-on ($i=90^{\circ}$) with fixed opening angle of $\theta=30^{\circ}$ or vary opening angle of the cones (right) from narrow ($\theta_{o}=30^{\circ}$) to a full sphere ($\theta=90^{\circ}$) with face-on inclination. The case of $i=0$ and $\theta=90^{\circ}$ (red line, right) is consistent with a filled sphere. For visualization, line profiles are all normalized to the spherical case ($i=0^{\circ}$, $\theta=90^{\circ}$).
2.4.3 Directed Outflows with Cavities
An additional geometry observed in young stellar objects (e.g., Padgett1999; Louvet2018; deValon2020), stellar winds (e.g. Sirianni1998), starbursts (e.g., Axon1978; Marlow1995; Heckman1995; Rich2010; Yoshida2010), and AGN (e.g., Phillips1983; Hjelm1996; Crenshaw2000a; Crenshaw2000b; Muller2011; Venturi2018) is that of hollow cones of outflowing material. We include this geometry, assuming some cavity angle $\theta_{c}$ and opening angle $\theta_{o}$ (see also models by Das2006; Muller2011; Bae2016; Luminari2024A&A). OutLines implements this geometry by subtracting the fractional angular length $\ell_{\rm cavity}$ of each deprojected velocity inscribed by the cavity from fractional the angular length inscribed by the total cone (see Appendix A). The difference will always be positive and greater than zero since $\theta_{o}>\theta_{c}$ by definition. We illustrate the impact of a hollowing of the outflow cone on the emergent line profile in Figure 12. Somewhat intuitively, an edge-on hollow-cone outflow loses flux from the least Doppler-shifted gas (truncating the core) while a face-on hollow-cone outflow loses flux from the most Doppler-shifted gas (truncating the wings).
Figure 12: Same as Figure 11 but for hollow bi-directional cones with cavity angle $\theta_{c}=\theta_{o}/2$.
2.4.4 An Obstructing Disk
The outflow may be launched from a region embedded in a disk of optically thick material, a phenomenon observed in a variety of astrophysical scenarios: expelled debris in planetary nebulae (e.g., Icke1981), protoplanetary material around young stellar objects (e.g., Padgett1999; Ilee2016), the dusty material concentrated around an AGN (e.g., Veilleux2001), or even the plane of a galaxy (e.g., Hjelm1996; Cronin2025). In such cases, the disk will obstruct some of the observer’s view of the outflow. OutLines optionally includes obstruction by a circular disk which is orthogonal to the central axis of the cones and which has some radius $x_{d}=r_{\rm disk}/R_{0}$. As with the cones, the disk must be projected onto the plane of the sky and subsequently compared to the deprojected velocity inscribed by the projected cone to determine the fractional arc length(s) $\ell_{\rm w}$ inscribed by the cone but not the projected disk. We elaborate on this formalism in Appendix A. Because a dense disk of material will likely suppress any outwardly moving material within its plane, OutLines only implements a disk if cones are present.
Figure 13: Same as Figure 11 but for bidirectional cones with fixed inclination and opening angle and an obscuring disk with varying radii, where a radius of 0 indicates a disk-free line profile. All profiles are normalized to the disk-free case for illustrative purposes. On the left, $i+\theta_{o}<90^{\circ}$, resulting in loss of the redshifted component. On the right, $i+\theta_{o}>90^{\circ}$, resulting in losses of both red- and blueshifted components when the disk radius is sufficiently high. For visualization, line profiles are normalized to the disk-free case in each panel.
We illustrate the effects of including an obstructing disk in Figure 13. By definition, the disk will only obstruct the posterior cone, so if $i+\theta_{o}<90^{\circ}$, the disk will only obstruct the receding outflow, solely obstructing the redshifted component (left of Figure 13). However, when $i+\theta_{o}>90^{\circ}$, the disk obstructs the lune projected by the posterior cone toward the observer, resulting in loss of the blueshifted component of the line profile as well (right of Figure 13).
2.5 Outflow Properties
Following Flury2023, the characteristic outflow radius $x_{out}$ corresponds to the maximum momentum density $\dot{\rho}_{p}=\mu m_{\rm H}n_{0}v_{\infty}n[x]w[x]$, i.e., where the wind effects are strongest. In the case of pulse-like density distributions, the center of the gas distribution $x_{1}$ will be close to, but not always the same as, $x_{out}$.The radius $x_{out}$ of maximum momentum density scales all other outflow properties, first and foremost the characteristic outflow velocity $v_{out}$ by determining the velocity at $x_{out}$ using the velocity field $w$ such that
$v_{out}=v_{\infty}w[x_{out}].$ (32)
The effective column density $N_{out}$ at the characteristic radius can be determined by integrating over the density profile from the base of the outflow ($x=1$) to the peak momentum density $x=x_{out}$ (e.g., Flury2023, their Equation 13 for the case of a CAK velocity field and power law density profile). Note that $N_{out}$ is not to be confused with $N$ from the absorption line profile, which is the total column density along the line of sight. $N_{out}$ necessarily depends on the space density $n_{0}$ and the radius at the base of the outflow $R_{0}$, which are unconstrained parameters in OutLines. As such, OutLines computes the relative column density $\mathcal{R}$ where
$\frac{N_{out}}{n_{0}R_{0}}=\mathcal{R}=\int\limits_{1}^{x_{out}}n[x]{\rm d}x.$ (33)
This quantity $\mathcal{R}$ becomes important for computing the strength of the outflow expressed via $\dot{M}$, $\dot{p}$, and $\dot{E}$.
The mass outflow rate $\dot{M}$ is given by the momentum flux $\mu m_{\rm H}n_{0}v_{out}$ with a geometric projection $4\pi R_{0}^{2}(\cos\theta_{c}-\cos\theta_{o})$ such that
$\dot{M}=4\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}x_{out}\mathcal{R}\mu m_{\rm H}v_{out},$ (34)
a more general case of Equation 12 from Flury2023. Such formalism is frequent in the literature with differences in implementation of the geometric projection (e.g. 2005ApJS..160...87R; 2005ApJS..160..115R; Crenshaw2012; Maiolino2012; Fiore2017; Luminari2018; Luminari2024A&A; Flury2023). The outflow momentum rate $\dot{p}$ expresses the rate at which momentum is injected into the outflow. Typically, $\dot{p}$ is defined as the product of outflow velocity and mass outflow rate (e.g., 2005ApJS..160..115R, their Equation 9) such that
$\dot{p}_{out}=v_{out}\dot{M}_{out}=4\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}v_{out}^{2}.$ (35)
The outflow mechanical luminosity $\dot{E}$ expresses the injection of kinetic energy into the outflow. Typically, $\dot{E}$ is defined as the product of outflow velocity and mass outflow rate (e.g., 2005ApJS..160..115R, their Equation 11) such that
$\dot{E}_{out}\begin{split}=&\frac{1}{2}\dot{M}\left(v_{out}^{2}+3\sigma_{v}^{2}\right)\\ =&2\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}\left(v_{out}^{3}+3v_{out}\sigma_{v}^{2}\right).\end{split}$ (36)
These values can be readily computed using constraints on $x_{out}$ and $v_{\infty}$ and possible geometric terms $\theta_{o}$ and $\theta_{c}$, all obtained from the line profile. The only two unknowns are $n_{0}$ and $R_{0}$, which cannot be directly informed by the line profile and require additional measurements (spatial for $R_{0}$, spectral for $n_{0}$) or assumptions based on prior knowledge. For instance, $n_{0}=10\rm~cm^{-3}$ and $R_{0}=0.5$ kpc may be reasonable for galactic outflows (e.g., 1985Natur.317...44C; Flury2023).
As implemented here, we compute $\dot{M}$, $\dot{p}$, and $\dot{E}$ normalized to $n_{0}R_{0}^{2}$ with $n_{0}$ in cm-3 and $R_{0}$ in kpc, meaning the characteristics of the outwardly moving gas are written as
$\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=0.191\mathcal{R}[\cos\theta_{c}-\cos\theta_{o}]\left[\frac{v_{out}}{\rm km~s^{-1}}\right]$ (37)
$\frac{\dot{p}}{\rm dyne}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=10^{-3.2}\left[\frac{v_{out}}{\rm km~s^{-1}}\right]\left[\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\right]$ (38)
$\frac{\dot{E}}{\rm erg~s^{-1}}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=10^{-6.5}\left[\frac{v_{out}}{\rm km~s^{-1}}\right]^{2}\left[\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\right]$ (39)
Typically, $n_{0}R_{0}^{2}$ is of order unity for the case studies in §4, with values ranging from about 1 to 5 cm-3 kpc2, although for Mrk 462, Flury2023 find $n_{0}R_{0}^{2}=12$ cm-3 kpc2.
3 The OutLines Code
OutLines, a portmanteau of outflows and spectral lines, implements all of the above model components in a single code for emission and absorption lines. The code is designed to be simultaneously efficient and user-friendly, with a straightforward user interface and ready scientific application. Below, we assess OutLines in the contexts of standard techniques, including empirical percentiles and multi-Gaussian fitting. Details of the architecture and use of the OutLines code can be found in Appendix C.
3.1 Empirical Percentiles
[…]
Figure 14: Velocity percentiles (black dashed lines) for an arbitrary emission (top) and absorption (bottom) line (purple) generated by OutLines for a directed, filled cone geometry (Geometry=’FilledCones’), default parameters, and otherwise default settings. The cumulative distribution function (fraction of integrated flux) is shown in red. For reference, we show the characteristic outflow velocity $v_{out}$ (Equation 32) in blue.
3.2 Modeling Line Profiles
Given that OutLines will sometimes require a large number of parameters (anywhere from four to nine for a given line depending on the geometry and density profile), we strongly recommend the use of Markov-chain Monte Carlo (MCMC) or other Gaussian-process samplers of the posterior on the parameters rather than the various least squares minimizers. Not only can this improve fit results but also allow users to incorporate informative priors, plus marginalization of particular nuisance parameter, which the trust-reflective algorithm, Levenberg-Marquardt, and other nonlinear least squares routines cannot. E.g., in the hollow cone scenario, $\theta_{o}>\theta_{c}$ necessarily and should be enforced.
To facilitate modeling of observed line profiles, the Profile classes contain several attributes and functions. Users can access the model line profile function directly via the Profile attribute, which receives rest wavelengths and parameters from the user and returns the total line profile predicted by the OutLines code. When instantiated, the Profile classes also contain initial parameter guesses and boundaries accessible via the get_params() and get_bounds() methods. For convenience, these two classes also contain the init_params function which will sample walkers within the current parameter bounds given the current parameter values. These initial parameter guesses can be readily seeded into an MCMC sampler while the boundaries can be folded into a prior. For this purpose, the log_probability() bound method includes both the likelihood and a uniform bounded prior incorporating the boundary values contained in the object. Users are able to update any or all of these bound methods as needed. An example of modeling with OutLines is provided in Appendix E.
While we encourage simultaneous fitting of multiple features from the same or very similar species, we caution against modeling lines from different species which may not necessarily share, say, the same density profile parameters due to ionization structure within a multiphase outflow.
OutLines does not presently include line-spread function (LSF) corrections in-house: such corrections are left to the user at this time. To assess the impact of the LSF on the recovery of line profiles, we generate $10^{4}$ mock observations varying $R$ from 1 000 to 10 000, $v_{\infty}$ from 100 to 1 000 $\rm km~s^{-1}$, and $\beta$ from 0 to 2.5 assuming the OutLines defaults. We convolve the predicted line profiles with a Gaussian LSF and add noise consistent with S/N between 3 and 10 in the continuum. We fit the mock observations using the scipy.least_squares trust reflective algorithm to attempt to recover the input parameters. Our simulation indicates that $R\sim 1500$ is sufficient to recover outflow properties to within $\la 15$% with a simple Gaussian LSF correction and $R\sim 3000$ to within 1010%, suggesting that facilities such as JWST/NIRSpec can be used with OutLines to recover key information about winds, bubbles, and outflows. Moreover, we find that $R\sim 6000$ can recover parameters within $\approx 5$% of the input values even without an LSF correction, suggesting LSF corrections in this regime are negligible.
Figure 15: Fit to a mock [O iii] $\lambda 5007$ outflow profile (see §3.2) with a double Gaussian (left) and with OutLines (right). While the fit quality is comparable in terms of $\chi^{2}$, the inferred fluxes and kinematics differ dramatically, leading to disparate scientific conclusions.
We illustrate the importance of using OutLines over traditional methods by comparing a fit to a mock [O iii] $\lambda 5007$ line profile profile at R=5  [^0] R=5\ 000 with Nyquist sampling and S/N=10=10 at the peak flux with OutLines and with a double Gaussian, both of which we show in Figure 15. For this purpose, we generate a mock line profile using OutLines for an acceleration power law velocity field with $\beta=1.5$ and $v_{\infty}=500$ $\rm km~s^{-1}$, power law density profile with $\alpha=1$, and filled cone geometry with $i=45^{\circ}$ and $\theta_{o}=30^{\circ}$. We add a static component with $\sigma_{v}=50$ $\rm km~s^{-1}$. Both the static and outflow components have equivalent integrated fluxes normalized to unity in the mock profile.
[…]
With this caveat in mind, we compare the results between the standard method and our new approach with OutLines in modeling an outflow. OutLines is able to recover all input parameters within the imposed 10% uncertainty due to noise. The double Gaussian model over-predicts the outflow flux by a factor of two and under-predicts the static gas flux by a factor of three while characterizing the outflow as red-shifted relative to the systemic velocity of the static gas. The double Gaussian model results lead to vastly different physical interpretations of the observations than those from OutLines: one would conclude the presence of a more massive outflow that is launched from a location traveling away from the static gas relative to the observer at $100$ $\rm km~s^{-1}$ but at characteristic velocities of $v_{out}=100$ $\rm km~s^{-1}$. One might even be inclined to interpret the double-Gaussian result as evidence for an inflow rather than an outflow, which is not the case. The OutLines best-fit model indicates an outflow launched from the same systemic velocity as the static gas with a bidirectional cone geometry and with characteristic velocities of $v_{out}=220$ $\rm km~s^{-1}$. Following Equation 34, accounting for the geometry (a factor of $1/(1-\cos\pi/6)\approx 7.5$), differences in line flux ($n_{0}\propto\sqrt{F_{0}}$ gives a factor $\sqrt{2}$), and difference in $v_{out}$ (factor of 0.5), the double Gaussian model yields a mass loss rate higher than that of OutLines by a factor of five. Thus, while the velocity inferred from the double Gaussian model is half that from OutLines, the double Gaussian model implies a substantially higher impact of the outflow on its environment than the OutLines model.
3.3 Absorption vs Emission
A key result of the development and applications of OutLines is that absorption and emission lines optimize constraints on different properties of outflows. Absorption lines are more sensitive to the line-of-sight, as evidenced in the line profiles (Figures 2 and 4) and the narrow range in velocities contributing to each wavelength bin due to the “down-the-barrel” effect (Figures 7 and 9), making them good probes of the density profile and velocity field but poor tracers of the geometry. Emission lines trace the global gas distribution since velocities from across the majority of the gas contribute to each velocity bin (Figure 9), making them good probes of the geometry (Figures 11, 12, 13) but, as a result, lose some constraining power on the density profile and velocity field. As such, combined modeling of both emission and absorption provides optimal constraints on the outflow properties.
4 A High Capacity for Science
Below, we demonstrate application of OutLines to a variety of phenomena on different spatial and temporal scales, all involving astrophysical winds: stellar wind bubbles in the brightest cluster of the giant extragalactic H ii region NGC 5471, radiative feedback and Lyman continuum escape from a super star cluster in Green Pea J1044+4012, blowout and the formation of a directed outflow in starburst Mrk 1486, and thermal/cosmic ray pressure from AGN-shocked bubbles in the Seyfert 2 nucleus of NGC 2992. Results from modeling these examples with OutLines are given in Table 1 and are discussed on a cases-by-case basis in the subsections below.
All modeling with OutLines is done here using emcee to sample the posterior on the parameters. The MCMC chain consists of a number of walkers equal to twice the number of free parameters plus one (the OutLines default), with each walker taking $10^{3}$ steps with an additional $200$ burn-in steps given the number of walkers (minimum of 10 walkers for at least 2000 individual burn-in steps), which we find to be sufficient to “forget” the initial guess parameters. The initial guess parameters, unless otherwise stated, are obtained using the scipy.optimize.least_squares trust-reflective algorithm fit. The prior is assumed to be uniform bounded unless otherwise stated. We show posterior sampling for each case in Appendix F.
Table 1: Best-fit parameters and properties of test-case objects: Knot A of giant extragalactic H ii region NGC 5471, super star cluster in the center of Green Pea J1044+3053, star-forming galaxy Mrk 1486, and Seyfert 2 NGC 2992. As discussed in §2.5, reported properties of $\dot{M}$, $\dot{p}$, and $\dot{E}$ are normalized to $n_{0}R_{0}^{2}$ with $n_{0}$ in units of cm-3 and $R_{0}$ in kpc (see Equations 37-39).
Object NGC 5471 Knot A J1044+0353 SSC Mrk 1486 NGC 2992 Sy2 nucl.
Scenario H ii region winds Lyman continuum escape galaxy-scale outflows AGN disk winds
[…]
OutLines model BetaCAK, PowerLaw, Sphere BetaCAK, PowerLaw, Sphere BetaCAK, DampedPulses, Sphere AccPlaw, PowerLaw, HollowCones, Disk
Parameters from MCMC Posterior Samples
$\alpha$ $0.98$  $\pm$ $0.31$ $1.39$  $\pm$ $0.76$ $1.42$  $\pm$ $0.37$
$\beta$ $4.07$  $\pm$ $1.24$ $2.27$  $\pm$ $1.10$ $0.71$  $\pm$ $0.25$ $2.48$  $\pm$ $0.42$
$v_{\infty}$ $\rm km~s^{-1}$] $396$  $\pm$ $75$ $213$  $\pm$ $26$ $624$  $\pm$ $38$ $1034$  $\pm$ $108$
$b$ or $\sigma_{v}$ $\rm km~s^{-1}$] $28$  $\pm$ 11 $34$  $\pm$ 2020 $50$  $\pm$ 2020 $77$  $\pm$ 11
Properties from MCMC Posterior Samples
$W_{50}$ (“FWHF”) $\rm km~s^{-1}$] $49$  $\pm$ 11 $87$  $\pm$ 44 $181$  $\pm$ $27$ $127$  $\pm$ 44
$W_{80}$ $\rm km~s^{-1}$] $109$  $\pm$ 33 $147$  $\pm$ 66 $305$  $\pm$ $77$ $289$  $\pm$ $12$
$x_{out}$ $5.16$  $\pm$ $2.04$ $2.63$  $\pm$ $^{1.78}_{1.09}$ $1.24$  $\pm$ $0.17$ $1.35$  $\pm$ $0.02$
$v_{out}$ $165$  $\pm$ $15$ $72$  $\pm$ $17$ $188$  $\pm$ $78$ $597$  $\pm$ $267$
$\dot{M}/n_{0}R_{0}^{2}$ $\rm M_{\odot}~yr^{-1}$] $105$  $\pm$ $59$ $22.1$  $\pm$ $^{33.6}_{14.2}$ $9.8$  $\pm$ $3.2$ $9.5$  $\pm$ $0.4$
$\dot{p}/n_{0}R_{0}^{2}$ $\rm 10^{34}~dyne$] $10.9$  $\pm$ $7.2$ $1.00$  $\pm$ $^{2.21}_{0.71}$ $0.95$  $\pm$ $0.54$ $3.59$  $\pm$ $0.25$
$\dot{E}/n_{0}R_{0}^{2}$ $\rm 10^{42}~erg~s^{-1}$] $0.90$  $\pm$ $0.69$ $0.04$  $\pm$ $^{0.11}_{0.03}$ $0.09$  $\pm$ $0.07$ $1.07$  $\pm$ $0.11$
4.1 Evolution of H ii Regions
H ii regions are dynamic environments with radiation and winds from O and B stars shaping and enriching the proximal ISM. Winds launched from stars carry out recently formed metals dredged up to the surface and subsequently sweep up surrounding material. Radiation heats gas, often through photoionization, and imparts momentum to the ISM, which can result in additional launching or driving of winds, bubbles, and outflowing gas (e.g., Castor1975b; Weaver1977; Dyson1977; Dyson1979; Oey1997; Dopita2005). While stellar population models often contain predictions for these feedback effects (e.g., 1999ApJS..123....3L; starburst99; Eldrige2017; Byrne2022; Hawcroft2025), empirically constraining the kinematics and related mechanical feedback in H ii regions remains an active area of research (e.g., models and simulations by Geen2020; Geen2023; Lancaster2025a; Lancaster2025b) with implications for understanding chemical enrichment (e.g., Esteban1995; LopezSanchez2007; Stock2011) and dispersal timescales (e.g., Dopita2005; Menon2024).
One such kinematically rich and complex system is the giant extragalactic H ii region NGC 5471 located in M101. Many knots within this region exhibit broad and sometimes asymmetric nebular lines which cannot be described by a single Gaussian (e.g., Chu1986; Bresolin2020). To investigate the evolution of H ii regions in this case-study, we apply OutLines to the [S ii] $\lambda\lambda 6716,6731$ doublet, which traces the density and kinematics of low ionization gas, in Knot A of NGC 5471, the brightest knot identified by Skillman1985, as observed with the $R=6000$ MEGARA IFU spectrograph on the Gran Telescopia Canarias (Arellano-Cordova in prep, PI Esteban). For our modeling, we use a 2.5′′ aperture extraction of the Knot A spectrum from the IFU cube. We obtain atomic data from NIST (NIST_ASD) for wavelengths (Kaufman1993) and spontaneous emission coefficients (transition probabilities) (Podobedova2009). We use collision strengths from Tayal2010. First, we fit the adjacent featureless continuum with a third-order polynomial, an appropriate approximation in the case of extended H ii regions which do not suffer appreciably from contamination such as stellar absorption except near permitted lines from, e.g., H i and He i. We subtract this continuum before modeling the line profile with OutLines with the default options. Then, we sample the posterior on the OutLines profile parameters to constrain the outflow and static gas properties. No LSF corrections are necessary as $R=6000$ is sufficiently high to constrain the kinematics (see §3.2). We compare the best-fit model to the observed line profiles in Figure 16.
Figure 16: Top: [S ii] $\lambda\lambda 6716,6731$ doublet in Knot A of NGC 5471. Grey dashed vertical line indicates the line center given by atomic data. The best-fit OutLines.Nebular model is shown in red with the outflow (dashes), static (dotted), and total (solid) line profiles. Bottom: residual difference between the best-fit OutLines profile and the observed spectrum.
From the best-fit line fluxes, we determine the electron density of both the nebula and outflow components from the [S ii] 6716/6731{6716}/{6731} flux ratio using PyNeb (2015A&A...573A..42L). The static gas flux ratio of 1.459 is in the low-density limit, requiring $n_{e}<10\rm\ cm^{-3}$. For the outflowing gas component, the doublet ratio is 0.831, which is where [S ii] is most sensitive simultaneous to density and to temperature. For a range in $T_{e}$ from 10 to 15 $\rm kK$, the inferred outflow density is anywhere from $970$ to $1140\rm~cm^{-3}$, with density values increasing with the assumed temperature. The implication from these densities is that the outflow is roughly two orders of magnitude more dense than the gas within the static nebular gas. While this increase in density may be due to a sweeping up of material in the wind, this dramatic change in density coupled with the supersonic flow could instead arise from compression by shocked gas as it rapidly cools via radiation (e.g., Flury2025b).
The slope of the velocity field, $\beta=4$, corresponds to gradual acceleration likely caused by optically thick radiation pressure from early type stars (e.g., 1977ApJ...213..737B). While Bresolin2020 argue that the non-thermal radio SEDs of Sramek1986 are clear-cut evidence for supernova-driven bubbles, the line profile shape indicates a continuous wind rather than shells or bubbles and, from modeling with OutLines, disfavors the supernova scenario. Indeed, the OutLines results are very much in agreement with the fact that Knot A features the a very young ($3\rm~Myr$) cluster–in fact the youngest in the entirety of NGC 5471 (e.g., GarciaBenito2011, based on HST photometric color-magnitude diagrams). Knot A is also very massive ($10^{4}\rm~M_{\odot}$) and requires a significant population of O stars to produce the $Q=4\times 10^{51}\rm~phot~s^{-1}$ associated with the star cluster (GarciaBenito2011). The age and mass of the stellar population indicate sufficient radiation available in the UV to radiatively drive a supersonic wind. Shocks triggered by this wind could explain the non-thermal emission in the radio in addition to the dense gas. Inhomogeneities in the gas distribution can readily dissipate the outward force due to pressure from the hot shocked gas (e.g., Dyson1977), allowing radiation pressure to dominate the acceleration of material out from the H ii region. Our results indicating escape of hot gas are consistent with the HarperClark2009 model for the Carina Nebula and the Lopez2011 multiwavelength study of the massive extragalactic H ii region 30 Doradus, both of which found strong evidence for radiation pressure driving and a deficit of X-ray flux associated with shocked gas.
With the velocity field and line fluxes inferred by modeling nebular lines with OutLines, we have obtained key insight into the nature of the outflowing material: (i) that it is a wind, not a bubble, (ii) that the wind exhibits gas densities far higher than that of the static nebula, and (iii) that the wind is driven by substantial radiation pressure from O stars. These results are consistent with the observed stellar population ages. Moreover, this wind is supersonic, which could explain the non-thermal emission and high gas densities via shocks.
4.2 Lyman Continuum Escape from a Super Star Cluster
Starburst galaxy J1044+0353 is a low mass ($M_{\star}\approx 10^{6.8}\rm~M_{\odot}$) Lyman continuum emitter candidate with $f_{esc}^{\rm Lyc}$ as high as 60% (Parker2025) and, as a nearby ($z=0.01287$) Green Pea, is a local analog for star-forming galaxies at cosmic dawn that are actively involved in re-ionizing the Universe (Martin2024). Recent IFU observations of J1044+0353 reveal expanding shells of gas at the periphery of the nebular gas flux distribution (Martin2024), indicative of blowout due to SNe from an aging stellar population 10-20 Myr after a burst of star formation (Peng2023). However, the central core of the galaxy hosts a super star cluster with several young (<2<2 Myr) stellar populations which may be responsible for driving additional feedback on shorter timescales (Olivier2022), providing a unique laboratory to explore the impact of young stars on their immediate environment in the context of stochastic, or bursty, star formation using OutLines.
To undertake this investigation, we draw on archival HST/COS G130M FUV spectroscopy (R=10  [^0] R=10\ 000) of the central starburst of J1044+0353 from the COS LegAcy Spectroscopic SurveY (CLASSY, CLASSY; CLASSY2, GO 15840), which allows for detailed modeling of outflow properties along the line of sight without significant effects from the LSF even after downsampling to the 0.073 Å resolution (see CLASSY; CLASSY2; Parker2024). The C ii $\lambda 1334$ absorption line is one of the strongest low ionization lines in the FUV due to the transition oscillator strength and comparatively high ionic abundance, making it an excellent tracer of outflows. Visual inspection of the C ii line profile suggests a classic blue-shifted continuous outflow plus a static ISM component (see, e.g., Figure 1 of 2015ApJ...809..147H). To obtain the normalized flux, we fit the starlight continuum within 3 Å of the feature with a third-order polynomial and divide the local spectrum by this fit (see, e.g., the VoigtFit software, VoigtFit, which takes a similar approach using Chebyshev polynomials).
For modeling the C ii line with OutLines, we use atomic data from Morton2003. No infilling is apparent in the observed line profile, suggesting either maximal aperture clipping where pure absorption is an appropriate description of the emergent feature (see Figure 8) or an optically thick wind. For simplicity, we assume the OutLines defaults, although we note that applying more sophisticated geometry or density profiles negligibly impacts the quality of the fit. Indeed, a filled cone model yields best-fit inclinations of $i\to 0^{\circ}$ and opening angles of $\theta_{o}\to 90^{\circ}$, consistent with the filled sphere geometry. Additional consideration of pulse-like density profiles consistently favors a peak location at the base of the outflow with wide dispersion, which combined are consistent with the power law approximation. Best-fit parameters are in Table 1.
[…]
Figure 17: MCMC fit of an OutLines.Absorption model to the C ii $\lambda 1334$ absorption line observed by HST/COS G130M (black, CLASSY). Line styles and colors as in Figure 16. Blue line indicates the intrinsic profile convolved with the HST/COS G130M line spread function (LSF).
The best-fit parameters indicate that the starbursting core of J1044+0353 contains a spherically expanding, isothermal wind with a velocity field characteristic of optically thick, radiatively-driven winds (e.g., 1977ApJ...213..737B; Flury2023). However, the relatively low wind velocity ($v_{\infty}=215$ $\rm km~s^{-1}$, comparable to Voigt modeling of the silicon lines, Parker2024, and the “broad” component of nebular lines, Peng2023) and weak $\dot{p}$ and $\dot{E}$ (marginally non-zero) suggests that mechanical feedback from massive stars is relatively weak or even suppressed. Such suppression can occur in cases where catastrophic cooling of swept up dense material prevents acceleration up to more “traditional” velocities closer to 500 $\rm km~s^{-1}$ or more, is known to occur in Green Pea galaxies with young stellar populations (2017ApJ...851L...9J), and could be associated with Lyman continuum (LyC) escape (e.g., 2017ApJ...851L...9J; Jaskot2019). Combined with evidence for recent SNe, the weak feedback inferred from OutLines models supports the bursty star formation scenario for LyC escape suggested by Flury2022b in their assessment of a large sample of star-forming galaxies with LyC detections and later demonstrated by Flury2025a with measurement of stellar populations and ISM conditions in stacks of UV spectra of LyC emitters. In their framework, SNe clear out channels in the ISM via mechanical feedback and a subsequent generation of stars 8-10 Myr later provides ionizing–but not mechanical–feedback to establish the density-bounded conditions necessary for Lyman continuum photons to escape anisotropically. Here, inferring underlying physics and wind properties from a single-aperture spectrum demonstrates the power and effectiveness of OutLines in understanding the role of mechanical feedback in LyC escape.
4.3 Outflow Formation and Blowout in a Starburst Galaxy
Mrk 1486 is a starburst galaxy exhibiting a tentative He ii $\lambda 4686$ Wolf-Rayet feature (e.g., Izotov1997), metal rich extraplanar biconical outflows beyond the galactic plane (e.g., Duval2016; Cameron2021), and >14>14 Myr stellar populations (Parker2025, although see Peng2025, who find from the same UV spectra that these populations are even older, $\approx 100$ Myr). The archival HST/COS spectrum from RiveraThorsen2015 exhibits the classic blue-shifted asymmetric feature associated with outflows, albeit with at least two troughs present in the outflow feature across several low ionization lines (Si ii $\lambda\lambda 1191,1193,1304$, O i $\lambda 1302$, and Al ii $\lambda 1670$, see RiveraThorsen2015 their Figure 3, Parker2024, their Figure 7). A multi-troughed absorption feature can indicate a ensemble of shells produced by a series of episodes or a system of many star clusters (see Figure 5, also RiveraThorsen2015).
For our analysis, we use the RiveraThorsen2015 spectrum reprocessed by CLASSY; CLASSY2, down-sampled to 0.073 Å resolution and normalized following the polynomial approach in §4.2. To determine the origin of this multi-episode outflow, we can use OutLines to model the absorption lines. The Si ii $\lambda\lambda 1191,1193$ doublet provides an excellent probe as both features arise from the same ion, thus providing additional constraining power. The stronger Si ii $\lambda 1260$ line is contaminated by telluric O i $\lambda 1302$ and cannot be used to constrain the fit. We note the presence of weak Si ii $\lambda 1194,1197$ fluorescent emission arising from the re-emission of Si ii $\lambda\lambda 1191,1193$ photon absorption, respectively. These lines suffer from apparent Milky Way contamination in the blue wings due to absorption by Galactic O iv $\lambda\lambda 1234,1237$, Si iii $\lambda 1235$, and Mn ii $\lambda\lambda 1234,1236$. We mask these features in order to include the fluorescent lines in our fit. Using atomic data from Kelleher2008 obtained from NIST (NIST_ASD) and the appropriate COS G130M LSF, we model the Si ii doublet complex of $\lambda\lambda 1191,1193$ resonant absorption and emission and $\lambda\lambda 1194,1197$ fluorescent emission using OutLines assuming a damped-pulses density profile to account for the multiple absorption troughs and otherwise adopting the default options of a spherical geometry and the $\beta$ CAK velocity field. The choice of spherical geometry is supported by the presence of relatively symmetric fluorescent Si ii $\lambda 1194,1197$ lines, although the COS LSF may complicate any inference of the geometry. As in §4.2, we incorporate the appropriate LSF corrections.
The two Si ii absorption lines have similar depths. Given the oscillator strengths of the two transitions, the Si ii $\lambda 1193$ line should be deeper than Si ii $\lambda 1191$. For the lines to have similar depths, the Si ii $\lambda 1193$ line must be saturated. Whether the Si ii $\lambda 1191$ line is saturated is not ascertainable solely from the relative line depths; however, the prominent $\lambda\lambda 1194,1197$ lines suggest optically thick lines consistent with trapping of the resonant emission, with many scattering events allowing photon escape solely through fluorescent channels, i.e., no resonant emission. Because $\lambda 1193$ is saturated and has a depth similar to that of $\lambda 1193$ (minimal infilling by the $\lambda 1194$ line), any transmitted flux must be due to starlight within the spectroscopic aperture that is not associated with the foreground gas. Traditionally, such residual flux is explained by a partial gas covering fraction, with the radiative transfer rewritten as
$\frac{F_{\lambda}}{F_{\lambda 0}}=1-C_{f}\left[1-\exp(-\tau_{0}\phi_{\lambda})\right]$ (40)
where $C_{f}$ expresses the fraction of background light which is not obstructed by the foreground gas (e.g., Netzer1982; Arav1999; Arav2005; RiveraThorsen2015). The COS LSF can affect the partial covering fractions of saturated absorption lines (see, e.g., assessment by Flury2025a, their Appendix D, also Figure 3 of Jennings2025). Incorporating both $C_{f}$ and $\epsilon$ as free parameters into the initial fit, we find that while the COS LSF and infilling by the $1194$ fluorescent line can account for some of the line profile depths, $C_{f}=0.82\pm 0.03$ is necessary to provide an optimal description to the absorption features (comparable to the $C_{f}=0.8$ obtained by RiveraThorsen2015). We also find $\epsilon=0.21\pm 0.03$ necessary to describe the fluorescent line profile.
We adjust the prior to require the column densities of each line component to agree within 0.05 dex standard deviation of the column densities of the same components in the other line, effectively enforcing the atomic data. Based on visual assessment of the line profiles, we further impose a Gaussian prior requiring the terminal velocity to agree with the apparent velocity at which the blue profile wing reaches the continuum ($\approx 400$ $\rm km~s^{-1}$, accounting for the LSF) and the central velocity of the first pulse to agree with the velocity of the deepest trough ($\approx 130$ $\rm km~s^{-1}$), both to within a standard deviation of four resolution elements. We also impose that each shell is relatively thin by requiring the width of each shell to be $15\pm$ 15% of the launch radius. We show the best-fit line profile in Figure 18 and list the results in Table 1. As a test of our assumption of spherical geometry, we predict the emergent fluorescent Si ii $\lambda\lambda 1194,1197$ lines using OutLines.Fluorescent and the best-fit parameters from the posterior, which we show in Figure 18.
Figure 18: MCMC fit of an OutLines model to the Si ii $\lambda\lambda 1191,1193,1194,1997$ doublet complex observed by HST/COS G130M (black, RiveraThorsen2015, GO 12583, PI Hayes). Shaded regions indicate suspected Milkyway contamination from Galactic Si iii, Mn $i$, and and O iv. Si ii $\lambda\lambda 1191,1193$ resonant emission lines computed using OutLines.Resonant are shown in orange, demonstrating strong radiative trapping. Si ii $\lambda\lambda 1194,1197$ fluorescent lines computed using OutLines.Fluorescent are shown in yellow. Total line profile corrected for the line spread function (LSF) is shown in blue. Other line styles, symbols, and colors as in Figure 16.
The posterior sampling indicates relatively small uncertainties on the order of 10% for each of the expanding shells, with the most prominent and lowest velocity shell having $v_{\rm shell}=188\pm 20$ $\rm km~s^{-1}$ (obtained from $v_{\rm shell}=v_{\infty}w[x_{1}]$ and propagating the uncertainties in $x_{1}$, $\beta$, and $v_{\infty}$). The system of shells suggest an outflow in formation, with regular events forming thin expanding superbubbles. Notably, the less pronounced absorption troughs in C ii $\lambda 1334$, which traces a higher ionization state than Si ii, O i, and Al ii, suggests that the ionized phase of the outflow is established more rapidly than the neutral phase and may indicate that the expanding bubbles are entrained in or even driven by a hotter phase of the ISM. The accumulation of superbubbles close to the terminal velocity suggests that an outflow forms from these successive events with sufficient speed to exceed the escape velocity ($v_{esc}\approx$100 $\rm km~s^{-1}$as implied by the $M_{\star}=10^{9.3}\rm~M_{\odot}$ from 2018MNRAS.481.1690C). The recurrance of many such expanding shells could produce the biconical outflow observed in emission once the superbubbles escape the gravitational potential and blow out from the galactic disk (see spatially resolved cases in, e.g., Martin1998; McQuinn2019, and Duval2016 specifically regarding Mrk 1486). Moreover, the successive events are removing gas at mass outflow rates of $\dot{M}=7.9\rm~M_{\odot}~yr^{-1}$, which may temporarily rob the galaxy of baryonic material necessary to sustain the production of stars and could explain the 30-40 Myr lapse between bursts of star formation.
Given that the outflows in Mrk 1486 are oxygen rich relative to their host galaxy (Cameron2021), it is useful to consider whether the outflow component of Si ii is metal-enhanced since, like oxygen, silicon is an $\alpha$ element and is predominantly produced by core-collapse supernovae (CCSNe, e.g., Kobayashi2020). From the best-fit line profiles from OutLines, the outflow $N_{o\ \rm Si^{+}}=10^{15.5}\rm~cm^{-2}$ while the static $N_{s\ \rm Si^{+}}=10^{13.5}\rm~cm^{-2}$. Ionization structure may account for some of the difference in Si+; nevertheless, the 2 dex increase in the outflow likely suggests substantial metal loading of the outflow (see, e.g., 2018MNRAS.481.1690C) due to enrichment by recent CCSNe.
It may be tempting to conclude that hot gas or cosmic rays from these supernovae are driving the shells’ expansion, especially since, for the lowest velocity shell at 169 $\rm km~s^{-1}$, the dynamical time to traverse the inner 500 pc is $\la 4$ Myr, well after the onset of CCSNe for a burst at least 10 Myr old. That being said, the age reported by Parker2025 from fits to the UV continuum that Mrk 1486 corresponds to a combination of a substantial young ($\leq 5$ Myr) stellar population and a far older (40 Myr) population. In light of the more detailed star formation history, while CCSNe may be the progenitors of the material in these expanding shells, they may not necessarily be responsible for their acceleration. Here, the shape of the velocity field from OutLines can provide insight into the driving mechanism. Indeed, the velocity index of $\beta=0.7\pm 0.2$ disfavors (but cannot fully not reject) the rapid acceleration associated with ram pressure driving ($\beta\la 0.5$, e.g., 2005ApJ...618..569M; 2016MNRAS.463..541C). However, these $\alpha$-rich shells are more likely continuously accelerated by optically thin line-driven radiation pressure associated with $\beta=0.8$ (e.g., 1986A&A...164...86P). Moreover, the total energy injection rate of $10^{40}\rm~erg~s^{-1}$ for the series of shells is roughly consistent with stellar winds, which have $L_{mech}\approx 10^{39}\rm~erg~s^{-1}$ for a given stellar population (see, e.g., Starburst99 or BPASS, starburst99; Byrne2022, respectively). Radiative driving by young stellar populations may also account for the less pronounced C ii troughs via ionizing feedback. Furthermore, given that the travel time across the central kpc is $\la 4$ Myr, the >40>40 Myr stellar populations are unlikely to be responsible for these superbubbles. Corroborating evidence for optically thick radiation pressure is the presence of the fluorescent $\lambda 1194,1197$ lines but the absence of resonant line infilling. Such a scenario occurs in the optically thick line limit where photon escape primarily occurs through the fluorescent channel (e.g., 2015ApJ...801...43S, their Figures 5-6).
The bursy star formation in Mrk 1486 could optimize the production of $\alpha$ elements and energetic expanding shells via CCSNe from earlier generations of stars while the most recent stellar populations provide the momentum injection via radiation pressure to form and accelerate optically thick enriched superbubbles, driving them beyond the galactic disk to form an outflow. OutLines is necessary here to infer the velocities of the shells as they accelerate under a collective driving mechanism, determine the relative importance of SNe and young stars in producing and accelerating these shells, and whether the shells are capable of eventually exceeding the escape velocity to form an outflow.
4.4 Thermal Winds in AGN Radio Lobes
The Seyfert 2 galaxy NGC 2992 has long been known to exhibit conical outflows (Allen1999; Veilleux2001). Using archival observations from the Siding Spring Seyfert Spectroscopic Snapshot Survey (S7, S7DR2) made using WiFES (WiFES), we have confirmed broad, asymmetric [O iii] $\lambda\lambda 4959,5007$ lines after summing over all the spaxels to boost signal to noise in the line profile and to mitigate any aperture effects. The low resolution ($R\approx$ 3 000 or FWHM of 105 $\rm km~s^{-1}$) requires corrections for the LSF by convolving with a Gaussian kernel with $\sigma_{\rm LSF}=$45 $\rm km~s^{-1}$; however, we note that the LSF predominantly impacts the narrow Gaussian core. The high signal in the wings ($S/N\approx 3$ in the blue broad extent of [O iii] $\lambda 4959$) might allow for tight constraints on the wind properties even without LSF corrections – indeed fitting with and without LSF corrections yields similar results for the OutLines parameters. Nevertheless, incorporating the LSF into the forward modeling with OutLines is essential to remain consistent with the observations. As a part of the LSF correction, we fold in the summed best-fit stellar continuum included in S7 to account for stellar contributions to the continuum, including any contamination of the line profile.
The bright intermediate portion of the wings suggests the acceleration power law for the velocity field. Visual assessment of the highly asymmetric line profile in the spectrum of NGC 2992 and comparison with models in Figures 11, 12, and 13 also suggests a hollow cone geometry with an obstructing disk. As such, we subsequently model the [O iii] doublet using OutLines assuming an acceleration power law, an open cone geometry, and an obstructing disk. We impose the following into the prior: $v_{\infty}$ within a 300 $\rm km~s^{-1}$ standard deviation of the apparent maximum velocity of 1100 $\rm km~s^{-1}$, a static velocity dispersion within 50 $\rm km~s^{-1}$ standard deviation of the apparent dispersion of 110 $\rm km~s^{-1}$, accounting for the LSF, line fluxes within 5% of the [O iii] 5007/4959 flux ratio of 2.98 given by atomic data (Storey2000), and angles within 10 standard deviation of the best guess values from the trust-reflective fit. With line profiles predicted by OutLines added to the continuum and subsequently convolved with the WiFES LSF (WiFES), we use emcee to sample the posterior on the outflow parameters, requiring $\theta_{o}-\theta_{c}>5^{\circ}$. Due to the large number of parameters, potential degeneracies, and sensitivity to initial conditions, we run the sampler four times longer, for a total of 40 000 samples post burn-in. Implementation of our modeling is demonstrated for reference in Appendix D with the best-fit profile shown in Figure 19 and results listed in Table 1. We note that the posterior sampling ultimately rejects the priors on the angles as the best-guess parameters fell within a local minimum.
Figure 19: MCMC fit of an OutLines.Nebular model to the [O iii] $\lambda\lambda 4959,5007$ doublet observed in the nucleus of Seyfert galaxy NGC 2992 as part of the Siding Spring Seyfert Spectroscopic Snapshot Survey (S7, S7DR2) DR2, assuming a hollow cone plus disk outflow geometry. Line styles and colors as in Figure 16. Blue line indicates the profile after LSF correction.
[…]
The slope $\beta=2.48$ indicates, for the acceleration power law velocity field, rapid acceleration to the terminal velocity (see Figure 2), which can be interpreted as thermal (ram pressure) or cosmic ray (magnetic pressure) driving as each can produce similar rapid-acceleration velocity fields based on the shape of the velocity field (see Figure 1, also 2005ApJ...618..569M; 2016MNRAS.463..541C; Flury2023). Thermal driving occurs when “cool" ($10^{4}\rm~K$) clouds are embedded in an adiabatically expanding hot ($10^{6}\rm~K$) gas (2005ApJ...618..569M). Cosmic ray occurs driving where clouds are accelerated by the diffusion (Breitschwerdt1993; Quataert2022a) or streaming (Ipavich1975; Breitschwerdt1991; Quataert2022b) of relativistic charged particles–launched by hot ionized media such as shocks or supernovae (Colgate1979; Blandford1980)–as they interact with magnetic fields in the ISM, resulting in a radially outward pressure gradient. NGC 2992 is known to contain $\approx 0.5$ kpc radio lobes forming an hourglass or “figure eight” shape, indicative of bi-directional bubbles of hot gas driven by nuclear activity (Wehrle1988; Allen1999). Combining optical and radio measurements indicates the presence of hot cavities which are heated (likely through shocks) by jets or accretion disk winds and within which are entrained clouds of cooler gas (Allen1999; Veilleux2001). The presence of such lobes and the associated hollow cavities suggests either or both (a plausible scenario, e.g., Breitschwerdt2002; Everett2008) mechanisms are involved in launching the outflow in NGC 2992.
Spatial analysis of the outflow in Veilleux2001 found an inclination of $i=68\pm 3^{\circ}$ and an opening angle of $\theta_{o}=62.5-67.5^{\circ}$. OutLines finds $i=53.19\pm 1.32^{\circ}$ and $\theta_{o}=68.60\pm 1.26^{\circ}$. While $\theta_{o}$ is in apparent agreement, ii is significantly discrepant, requiring further assessment of the fit results. To test the reliability of the geometry inferred from the OutLines model, we project the best-fit bidirectional cones onto the plane of the sky and compare with the spatially resolved continuum-subtracted [O iii] doublet flux in the WiFES IFU in Figure 20. We find good agreement between the cone geometry implied by the integrated line profile using OutLines and the observed flux distribution, indicating that modeling with OutLines can recover the spatial geometry without any knowledge of that geometry apriori. Not only have we accurately recovered the spatial flux distribution with OutLines, but we have also inferred the presence of a cavity comprising the inner two thirds of the outflow cone ($\theta_{c}=48.64\pm 1.82^{\circ}$), suggesting that cooler line-emitting clouds can only survive towards the edges of the hot cavity where the thermal wind may be somewhat cooler as it expands and contacts the surrounding ISM.
Figure 20: Greyscale map across the central kiloparsec of NGC 2992 of the continuum-subtracted, integrated [O iii] doublet flux in the S7 DR2 (S7DR2). Drawn cones indicate the geometry inferred from the best-fit OutLines profile to the integrated emission line profile. No spatial information was used to determine the shape of the projected cones. Blue indicates the anterior cone, red indicates the posterior cone. Indigo indicates the disk, here with arbitrary radius scaled for visualization. A PA of 75 was assumed to rotate the projected geometry to the PA of the observation.
With the best-fit parameters, OutLines computes high mass loss rates of at least 40 $\rm M_{\odot}~yr^{-1}$. From the $v_{0}\sim 300$ $\rm km~s^{-1}$ rotational velocity of the galaxy (Veilleux2001), the outflow must be capable of ejecting material from the host galaxy with the characteristic velocity exceeding the escape velocity by a factor of 3. Together, these results suggest NGC 2992 is in a state of AGN-triggered inside-out quenching with the combined effects of hot cavities and the evacuation of cool gas away from the plane of the galaxy, both effects preventing future star formation. Interestingly, this quenching feedback is not sufficiently columnated to be driven by a jet but instead is most likely due to expanding gas heated by winds launched from the accretion disk. While the implications of quenching are clear for galaxy evolution, the association of hot winds and shocks also has implications for the measurement of chemical abundances in AGN (e.g., Dors2021). Here, inferring underlying physics, geometry, and outflow strength from a single-aperture spectrum demonstrates the power and effectiveness of OutLines in the context of AGN feedback.
5 Conclusion
Here, we present a novel method for modeling emission and absorption lines arising from astrophysical winds, bubbles, and outflows. Our method has many advantages over previous methods, such as multi-Gaussian and Gauss-Hermite models or velocity quantiles, in that it is physically motivated but also agnostic to underlying conditions or assumptions about the driving mechanism(s). Moreover, our model is readily applicable to absorption and emission line profiles. Improvements over a previous implementation of this method include i) multiple approximations to the velocity field from the wind equations of motion: a velocity power law, an acceleration power law, and the CAK theory $\beta$ law; ii) a variety of density profiles expressing continuous (power law, double power law exponential) and single-episode (normal, log normal, fast-rise exponential decay) scenarios as well as a composite of many expanding shells or bubbles and iii) different geometries (spherical, filled cones, and hollow cones, with cones accompanied by an optional disk). All these scenarios are consistent with conditions observed in stars, nebulae, and galaxies.
We have made this modeling approach publicly accessible through the open-source software OutLines in python, available via pip and GitHub facilities. To date, OutLines is the only line profile code with such extensive and interchangeable geometries, density profiles, and velocity fields, making it by far the most comprehensive and versatile tool available for the study of outflows, winds, and bubbles. We illustrate the usefulness of the OutLines software with application to an H ii region, a super star cluster, a starburst galaxy, and an AGN, deriving properties based on the line profile models. Furthermore, we demonstrate how the results from modeling with OutLines provide insight into the underlying physics of winds, bubbles, and outflows, obtaining the following results specifically made possible through our custom line profiles:
  • Young (<4<4 Myr) H ii regions undergoing expansion can exhibit two orders of magnitude in gas compression due to wind-triggered shocks even while optically thick radiation pressure is the main driving mechanism of the bulk motion of the gas. Thus, while shocks may dominate the radio continuum, it is radiation pressure from stars which primarily shapes the evolution of H ii regions.
  • Suppression of optically thin radiative feedback occurs in young (<2<2 Myr) super star clusters in dwarf galaxies like the Green Peas indicates that the most recent stellar populations cannot mechanically evacuate optically thin channels for Lyman continuum photons to escape. Mechanical feedback from previous episodes of star formation may be necessary to shape the gas geometry for subsequent populations to leak ionizing radiation. Thus it is ionizing feedback from the most recent stellar populations which produces the density-bounded conditions in low gas density channels through which Lyman continuum photons escape.
  • Expanding superbubbles produced by core-collapse supernovae can, in succession, accumulate up to produce an outflow as the shells relax or dissipate into a more continuous flow; however, optically thick radiation pressure is necessary for these shells to accelerate to the terminal velocity before blowout and formation of bidirectional cones extended beyond the galactic disk. The formation of outflows from systems of metal-rich shells indicates that radiative driving is a key mechanism in the baryon cycle which shapes the scaling relations of galaxies across cosmic time.
  • Accretion disk winds of a Seyfert galaxy are sufficiently fast to trigger substantial shocks in the central kiloparsec. The resulting hot fluid drives an outflow via pressure generated from the hot gas itself or from cosmic rays accelerated by the shock. The outflow consists of “cold” clouds entrained in the periphery of hollow cones and has the potential to cause inside-out quenching as it impacts the host galaxy.
This broad applicability and deep insight into the physics of outwardly moving gas demonstrates that OutLines is powerful tool which will prove to be useful across the astronomical community. Future work will involve application to massively multiplexed fiber spectroscopic surveys like WEAVE-LOFAR and 4MOST/WAVES for statistical assessments and to hydrodynamical simulations to explore the underlying physics.
Acknowledgements
We thank the referee for comments which improved this manuscript. We thank the beta testers, including K. Z. Arellano Cordova, K. Duncan, and T. Horn, whose efforts helped to optimize the use of this code. We thank K. Z. Arellano Cordova, C. Esteban, and J. Garcia-Rojas for providing access to Knot A extractions from their reduced MEGARA data. We thank K. Parker for sharing results from the recently accepted Parker2025. We thank K. Duncan for insight regarding organization of this manuscript. We also thank M. Divakara and G. Romero-Cruz for insightful discussions about outflow geometry at the National Institute of Astrophysics, Optics and Electronics (INAOE) and J. Chisholm, C. Martin, and N. Murray for insightful discussions about winds, massive stars, feedback, and HST/COS at the Aspen Center for Physics.
This work was performed in part at the Aspen Center for Physics, which is supported by a grant from the Simons Foundation (1161654, Troyer) and by National Science Foundation grant PHY-2210452.
Support for HST GO programs 12583 and 15840 was provided by NASA through a grant from the STScI, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555.
Facilities: HST, GTC, SSO
Software, repositories, and other digital resources: corner (corner), emcee (emcee), matplotlib (matplotlib), MAST, NIST/ASD (NIST_ASD), numpy (numpy), OutLines (this work, see also 10.5281/zenodo.11238265), PyNeb (2015A&A...573A..42L), S7 DR2 (S7DR2), scipy (scipy), vygrboi
AI use: No language learning model or otherwise generative artificial intelligence was used in this work.
Data Availability
The data underlying this article for NGC 5471 were provided by K. Z. Arellano Cordova, C. Esteban, and J. Garcia-Rojas by permission. Data will be shared on reasonable request to the corresponding author with permission of the above owners. The data underlying this article for J1044+4012 and Mrk 1486 are publicly available from MAST. The data underlying this article for NGC 2992 are publicly available from S7 DR2.
The OutLines code presented here is publicly available on github.com/sflury/OutLines under DOI 10.5281/zenodo.11238265 or from PyPI via pip install SpecOutLines.
Appendix A Geometry of Directed Outflows
[…]
A.1 Spherical Cap Projection
The ellipse is the projected spherical cap with axes $g$ and $h$ and is offset from the projected outflow source by some distance $S$. These quantities can be expressed in terms of the inclination and opening angle of the ouflow center such that
$S$ $=\sin{i}\cos{\theta}$ (41)
$g$ $=\sin{\theta}$ (42)
$h$ $=\cos{i}\sin{\theta}.$ (43)
In the case where $i+\theta>90^{\circ}$, the projection is no longer an ellipse but rather an ellipse plus a lune to account for the projection of the “top” of the spherical cap and an additional lune to the opposing side to account for the projection of the second cone. We visualize such a scenario in the last two panels of Figure 21.
A.2 Cone Inscription of Deprojected Velocities
The fractional arc length $\ell_{\rm w}$ of the deprojected velocity inscribed by the cone can be determined by solving the system of linear equations for a circle and the projected spherical cap area to determine where they intersect, which in turn determine the bounds on the integral(s) over the azimuthal coordinate of the solid angle. The azimuthal “ring” of gas with velocity $w$ is defined by the sine of the angle between the observed and deprojected velocities $u$ and $w$, respectively, such that its radius is
$r_{\rm w}=\sqrt{1-(u/w)^{2}}.$ (44)
Solving the system of equations of the projected cap and the deprojected velocity ring gives intersection points of ($p$, $q$) such that
$p=\pm\sqrt{r_{\rm b}^{2}-q^{2}}$ (45)
and, from substituting $p$ and employing the quadratic equation,
$q=\frac{-B\pm\sqrt{B^{2}-4AC}}{2A}$ (46)
where the coefficents are
$A$ $=\left(\frac{g}{h}\right)^{2}-1$ (47)
$B$ $=-2S\left(\frac{g}{h}\right)^{2}$ (48)
$C$ $=r_{\rm b}^{2}+\left(\frac{gS}{h}\right)^{2}-g^{2}.$ (49)
To reduce computation time, $A$ and $B$ are calculated outside of the integral as only $C$ depends on the de-projected velocity $w$. The intersection points ($p$, $q$) given by $A$, $B$, and $C$ yield the arc of the enclosed deprojected velocity such that
$\pi\ell_{\rm w}=\begin{cases}\arccos{\frac{q}{r_{\rm b}}}&q>0\\ \pi-\arccos{\frac{q}{r_{\rm b}}}&q<0\end{cases}$ (50)
which notably avoids the additional step of calculating $p$. If multiple real solutions for ($p$, $q$) exist due to the lune from the projection of the second cone into the plane, then OutLines simply sums the two arcs.
We show example projections in Figure 21 for several values of $u/w$, $i$, and $\theta$. As a tool, we also include in OutLines a module to produce such visualizations for any given combination of $i$ and $\theta$. Calling OutLines.PlotConeProjection(i,thetaO) will draw the geometry and projection as shown in Figure 21 for any inclination and opening angle.
Figure 21: Directed outflows (left of each panel) projected onto the plane of the sky (right of each panel) for various combinations of $i$ and $\theta_{o}$ (orange). Circles indicate deprojected velocities of u/w=u/w=0.4 (green), 0.6 (blue), and 0.8 (purple). Opaque segments indicate where a deprojected velocity is inscribed by the projected spherical cap. Arrow in the left of each panel indicates direction of the projections onto the plane of the sky (i.e., the line of sight).
A.3 Cavity Effects
As discussed in §2.4, some outflows exhibit a bidirectional cones with cavities, resulting in a hollow cone geometry. The fractional arc length $\ell_{\rm cavity}$ lost due to the cavity can be calculated in the same manner as $\ell_{\rm w}$ and subtracted, which we show in Figure 22. Calling OutLines.PlotConeProjection(i,thetaO,thetaC) will illustrate projections similar to those of the filled cones but for the hollow cone geometry.
Figure 22: Same as Figure 21 but for a directed outflow with an inner cavity.
A.4 Obstruction by a Disk
In §2.4, we present the possible inclusion of a disk which encircles the source of the outflow and is inclined orthogonally to the cone axis. This disk obstructs the posterior cone but may affect either the red-shifted or both wings of the line profile. OutLines projects the disk onto the line of sight and then determines the length of each deprojected velocity which is still observed after accounting for the disk effect. Since the disk is centered at the source of the outflow and assumed to be orthogonal to the outflow cone, the projected disk is simply an ellipse with semi-major and semi-minor axes of ($w[x_{d}|\beta],w[x_{d}|\beta]\sin i$). Thus, the intersection points ($p_{d}$$q_{d}$) between the deprojected velocity and the disk can be written as
$p_{d}=\pm\sqrt{r_{\rm b}^{2}-q_{d}^{2}}$ (51)
and
$q_{d}=\pm\sqrt{\frac{r_{\rm b}^{2}-w^{2}[x_{d}|\beta]}{\csc^{2}i-1}}.$ (52)
Rather than compute $\ell_{\rm w}$ and $\ell_{\rm disk}$ separately, which would needlessly incur a computational overhead, we can instead simply compare $q$ and $q_{d}$ to determine the final $\ell_{\rm w}$. For the posterior cone, $\ell_{\rm w}$ is given by
$\pi\ell_{\rm w}=\begin{cases}\arccos{\frac{q}{r_{\rm b}}}&q>q_{d}\\ \arccos{\frac{q_{d}}{r_{\rm b}}}&-q_{d}<q<q_{d}\\ 2\arccos{\frac{q_{d}}{r_{\rm b}}}-\arccos{\frac{q}{r_{\rm b}}}&q<-q_{d}\\ \end{cases}.$ (53)
If $q$ has multiple solutions due to the projected lune from the anterior cone, OutLines will add to $\ell_{\rm w}$ the first case evaluated using the second solution for $q$ with the imposition that $q>q_{d}$ since the anterior cone is not obstructed by the disk. Note that, in the case of $i+\theta>90^{\circ}$, the red component will never be fully obstructed. For the anterior cone, $\ell_{\rm w}$ is given by
[…]
where the $q_{1}$ criterion only applies in the case of a lune projected from the posterior cone, which can be obstructed by the disk. Note that, in the case of $i+\theta>90^{\circ}$, the blue component can be partially obstructed, primarily affecting the line core. Under this formalism, $\ell_{\rm cavity}$ can still be subtracted to determine the hollow cone geometry as both $\ell_{\rm w}$ and $\ell_{\rm cavity}$ will account for any disk effects on the deprojected velocity apriori.
Appendix B Non-zero Launch Velocities
While the OutLines default assumes that gas is launched from rest, we have allowed for the possibility that the gas starts from some launch velocity $v_{0}$, which we include as an optional free parameter. The velocity limits are accordingly updated so that the minimum velocity $w_{\ell}$ which contributes to an observed velocity $|u|$ is never below $v_{0}/v_{\infty}$. The velocity fields in Table 4 and Equations 15-17 are modified by a factor of $1-v_{0}/v_{\infty}$ with the addition of a constant $v_{0}/v_{\infty}$ such that the relative velocity not from rest $w_{\rm nfr}$ is
$w_{\rm nfr}=w[x|\beta]\left(1-\frac{v_{0}}{v_{\infty}}\right)+\frac{v_{0}}{v_{\infty}}.$ (55)
Emergent profiles are similar to those presented in TenorioTagle1996, which included distributions of non-zero shell ages, and to those presented in 2015ApJ...801...43S, which account for non-zero launch velocities. We show example line profiles with varying launch velocities in Figure 23. Future work will revisit this possibility in greater detail.
Figure 23: Effect of varying the launch velocity from rest ($v_{0}=0$) to 30% of the terminal velocity on nebular emission (left) and absorption (right) line profiles, here normalized to the case of launching from rest for visualization.
Table 2: Left: the Profile object classes in OutLines with which users can generate model profiles for different spectral line types. Right: keyword settings for the Profile classes to customize models.
class use atomic data keywords options default
Nebular nebular emission lines $\lambda_{0}$ Geometry
VelocityField
DensityProfile
AddStatic
FromRest
Disk
Aperture
Table 3
Table 4
Table 5
boolean
boolean
boolean
boolean
Spherical
BetaCAK
PowerLaw
False
True
False
False
Resonant resonant emission lines $\lambda_{0},\ f,\ p_{r}$
Fluorescent fluorescent emission lines $\lambda_{0},\ f,\ p_{r},\ p_{f}$
Absorption resonant absorption lines $\lambda_{0},\ f$
Appendix C Implementation of OutLines
C.1 Installation
Installing OutLines can be done in the terminal command line using pip as shown below.
1pip3 install SpecOutLines
Alternatively, the GitHub repository can be cloned as below
1git clone https://github.com/sflury/OutLines.git
Dependencies include python >=3.12.9, which may require updating of python globally or creating, say, an anaconda environment. Other dependencies, including numpy, scipy, and matplotlib will be automatically managed by pip, including within conda environments.
C.2 The Profile Classes
The primary front-end user interaction with OutLines is through the various Profile object classes, which we list in Table 2. Calling one of these classes instantiates an object containing the line profile model, parameter guesses, and parameter bounds. When calling a class, users need only provide the relevant atomic data and any optional keyword arguments specifying the velocity field, geometry, and density profile to customize the model (see Table 2). The terms and related keyword arguments for the profile model are listed in Table 3 for the geometry, Table 4 for the velocity field, and Table 5 for the density profile, with the OutLines defaults indicated by a .
Table 3: Geometry keyword options and related required arguments for OutLines. indicates the OutLines default. indicates geometry only applicable to emission lines.
Geometry terms assumptions
‘Spherical’  none $i=0$, $\theta_{c}=0^{\circ}$, $\theta_{o}=90^{\circ}$
‘FilledCones’ $i$, $\theta_{o}$ $\theta_{c}=0^{\circ}$
‘HollowConesFixedCavity’ $i$, $\theta_{o}$ $\theta_{c}=\theta_{o}-10^{\circ}$
‘HollowCones’ $i$, $\theta_{c}$, $\theta_{o}$ none
‘OpenCones’
Table 4: Velocity field keyword options for OutLines, each paramaterized by an exponent $\beta$. indicates the OutLines default.
VelocityField equation references
‘BetaCAK’  15 1975ApJ...195..157C; Castor1979
‘AccPlaw’ 16 2010ApJ...717..289S
‘VelPlaw’ 17 FaucherGiguere2012
Table 5: Density profile keyword options and related required arguments for OutLines. indicates the OutLines default.
DensityProfile terms equation
‘PowerLaw’  $\alpha$ 18
‘Exponential’ $\gamma$ 19
‘Normal’ $x_{1}$, $\sigma_{x}$ 21
‘LogNormal’ $\log x_{1}$, $\log\sigma_{x}$ 22
‘Shell’ $x_{1}$, $\sigma_{x}$ 24
‘FRED’ $x_{1}$, $\tau_{1}$, $\tau_{2}$ 23
‘DoublePowerLaw’ $\alpha_{1}$, $\alpha_{2}$, $x_{1}$ 20
‘Pulses’ $x_{1}$, $\sigma_{x}$, $x_{k}$ 25
‘DampedPulses’ $\gamma$, $x_{1}$, $\sigma_{x}$, $x_{k}$ 26
Users will primarily use OutLines by instantiating one of the several Profile classes: Nebular, Resonant, Fluorescent, and Absorption. The object-bound methods of the Profile classes can be used to obtain or update parameter values and to generate line profiles. We provide examples of using these classes to generate line profiles in Appendix D. To obtain line profiles computed by OutLines, users can call the get_profile() bound method for either class, which requires only an array of wavelengths and computes a line profile for the parameters stored in the object. Alternatively, the line profile function is directly accessible through the Profile attribute, which requires a wavelength array just as get_profile() but also takes the full set of parameters. For purposes of modeling observations, users may prefer to use the Profile function where the input parameters may be readily varied without updating the corresponding attribute in the object.
The get_params() and get_bounds() methods will pass the relevant values from the object to the user. The print_settings(), print_params(), and print_bounds() will print the selected profile settings (velocity field, geometry, density profile, and optional static line component), the corresponding current parameters, and the associated boundaries. The update_params() and update_bounds() methods will allow users to update any of the profile parameters and their related bounds. All Profile classes can each support multiple lines, something we encourage users to consider.
C.3 The Properties Class
After modeling a wind or outflow from observed spectral lines, users will want to derive important diagnostic properties to determine what drives the wind, how, and to what extent the outflow impacts its environment. Because the relevant information is already contained in the Profile classes, Properties is designed to inherit one of the Profile classes and will automatically compute relevant properties defined in 2.5, with the caveat that $n_{0}$ and $R_{0}$ remain unknown.
Users can access these results directly from the object class once instantiated, either through the stored dictionary or by printing to the command line via the bound method print_propts(). As with the Emisison and Absorption classes, Properties also has the bound method update_params(), which behaves similar to the bound method in the Profile classes by allowing users to change parameters associated with the line profile.
Table 6: Outflow properties computed by the OutLines.Properties object class. Property names listed here correspond to the dictionary keys in props attribute and the terminal output by print_props method.
name equation property
x.out $x_{out}$ characteristic radius
v.out 32 $v_{out}$ characteristic radius
Rcal 33 $\mathcal{R}$ relative column density
Mdot 37 $\dot{M}/n_{0}R_{0}^{2}$ mass outflow rate
pdot 38 $\dot{p}/n_{0}R_{0}^{2}$ momentum flux
Edot 39 $\dot{E}/n_{0}R_{0}^{2}$ mechanical luminosity
C.4 Visualization
To assist with interpretation of outflow properties and geometry, OutLines contains several functions to generate figures like the ones presented here. To produce isocontours as in Figure 6, we provide the PlotIsoContours function. To produce the edge-on geometric cartoon as in Figure 10, we provide the PlotGeometry function. To produce the velocity projections for filled and hollow cones as in Figures 21 and 22, we provide the PlotConeProjection function.
C.5 Numerical Methods
Some equations require root-finding where analytic solutions are not possible, including the “source-effect” and aperture velocity limits (Equations 28 and 30, respectively) and the characteristic outflow radius $x_{out}$ where the momentum density is maximized. As in Flury2023, we employ Brent’s method implemented by scipy.optimize to solve for the root in each case.
The initial profile model of Flury2023 solved the integral for the emission line profile (Equation 10) using the adaptive Romberg method, which included Richardson extrapolation under the Bulirsch-Stoer method. As implemented by the now-deprecated scipy.integrate.romberg, this method can be costly and in later scipy versions has loss of accuracy in non-spherical geometries due to poor sampling of the integrand. For improvements to both speed and accuracy, we now adopt Gaussian quadrature with a fixed number of nodes (as implemented by scipy.integrate.fixed_quad) to solve each of the integrals for $\phi_{\lambda}$. We find at least 16 nodes necessary to obtain reliable profile solutions in many cases but conservatively assume 256 nodes for optimal accuracy across the profile parameter spaces.
For a single core on a conventional laptop computer, benchmarking indicates that OutLines can calculate an emission line profile at $R=5,000$ in $\approx$20 ms for any assumed geometry. Given reasonable guess parameters, iterative non-linear least squares fitting routines like the trust-reflective algorithm (see scipy.optimize.least_squares) perform rapidly, typically converging on a solution within 10 to 30 s depending on the data and the assumed geometry. For 25 walkers, a typical MCMC sampler (here benchmarked using emcee, emcee) takes anywhere from $0.1$ to $2$ s per step to run on a conventional laptop computer when fitting an OutLines model to an observed line profile, even when accounting for complex geometries and LSF effects. The addition of convolution with a line spread function does increase the computation time; however, the extent to which this occurs depends on the resolution, size of the data set, and the implementation of the convolution.
Appendix D Example Line Profiles with OutLines
Below, we provide example usages of the Absorption, Nebular, and Fluorescent line profile object classes in OutLines. We demonstrate how to select different options for the line profile model, update model parameters and bounds associated with the line profile, display documentation, print various model settings and parameters, and extract computed line profiles for visualization. We also demonstrate how to compute outflow properties from the Properties class and print the results to the command line.
D.1 C ii $\lambda 1334$
Example instantiations of Absorption, Resonant, and Fluorescent objects from the Profile class using atomic data from Morton2003. Here, the object classes are instantiated, documentation, setting, and parameters are printed to the terminal, and the column density and terminal velocity are updated.
1# import the OutLines package
2import OutLines as OL
3# instantiate the absorption line
4# Profile object class
5modelAbs = OL.Absorption(1334.519,0.128)
6# print documentation of the model object
7modelAbs.docs()
8# print model settings
9modelAbs.print_settings()
10# print the current parameters
11model.print_params()
12# instantiate the resonant and fluorescent
13# Profile object classes
14modelRes = OL.Resonant(1334.532,0.128,0.456)
15modelFlu = OL.Fluorescent(1335.663,0.115,0.456,0.544)
16# updates for all three objects
17for model in [modelAbs,modelRes,modelFlu]:
18 # change the log column density to 1e15 cm^-2
19 model.update_params('LogColumnOutflow1',15)
20 # change the velocity to 500 km s^-1
21 model.update_params('TerminalVelocity', 500.00)
D.2 [O iii] $\lambda 4959,5007$
Example insantiation of the Nebular object from the Profile class in OutLines using atomic data from NIST (NIST_ASD) and Storey2000. Here, we illustrate how to customize the model with keyword arguments passed to the object class on instantiation, updating parameters, printing setting and parameters, computing and printing outflow properties, and visualizing a line profile as in Figure 24.
1# import packages
2import OutLines as OL
3from numpy import linspace
4import matplotlib.pyplot as plt
5# instantiate the nebular line
6# profile object
7kwargs = dict(Geometry='HollowCones',AddStatic=True,\ VelocityField='AccPlaw',Disk=True)
8model = OL.Nebular([4958.911,5006.843],**kwargs)
9# update some parameters
10model.update_params(['TerminalVelocity'],[500,30,15,45])
11model.update_params(['OpeningAngle','CavityAngle',\ 'Inclination'],[30,15,45])
12model.update_params(['FluxOutflow1','FluxStatic1'],\ [1/2.98,1/2.98]) # Storey & Zeppen 2000
13# print model settings and current parameters
14model.print_settings()
15model.print_params()
16# compute outflow properties
17props = OL.Properties(model)
18# print outflow properties
19props.print_props()
20# plot the profile
21wave = linspace(4950,5015,651)
22plt.plot(wave,model.get_outflow(wave),label='outflow')
23plt.plot(wave,model.get_static(wave),label='static')
24plt.plot(wave,model.get_profile(wave)+0.1,label='total')
25plt.show()
Figure 24: Example [O iii] $\lambda\lambda 4959,5007$ line doublet profile produced using OutLines as shown in Appendix D.2.
Appendix E Example Modeling
Below, we provide examples for modeling absorption and nebular emission line features using line profiles computed by OutLines. We indicate suggested methods for obtaining best-guess parameters and demonstrate how to use OutLines in tandem with MCMC samplers using the initial value generator and logarithmic posterior probability contained in the object instance of the profile class.
E.1 Si ii $\lambda 1190,1193$ Absorption Doublet
1# import the OutLines package
2import OutLines as OL
3# MC sampler for posterior
4import emcee
5# wavelength, continuum-subtracted flux, and uncertainty
6wave,flux,flux_error = data
7# set up OutLines model using NIST vacuum data
8kwargs = dict(DensityProfile='DampedPulses',\ AddStatic=True)
9model = OL.Absorption([1190.416,1193.290],[0.277,0.575],\ **kwargs)
10# initialize walkers and sampler and run the sampler
11init,nwalk = model.init_params()
12# run the MCMC sampler
13sampler = emcee.EnsembleSampler(nwalk,model.npar,\ model.log_probability,args=(wave,flux,flux_error))
14sampler.run_mcmc(init,1200, progress=True)
15samples = sampler.get_chain(discard=200,flat=True)
16# extract parameters from posterior sample
17for i,param in enumerate(model.get_param_names()):
18 value = nanquantile(samples[:,i],0.5)
19 model.update_params(param,value)
20# print line profile parameters to terminal
21model.print_params()
E.2 [S ii] $\lambda 6716,6731$ Emission Doublet
1# import the OutLines package
2import OutLines as OL
3# scipy nonlinear least squares for initial guess
4from scipy.optimize import curve_fit
5# MC sampler for posterior
6import emcee
7# wavelength, continuum-subtracted flux, and uncertainty
8wave,flux,flux_error = data
9# set up OutLines model using NIST air wavelengths
10mod_kwargs = dict(AddStatic=True)
11model = OL.Nebular([6716.440,6730.815],**mod_kwargs)
12# simple trust-reflective fit for initial guess
13tr_kwargs = dict(p0=model.get_params(),\ bounds=model.get_bounds(),sigma=flux_error)
14p0,c0 = curve_fit(model.Profile,wave,flux,**tr_kwargs)
15model.update_params(model.get_param_names(),p0)
16# initialize walkers and sampler and run the sampler
17init,nwalk = model.init_params()
18# run the MCMC sampler
19sampler = emcee.EnsembleSampler(nwalk,model.npar,\ model.log_probability,args=(wave,flux,flux_error))
20sampler.run_mcmc(init,1200, progress=True)
21samples = sampler.get_chain(discard=200,flat=True)
22# extract parameters from posterior sample
23for i,param in enumerate(model.get_param_names()):
24 value = nanquantile(samples[:,i],0.5)
25 model.update_params(param,value)
26# print line profile parameters to terminal
27model.print_params()
28# obtain outflow properties
29props = OL.Properties(model)
30# print outflow properties to terminal
31props.print_props()
E.3 [O iii] $\lambda 4959,5007$ Emission Doublet for the Seyfert 2 in NGC 2992
1# import the OutLines package
2import OutLines as OL
3# MethodType for changing prior
4from types import MethodType
5# scipy nonlinear least squares for initial guess
6from scipy.optimize import curve_fit
7# MC sampler for posterior
8import emcee
9# log prior -- uniform bounded
10def log_prior(self,theta):
11 tlower,tupper = self.get_bounds()
12 tbound = list(map(self.__fun_bound__,theta,tlower,tupper))
13 # enforce theta_o - theta_c >= 5 degrees
14 # suggest atomic data to within 5%
15 # suggest that the best-guess angles
16 # are accurate within 10 degrees
17 if all(tbound) and theta[8]-theta[9] > 0.087:
18 return -0.5*((2.99e5*theta[1]-1100)/300)**2 \
19 -0.5*((2.99e5*theta[0]- 75)/50)**2 \
20 -0.5*((theta[7]-1.25)/0.2)**2 \
21 -0.5*((theta[8]-0.95)/0.2)**2 \
22 -0.5*((theta[9]-0.40)/0.2)**2 \
23 -0.5*((theta[4]/theta[3]-2.98)/0.05)**2 \
24 -0.5*((theta[6]/theta[5]-2.98)/0.05)**2
25 else:
26 return -inf
27# instantiate the nebular line
28# profile object using
29# rest wavelengths from NIST
30kwargs = dict(Geometry='HollowCones',AddStatic=True,\ VelocityField='AccPlaw',Disk=True)
31model = OL.Nebular([4958.911,5006.843],**kwargs)
32# update some parameters
33model.update_params(['TerminalVelocity'],[500,30,15,45])
34model.update_params(['OpeningAngle','CavityAngle',\ 'Inclination'],[60,30,60])
35model.update_params(['FluxOutflow1','FluxStatic1'],\ [1/2.98,1/2.98]) # Storey & Zeppen 2000
36# simple trust-reflective fit for initial guess
37tr_kwargs = dict(p0=model.get_params(),\ bounds=model.get_bounds(),sigma=flux_error)
38p0,c0 = curve_fit(model.Profile,wave,flux,**tr_kwargs)
39model.update_params(model.get_param_names(),p0)
40# initialize walkers and sampler and run the sampler
41init,nwalk = model.init_params()
42# use the custom prior
43model.log_prior = MethodType(log_prior,model)
44# run the MCMC sampler
45sampler = emcee.EnsembleSampler(nwalk,model.npar,\ model.log_probability,args=(wave,flux,flux_error))
46sampler.run_mcmc(init,1200, progress=True)
47samples = sampler.get_chain(discard=200,flat=True)
48# extract parameters from posterior sample
49for i,param in enumerate(model.get_param_names()):
50 value = nanquantile(samples[:,i],0.5)
51 model.update_params(param,value)
52# print line profile parameters to terminal
53model.print_params()
54# obtain outflow properties
55props = OL.Properties(model)
56# print outflow properties to terminal
57props.print_props()
Appendix F Posterior Sampling
All posterior samples for each of the OutLines modeling cases were obtained using emcee (emcee) and visualized with figures produced by corner (corner) and subsequently adapted. Results are shown for NGC 5471 Knot A in Figure 25, the super star cluster of Green Pea J1044+0353 in Figure 26, Mrk 1486 in 27, and the Seyfert 2 nucleus of NGC 2992 in 28.
[…]

### v2

@@
 Abstract
-Common methods for studying the kinematics and geometry of outflowing gas rely on modeling emission and absorption lines in integrated spectra using methods that are not physically motivated, including empirical quantiles or fitting multiple Gaussian or Voigt profiles.
-Such methods are not always consistent with the interpretation of these features and, as a result, miss key underlying physics and can even lead to inaccurate interpretations of observations.
-To address this problem, we present the publicly available python code OutLines, which provides astrophysical models of spectral emission and absorption line profiles produced by outflows in a variety of environments.
-The OutLines code accounts for differences in parameterization of the velocity field and density profile while allowing for different outflow geometries, making OutLines versatile and useful for a wide variety of astrophysical phenomena.
+Studies of kinematics and geometry of outflowing gas rely on modeling features in integrated spectra using empirical quantiles or fitting multiple Gaussian or Voigt profiles.
+Such methods can miss key underlying physics and even lead to spurious interpretations of observations.
+To address this problem, we present the public python code OutLines, which predicts spectral emission and absorption line profiles produced by winds, bubbles, and outflows.
+By design, OutLines is phenomenologically and scale agnostic to bridge the gap between observations and simulations across a broad swath of astrophysics.
+The OutLines code accounts for differences in velocity field, density profile, and outflow geometries, making OutLines versatile for a wide variety of astrophysical phenomena.
 We demonstrate the wide applicability of OutLines by using the code to model line profiles in an H ii region knot, super star clusters, a starburst galaxy, and an AGN.
-In each of these contexts, we illustrate how OutLines can illuminate key underlying physics in ways that improve our scientific understanding and address important open questions in astronomy, including the key mechanisms in the baryon cycle, the evolution of H ii regions and galaxies, and even Lyman continuum escape.
-OutLines will be a critical resource as massively multiplexed spectroscopic surveys like WEAVE-LOFAR and 4MOST/WAVES come online, providing the means to probe feedback kinematics with deeper, higher resolution spectroscopy for unprecedented large samples of galaxies.
+In each of these contexts, we illustrate how OutLines can illuminate key underlying physics in ways that improve scientific understanding and address important open questions in astronomy, including the key mechanisms in the baryon cycle, the evolution of H ii regions and galaxies, and even Lyman continuum escape.
+OutLines will be a critical resource as massively multiplexed spectroscopic surveys with facilities like WEAVE and 4MOST come online, providing the means to probe feedback with deeper, higher resolution spectroscopy for unprecedented large samples of objects.
 keywords:
 ISM: jets and outflows – ISM: bubbles – ISM: kinematics and dynamics – galaxies: kinematics and dynamics
-pubyear: 2025pagerange: OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows28
+pubyear: 2026pagerange: OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows29
 1 Introduction
 Outflows, winds, and bubbles are ubiquitous astrophysical phenomena occurring on a dynamic range of spatial scales and in a variety of contexts, from young stellar objects (YSOs) to nebulae, active galactic nuclei (AGN), galaxies, and even galaxy clusters.
-This feedback has the capacity to shape many key astrophysical systems, from YSO winds affecting the formation of exoplanets (e.g., Suzuki2009; Manara2023) to massive stars redistributing the leftover material from their birth clouds (e.g., Dopita2005; Menon2024) to AGN and starbursts driving the baryon cycle of galaxies (e.g., 2005ApJ...618..569M; Tumlinson2017).
-Many physical mechanisms can accelerate, or drive, winds and bubbles, from cosmic rays (e.g., Ipavich1975) to hot expanding gas (e.g., 1985Natur.317...44C) to radiation pressure on dust (e.g., Poynting1904; LyndenBell1971) and ions (so-called line driving, e.g., Lucy1970), each with direct implications for how feedback shapes the surrounding environment (see, e.g., 2020A&ARv..28....2V; Thompson2024, and references therein).
-A common method of detecting and studying feedback is through broad, often asymmetric spectroscopic features (e.g., Castor1970; 1975ApJ...195..157C; 1974ApJ...193..651N; 1980A&A....81..172P; 1981MNRAS.195..787P; 1981ApJ...247..403H; Chu1986), perhaps most famously with P Cygni wind lines in stellar atmospheres (e.g., 1986A&A...164...86P; 1987ApJ...314..726L; 2001A&A...369..574V).
+This feedback has the capacity to shape many key astrophysical systems, from YSO winds affecting the formation of exoplanets (e.g., Suzuki and Inutsuka, 2009; Manara et al., 2023) to massive stars redistributing the leftover material from their birth clouds (e.g., Dopita et al., 2005; Menon et al., 2024) to AGN and starbursts driving the baryon cycle of galaxies (e.g., Murray et al., 2005; Tumlinson et al., 2017).
+Many physical mechanisms can accelerate, or drive, winds and bubbles, from cosmic rays (e.g., Ipavich, 1975) to hot expanding gas (e.g., Chevalier and Clegg, 1985) to radiation pressure on dust (e.g., Poynting, 1904; Lynden-Bell and Rees, 1971) and ions (so-called line driving, e.g., Lucy and Solomon, 1970), each with direct implications for how feedback shapes the surrounding environment (see, e.g., Veilleux et al., 2020; Thompson and Heckman, 2024, and references therein).
+A common method of detecting and studying feedback is through broad, often asymmetric spectroscopic features (e.g., Castor, 1970; Castor et al., 1975a; Noerdlinger and Rybicki, 1974; Pelat and Alloin, 1980; Pelat et al., 1981; Heckman et al., 1981; Chu and Kennicutt, 1986), perhaps most famously with P Cygni wind lines in stellar atmospheres (e.g., Pauldrach et al., 1986; Lamers et al., 1987; Vink et al., 2001).
 Physically motivated models for these line profiles is thus of critical importance to understanding how energetic phenomena shape their environment.
-While the physics and formalism for spectroscopic features arising from winds, bubbles, and outflows is not new (e.g., Beals1931; 1960mes..book.....S; Castor1970; 1975ApJ...195..157C; 1986A&A...164...86P; 1987ApJ...314..726L), their application is relatively sparse outside the phenomena of stellar atmospheres (e.g., 2001A&A...369..574V; 2021MNRAS.504.2051V; Sander2017; Puls2020).
-Instead, fitting multiple Gaussian or Voigt profiles to broad, asymmetric lines has become the standard (e.g., 1980A&A....81..172P; 1981MNRAS.195..787P) for a wide variety of objects, including H ii regions (e.g., Bresolin2020; Yarovova2025), AGN (e.g., Liu2020), and galactic winds (e.g., Hogarth2020; Peng2023; Peng2025), despite early evidence that physics-based profiles are not only more appropriate but also demonstrably different in shape while still providing good descriptions of observations (e.g., Zheng1990; TenorioTagle1996).
-Some studies have skirted the issue by adopting a more empirical approach using velocity quantiles (e.g., 1981ApJ...247..403H; 1985MNRAS.213....1W; 1991ApJS...75..383V) or by fitting the line profile wings with a log-linear (e.g., 2021ApJ...920L..46K; Komarova2025) or exponential (e.g., Rusakov2025) functions to characterize the steepness of the non-Gaussian flux distribution.
+While the physics and formalism for spectroscopic features arising from winds, bubbles, and outflows is not new (e.g., Beals, 1931; Sobolev, 1960; Castor, 1970; Castor et al., 1975a; Pauldrach et al., 1986; Lamers et al., 1987), their application is relatively sparse outside the phenomena of stellar atmospheres (e.g., Vink et al., 2001; Vink and Sander, 2021; Sander et al., 2017; Puls et al., 2020).
+Instead, fitting multiple Gaussian or Voigt profiles to broad, asymmetric lines has become the standard (e.g., Pelat and Alloin, 1980; Pelat et al., 1981) for a wide variety of objects, including H ii regions (e.g., Bresolin et al., 2020; Yarovova et al., 2025), AGN (e.g., Liu et al., 2020), and galactic winds (e.g., Hogarth et al., 2020; Peng et al., 2023, 2025), despite early evidence that physics-based profiles are not only more appropriate but also demonstrably different in shape while still providing good descriptions of observations (e.g., Zheng et al., 1990; Tenorio-Tagle et al., 1996).
+Some studies have skirted the issue by adopting a more empirical approach using velocity quantiles (e.g., Heckman et al., 1981; Whittle, 1985; Veilleux, 1991) or by fitting the line profile wings with a log-linear (e.g., Komarova et al., 2021, 2025) or exponential (e.g., Rusakov et al., 2025) functions to characterize the steepness of the non-Gaussian flux distribution.
 The benefit of such methods is that they require no physical model while still providing insight into kinematics.
 However, bridging the gap between these approaches and the underlying physics is not always straightforward.
-Some effort towards a physically motivated approach has emerged in recent years, including detailed treatment of radiation pressure and conservation effects (e.g., 2017MNRAS.471.4061K), turbulent mixing layers (e.g., 2009A&A...500..817B), and relativistic accretion disk winds (e.g., DISKWIND1; DISKWIND2; XRADE) along with more generalized models (e.g., 2010ApJ...717..289S; 2011ApJ...734...24P; Martin2013; 2015ApJ...801...43S; 2016MNRAS.463..541C; Luminari2018; Flury2023).
+Some effort towards a physically motivated approach has emerged in recent years, including detailed treatment of radiation pressure and conservation effects (e.g., Krumholz et al., 2017), turbulent mixing layers (e.g., Binette et al., 2009), and relativistic accretion disk winds (e.g., Sim et al., 2008, 2010; Matzeu et al., 2022) along with more generalized models (e.g., Steidel et al., 2010; Prochaska et al., 2011; Martin et al., 2013; Scarlata and Panagia, 2015; Chisholm et al., 2016; Luminari et al., 2018; Flury et al., 2023).
 Despite these exciting advancements, wide-spread adoption of these methods has stalled.
-A key to furthering the use of physically-motivated line profiles is versatile, well-documented, and user-friendly software tools which make such models readily accessible to the community to generate spectral line profiles linking observations to outflow physics.
-In this paper, we present OutLines, a python-based code for generating model emission, absorption, and P Cygni line profiles arising from astrophysical winds, bubbles, and outflows.
+A key to furthering the use of physically-motivated line profiles is versatile, well-documented, and user-friendly software tools which make such models readily accessible to the community to generate spectral line profiles linking observations to outflow physics without forward-modeling simulations.
+In this paper, we present OutLines, a python-based code for generating model emission, absorption, and P Cygni line profiles arising from astrophysical winds, bubbles, and outflows on arbitrary spatial scales and independent of the phenomenon.
 The input geometry and wind physics can be readily customized, and functionality for modeling is provided.
@@
 We demonstrate the versatile applicability and scientific usefulness of OutLines in §4 by modeling broad spectral lines in an H ii region knot, a super star cluster (SSC), a starbursting galaxy, and an active galactic nucleus (AGN), with profiles generated by OutLines.
-Finally, we summarize OutLines and its potential for high scientific impact in §5, providing worked examples in Appendix D and E.
+In the final case, we demonstrate that OutLines can successfully recover wind geometry through comparison with observations.
+Finally, we summarize OutLines and its potential for high scientific impact in §5, providing details of code setup and installation in Appendix B and worked examples in Appendices C and D.
 2 Physical Model
-Within the framework of the Sobolev approximation (1944SvA.....21..143S; 1957SvA.....1..678S; 1960mes..book.....S), small scale Doppler motions of gas within clouds can be ignored in the context of the large-scale motion of the clouds themselves.
-The Sobolev approach has a long-standing precedent in the study of winds (e.g., Castor1970; 1974ApJ...193..651N; 1975ApJ...202..465M; 1987ApJ...314..726L), with successful application to galactic winds and outflows over the past 15 years (e.g., 2010ApJ...717..289S; 2011ApJ...734...24P; Martin2013; 2015ApJ...801...43S; 2016MNRAS.463..541C; Flury2023).
-Under this assumption, the key characteristics of winds, outflows, and bubbles which dictate the line profile are the velocity field, density profile, and geometry of the clouds (e.g., Zheng1990).
-Below, we elaborate how the observed projected velocity $u$ in a spectral line, given by
-$u=\frac{\lambda-\lambda_{0}}{\lambda_{0}}\frac{c}{v_{\infty}}$ (1)
-for a line with rest wavelength $\lambda_{0}$ and normalized to the wind terminal velocity $v_{\infty}$, relates to the intrinsic wind velocity $w$ given by some radially-symmetric velocity field
+Within the framework of the Sobolev approximation (Sobolev, 1944, 1957, 1960), small scale Doppler motions of gas within clouds can be ignored in the context of the large-scale motion of the clouds themselves.
+The Sobolev approach has a long-standing precedent in the study of winds (e.g., Castor, 1970; Noerdlinger and Rybicki, 1974; Mihalas et al., 1975; Lamers et al., 1987), with successful application to galactic winds and outflows over the past 15 years (e.g., Steidel et al., 2010; Prochaska et al., 2011; Martin et al., 2013; Scarlata and Panagia, 2015; Chisholm et al., 2016; Flury et al., 2023).
+Under this assumption, the key characteristics of winds, outflows, and bubbles which dictate the line profile are the velocity field, density profile, and geometry of the clouds (e.g., Zheng et al., 1990).
+Below, we elaborate how the observed projected velocity $u$ in a spectral line, given by the longitudinal relativistic Doppler shift (Einstein, 1905, also Sher 1968 their Eqn 4) such that
+$u=\frac{c}{v_{\infty}}\frac{(\lambda/\lambda_{0})^{2}-1}{(\lambda/\lambda_{0})^{2}+1}$ (1)
+for a line with rest wavelength $\lambda_{0}$ and normalized to the wind terminal velocity $v_{\infty}$ and reducing to the classical Doppler shift $u=c(\lambda/\lambda_{0}-1)/v_{\infty}$ when $v_{\infty}\ll c$, relates to the intrinsic wind velocity $w$ given by some radially-symmetric velocity field
 $w=\frac{v_{\rm wind}[x,\beta]}{v_{\infty}}$ (2)
@@
 […]
-Under the Sobolev approximation, instead one must evaluate the relative column density at each velocity $v$ in the wind or bubble that is projected onto the line of sight velocity $u$.
-The optical depth profile for the outflow component is given by the Sobolev approximation such that
-$\phi_{\lambda}=K^{-1}\int n[x]\frac{{\rm d}x[w]}{{\rm d}w}{\rm d}\Omega$ (5)
-over some differential solid angle ${\rm d}\Omega$ given by the geometry (see Rybicki1983, their Equation 2.14) and where $K$ is a normalization constant ensuring $\int_{\lambda}\phi_{\lambda}{\rm d}\lambda=1$ (Rybicki1979, their Equation 1.65).
-The differential term $\frac{{\rm d}x}{{\rm d}w}$ arises from integrating over all column densities associated with the observed velocity $u$ (Castor1970, their Equation 6, 2016MNRAS.463..541C, their Equation 4, assuming no stimulated emission).
+Under the Sobolev approximation, instead one must evaluate the relative column density at each velocity $w$ in the wind or bubble that is projected onto the line of sight velocity $u$.
+The optical depth profile for the outflow component given by the Sobolev approximation is thus
+$\phi_{\lambda}=K^{-1}\int n[x]\frac{{\rm d}x[w]}{{\rm d}w}{\rm d}\Omega^{\prime}$ (5)
+over some observed differential solid angle ${\rm d}\Omega$ given by the geometry (see Rybicki and Hummer, 1983, their Equation 2.14) and where $K$ is a normalization constant ensuring $\int_{\lambda}\phi_{\lambda}{\rm d}\lambda=1$ (Rybicki and Lightman, 1979, their Equation 1.65).
+The differential term $\frac{{\rm d}x}{{\rm d}w}$ arises from integrating over all column densities associated with the observed velocity $u$ (Castor 1970, their Equation 6, Chisholm et al.
+2016, their Equation 4, assuming no stimulated emission).
 The corresponding integral limits are given by constraints from the geometry (§2.4).
+Following the equation of radiative transfer, one must sum up contributions of each velocity slice to the optical depth, not to the emergent flux, or else risk falling victim to Jensen’s inequality ($\exp[-\int\tau]<\int\exp[-\tau]$) by over-estimating the column density and optical depth (e.g., Jennings et al., 2025).
 The emergent flux, including a static component with a Voigt profile, is therefore given by
@@
 In the case of an outflow, we must first consider how the line emission depends on density and velocity.
-The line emission coefficient $j_{\lambda}$ is typically defined as proportional to the density of the populated upper state producing the transition (Rybicki1979; source:osterbrock2006).
-Since nebular lines emerge from upper electronic states populated by recombination or collisional excitation, $j\propto n[x(w)]^{2}$ (e.g., source:osterbrock2006; Peimbert_2017).
+The line emission coefficient $j_{\lambda}$ is typically defined as proportional to the density of the populated upper state producing the transition (Rybicki and Lightman, 1979; Osterbrock and Ferland, 2006).
+Since nebular lines emerge from upper electronic states populated by recombination or collisional excitation, $j\propto n[x(w)]^{2}$ (e.g., Osterbrock and Ferland, 2006; Peimbert et al., 2017).
 Thus, we can simply express the line profile as
-$\phi_{\lambda}=K^{-1}\int n\left[x[w]\right]^{2}{\rm d}\Omega$ (10)
-where $K$ again is a normalization constant ensuring $\int_{\lambda}\phi_{\lambda}{\rm d}\lambda=1$ (see Flury2023). As with absorption, limits on the integral are given by the geometry in §2.4.
+$\phi_{\lambda}=K^{-1}\int n\left[x[w]\right]^{2}{\rm d}\Omega^{\prime}$ (10)
+where $K$ again is a normalization constant ensuring $\int_{\lambda}\phi_{\lambda}{\rm d}\lambda=1$ (see Flury et al., 2023).
+As with absorption, limits on the integral are given by the geometry in §2.4.
 The emergent flux, including a static component with a Gaussian profile, is then given by
@@
 However, photons emitted in this fashion can be re-absorbed and subsequently re-emitted in an arbitrary direction, effectively a scattering process, which will alter the line profile.
-Resonant and fluorescent lines re-emit light which has previously been absorbed isotropically, meaning that the line flux is determined by the source function $S_{\lambda}=1-\exp[-\tau_{\lambda}]$ (see, e.g., Castor1970, their §6, Rybicki1983, their §3) such that
+Resonant and fluorescent lines re-emit light which has previously been absorbed isotropically, meaning that the line flux is determined by the source function $S_{\lambda}=1-\exp[-\tau_{\lambda}]$ (see, e.g., Castor, 1970, their §6, Rybicki and Hummer 1983, their §3) such that
 $\frac{F_{\lambda}}{F_{\lambda 0}}=\epsilon f_{\rm esc,\lambda}\left(1-\exp\left[-\tau_{\lambda}\right]\right)$ (12)
@@
 The escape fraction of photons depends on the fraction $p_{i}=A_{i}/\sum A$ of photons which are produced by a transition $i$ given by the transition rate $A_{i}$ divided by the total radiative depopulation rate $\sum A$, where each $i$ corresponds to a resonant ($p_{res}$) or fluorescent ($p_{fluor}$) channel.
-Following formalism from Castor1970, the probability that a resonant photon can escape from radiative “trapping” through resonant scattering is given by the ratio of emission to absorption $(1-\exp[-\tau_{\rm res,\lambda}])/\tau_{\rm res,\lambda}$.
+Following formalism from Castor (1970, their Equation A12, see also , their Equation 3.19), the probability that a resonant photon can escape from radiative “trapping” through resonant scattering is given by the ratio of emission to absorption $(1-\exp[-\tau_{\rm res,\lambda}])/\tau_{\rm res,\lambda}$.
 Trapped photons either scatter through emission and re-absorption in the resonant channel or re-emit in the fluorescent channel.
@@
 $f_{\rm esc,\lambda}^{\rm fluor}=p_{\rm fluor}\left(1-p_{\rm res}\frac{1-\exp[-\tau_{\rm res,\lambda}]}{\tau_{\rm res,\lambda}}\right)^{-1}$ (14)
-for fluorescent photons under the infinite scattering limit given by 2015ApJ...801...43S.
+for fluorescent photons under the infinite scattering limit given by Scarlata and Panagia (2015, their Equations 20a,b).
 2.2 Solutions to Wind Equations of Motion
-Velocity profiles for winds can take on a variety of shapes arising from different treatments of the wind equations of motion (e.g., CAK theory 1975ApJ...195..157C; 1976ApJS...32..715L; 1977PhDT.........3A; 1977ApJ...213..737B; Castor1979).
+Velocity profiles for winds can take on a variety of shapes arising from different treatments of the wind equations of motion (e.g., CAK theory Castor et al., 1975a; Lamers and Morton, 1976; Abbott, 1977; Barlow and Cohen, 1977; Castor and Lamers, 1979).
 The simple $\beta$ law approximation
 $w[x]=\left(1-x^{-1}\right)^{\beta}{\rm,\hskip 30.00005ptBetaCAK}$ (15)
-first appears in early treatment of Wolf-Rayet P Cygni lines for $\beta=0.5$ (Equation 23 Castor1970) and arises from a generalization of CAK theory solutions (see, e.g., Equation 47 of 1975ApJ...195..157C) without loss of physical insight (see, e.g., 1977ApJ...213..737B; Castor1979; 1986A&A...164...86P).The CAK $\beta$ law approximation is ubiquitous in the study of stellar winds (e.g., 2001A&A...369..574V; 2021MNRAS.504.2051V; Telford2024) and has even seen application to galactic outflows (e.g., 2016MNRAS.463..541C; Flury2023).
+first appears in early treatment of Wolf-Rayet P Cygni lines for $\beta=0.5$ (Equation 23 Castor, 1970) and arises from a generalization of CAK theory solutions (see, e.g., Equation 47 of Castor et al., 1975a) without loss of physical insight (see, e.g., Barlow and Cohen, 1977; Castor and Lamers, 1979; Pauldrach et al., 1986).The CAK $\beta$ law approximation is ubiquitous in the study of stellar winds (e.g., Vink et al., 2001; Vink and Sander, 2021; Telford et al., 2024) and has even seen application to galactic outflows (e.g., Chisholm et al., 2016; Flury et al., 2023).
 Given its widespread adoption and underlying physical motivation, the CAK $\beta$ law is our default and fiducial velocity field.
 In addition to the $\beta$ law approximation to CAK theory, OutLines includes two alternate solutions to the wind equations proposed in the literature.
-The first, from 2010ApJ...717..289S, assumes a power law approximation to the radial acceleration field to obtain
+The first, from Steidel et al.
+(2010), assumes a power law approximation to the radial acceleration field to obtain
 $w[x]=\left(1-x^{1-\beta}\right)^{1/2}{\rm,\hskip 30.00005ptAccPlaw}$ (16)
 where the normalization terms are implicit in $w$ and $x$ and $r_{min}=1$.
-While those authors do not provide a corresponding physical model, their acceleration power law has a velocity field qualitatively similar to those of Ipavich1975, which assumed cosmic ray driving, and 1985Natur.317...44C, which assumed constant momentum and energy deposition into a steady-state wind from a fluid in which it is entrained (see also, e.g., Owocki2025), and formally akin to radiative acceleration of expanding clouds (see Equations 16-17 Mathews1974).
+While those authors do not provide a corresponding physical model, their acceleration power law has a velocity field qualitatively similar to those of Ipavich (1975), which assumed cosmic ray driving, and Chevalier and Clegg (1985), which assumed constant momentum and energy deposition into a steady-state wind from a fluid in which it is entrained (see also, e.g., Owocki, 2025), and formally akin to radiative acceleration of expanding clouds (see Equations 16-17 Mathews, 1974).
 Note that the acceleration power law field with $\beta=2$ is identical to that of the CAK $\beta$ law when $\beta=0.5$.
-The second velocity field alternative, from various sources (e.g., FaucherGiguere2012), constitutes a power law approximation to the radial velocity field such that
+The second velocity field alternative, from various sources (e.g., Faucher-Giguère and Quataert, 2012), constitutes a power law approximation to the radial velocity field such that
 $w[x]=A(x-1)^{\beta}{\rm,\hskip 30.00005ptVelPlaw}$ (17)
@@
 We set $A=0.5$ to allow clouds to reach 50% terminal velocity in roughly the same range of radii as the CAK theory $\beta$ law and generally overlap with the same region in $w$-$x$ space as both the CAK $\beta$ law and acceleration power law fields.
-Single-shell models of stellar wind bubbles can give rise to power law velocity fields from the shared time dependence of $r\propto t^{3/5}$ and $v\propto t^{-2/5}$ ($v\propto r^{-2/3}$, e.g., Weaver1977; Dyson1977).
-Building on this framework, FaucherGiguere2012 arrive at a velocity power law solution assuming continuous production of “shells” by an AGN to produce a wind.
+Single-shell models of stellar wind bubbles can give rise to power law velocity fields from the shared time dependence of $r\propto t^{3/5}$ and $v\propto t^{-2/5}$ ($v\propto r^{-2/3}$, e.g., Weaver et al., 1977; Dyson, 1977).
+Building on this framework, Faucher-Giguère and Quataert (2012) arrive at a velocity power law solution assuming continuous production of “shells” by an AGN to produce a wind.
+While the OutLines default assumes that gas is launched from rest, we have allowed for the possibility that the gas starts from some launch velocity $v_{0}$, which we include as an optional free parameter.
+The velocity limits are accordingly updated so that the minimum velocity $w_{\ell}$ which contributes to an observed velocity $|u|$ is never below $v_{0}/v_{\infty}$.
+The velocity fields in Table 4 and Equations 15-17 are modified by a factor of $1-v_{0}/v_{\infty}$ with the addition of a constant $v_{0}/v_{\infty}$ such that the relative velocity not from rest $w_{\rm nfr}$ is
+$w_{\rm nfr}=w[x|\beta]\left(1-\frac{v_{0}}{v_{\infty}}\right)+\frac{v_{0}}{v_{\infty}}$ (18)
+(e.g., Castor and Lamers, 1979).
+Emergent profiles are similar to those presented in Tenorio-Tagle et al.
+(1996), which included distributions of non-zero shell ages, and to those presented in Scarlata and Panagia (2015), which account for non-zero launch velocities.
+For visualization purposes, we assume in example profiles in §2 that $v_{0}=0.1v_{\infty}$.
 Figure 1 depicts the different velocity fields for a range of $\beta$.
@@
 Moreover, the power law velocity field does not approach $v_{\infty}$ asymptotically as do the other velocity fields, meaning that the stopping condition at $v_{\infty}$ is imposed solely by integral limits rather than the formalism for the velocity field.
-We note that physical wind models almost always exhibit an acceleration which gradually decreases as the outflow asymptotically approaches some terminal velocity (e.g., 1975ApJ...195..157C; 1977ApJ...213..737B; 1985Natur.317...44C; Thompson2015; 2017MNRAS.471.4061K; 2018MNRAS.476..512I; Quataert2022a), indicating that the underlying wind physics strongly favors a CAK $\beta$ law or acceleration power law field over a velocity power law field.
-While individual wind bubble models indicate a velocity power law field (e.g., Weaver1977; Dyson1977), in ensembles of shells, the distribution of shell ages will dominate the net velocity field (e.g., TenorioTagle1996) and is likely determined by the physical driving mechanism (e.g., Chu1994), suggesting that the CAK theory or acceleration power law velocity fields provide a more physically appropriate description for wind scenarios.
-That being said, the velocity power law field prescription can be consistent with turbulent streaming including the classic Kolmogorov cascade, with the case of $\beta=1$ indicative of Couette flow (e.g., 2009A&A...500..817B), although such scenarios may not capture the underlying driving mechanism.
+We note that physical wind models almost always exhibit an acceleration which gradually decreases as the outflow asymptotically approaches some terminal velocity (e.g., Castor et al., 1975a; Barlow and Cohen, 1977; Chevalier and Clegg, 1985; Thompson et al., 2015; Krumholz et al., 2017; Ishibashi et al., 2018; Quataert et al., 2022b), indicating that the underlying wind physics strongly favors a CAK $\beta$ law or acceleration power law field over a velocity power law field.
+While individual wind bubble models indicate a velocity power law field (e.g., Weaver et al., 1977; Dyson, 1977), in ensembles of shells, the distribution of shell ages will dominate the net velocity field (e.g., Tenorio-Tagle et al., 1996) and is likely determined by the physical driving mechanism (e.g., Chu and Kennicutt, 1994), suggesting that the CAK theory or acceleration power law velocity fields provide a more physically appropriate description for wind scenarios.
+That being said, the velocity power law field prescription can be consistent with turbulent streaming including the classic Kolmogorov cascade, with the case of $\beta=1$ indicative of Couette flow (e.g., Binette et al., 2009), although such scenarios may not capture the underlying driving mechanism.
 Figure 1: Radial velocity fields included in OutLines and given by Equations 15-17.
 Acceleration from rest at the base $R_{0}$ of the wind ($x=1$) proceeds until the wind reachs the terminal velocity $v_{\infty}$ ($w[x]=1$).
-Velocity fields include the 1986A&A...164...86P $\beta$ law approximation to CAK theory (1975ApJ...195..157C) (blue, Equation 15), the 2010ApJ...717..289S power law acceleration (green, Equation 16), and the power law velocity field (purple, Equation 17).
+Velocity fields include the Pauldrach et al.
+(1986) $\beta$ law approximation to CAK theory (Castor et al., 1975a) (blue, Equation 15), the Steidel et al.
+(2010) power law acceleration (green, Equation 16), and the power law velocity field (purple, Equation 17).
 Dash length increases with increasing $\beta$.
@@
 2.3 Radial Density Profiles
-The gas / clouds within a wind likely vary with projected radial distance from the launching source due to a variety of factors, including conservation of cloud surface area (e.g., 2017MNRAS.471.4061K), fountains (e.g., 2015ApJ...814...83L), or an isothermal mass-conserving gas (e.g., 2005ApJS..160..115R; 2011ApJ...734...24P).
-For simplicity, gas number density $n$ can also be assumed to follow a power law such that
+The gas / clouds within a wind likely vary with projected radial distance from the launching source due to a variety of factors, including conservation of cloud surface area (e.g., Krumholz et al., 2017), fountains (e.g., Leroy et al., 2015), or an isothermal mass-conserving gas (e.g., Rupke et al., 2005b; Prochaska et al., 2011).
+We adopt a smooth function to describe a continuous distribution of delta functions, with each delta function “pulse” representing an instantaneous burst or ejection of material with gas density $n$ from the launch site and projected radially.
+For simplicity, $n$ can be assumed to follow a power law such that
 $n[x]=x^{-\alpha}{\rm,\hskip 30.00005ptPowerLaw}$ (19)
-a reasonable approximation for stellar and galactic winds (e.g., Castor1979; 1985Natur.317...44C, respectively) and a common choice in the literature (e.g., FaucherGiguere2012; 2015ApJ...801...43S; 2016MNRAS.463..541C; Flury2023; Luminari2024A&A) arising from power law solutions to differential equations (as also the case for the velocity profile) and can be indicative of momentum or energy conservation ($\beta=(\alpha-2)/2$ and $\beta=(\alpha-2)/3$, respectively FaucherGiguere2012, for the power law velocity field $\beta$).
+a reasonable approximation for stellar and galactic winds (e.g., Castor and Lamers, 1979; Chevalier and Clegg, 1985, respectively) and a common choice in the literature (e.g., Faucher-Giguère and Quataert, 2012; Scarlata and Panagia, 2015; Chisholm et al., 2016; Flury et al., 2023; Luminari et al., 2024) arising from power law solutions to differential equations (as also the case for the velocity profile) and can be indicative of momentum or energy conservation ($\beta=(\alpha-2)/2$ and $\beta=(\alpha-2)/3$, respectively Faucher-Giguère and Quataert, 2012, for the power law velocity field $\beta$).
+Perhaps intuitively, $\alpha=2$ represents a mass-conserving scenario with density changing only as a $r^{-2}$ projection effect.
 As an alternative, we can adopt an exponential profile where
@@
 with a break-point or “knee” at radius $x_{1}$, inner slope $a_{1}$ and outer slope $a_{2}$.
-Examples of a broken power law include the shock-driven outflow launched by the interaction of winds from massive stars in proximity to Sag A ($a_{1}<a_{2}$, drop off in density profile, e.g., Quataert2004) and the relativistic outflow triggered by the waking/activating supermassive black hole in Swift J164449.3+57345 ($a_{1}>a_{2}$, flattening in density profile, e.g., Berger2012).
+Examples of a broken power law include the shock-driven outflow launched by the interaction of winds from massive stars in proximity to Sag A ($a_{1}<a_{2}$, drop off in density profile, e.g., Quataert, 2004) and the relativistic outflow triggered by the waking/activating supermassive black hole in Swift J164449.3+57345 ($a_{1}>a_{2}$, flattening in density profile, e.g., Berger et al., 2012).
 For a variety of physical scenarios, including swept up shells and blast waves, we also allow in Outlines the possibility of pulse-like radial density profiles consistent with a shell or bubble with a scaled radius $x_{1}=R_{\rm shell}/R_{0}$ and characteristic width $\sigma_{x}=\sigma_{r}/R_{0}$, including a shell with gas in a normal distribution
@@
 $n[x]=\left(2\pi(\log\sigma_{x})^{2}\right)^{-1/2}\exp\left[-\frac{(\log x-\log x_{1})^{2}}{2(\log\sigma_{x})^{2}}\right]{\rm,\hskip 10.00002ptLogNormal,}$ (23)
-a fast-rise exponential decay (or FRED, Equation 23, see Norris2005) distribution
+a fast-rise exponential decay (or FRED, Equation 24, see Norris et al. 2005) distribution
 $\begin{split}n[x]=&\exp[2\mu]\exp\left[-\frac{\tau_{1}}{x+\mu-x_{1}}-\frac{x+\mu-x_{1}}{\tau_{2}}\right],\\ &\mu=\sqrt{\tau_{1}\tau_{2}}\end{split}{\rm\hskip 10.00002ptFRED,}$ (24)
+a logistic column density distribution
+$n[x]=\frac{k\exp[-k(x-x0)]}{(1+\exp[-k(x-x0)])^{2}}{\rm,\hskip 10.00002ptDLogistic,}$ (25)
 and a uniform distribution
@@
 The emission line profile wings for the shell or pulse-like density distributions are distinct from the continuous distributions; however, distinctions among the various subsets are less clear, likely owing to the geometric dilution of the density profile by integrating over the entire volume of the line-emitting gas.
-The effect of pulse profiles on the emission line profile is largely consistent with early fixed-velocity single expanding shell models for Wolf-Rayet features and stellar wind bubbles wherein no line core emerges (e.g., Beals1931; TenorioTagle1996).
-Such marked differences in radial gas distribution may explain the line profiles of ionized gas around pre- and post-SNe stellar populations as in the different clusters in the Sunburst Arc (e.g., Mainali2022).
+The effect of pulse profiles on the emission line profile is largely consistent with early fixed-velocity single expanding shell models for Wolf-Rayet features and stellar wind bubbles wherein no line core emerges (e.g., Beals, 1931; Tenorio-Tagle et al., 1996).
+Such marked differences in radial gas distribution may explain the line profiles of ionized gas around pre- and post-SNe stellar populations as in the different clusters in the Sunburst Arc (e.g., Mainali et al., 2022).
 The absorption line profiles are more distinct for each individual density profile, likely owing to the direct tracing of the line-of-sight gas distribution without any geometric dilution.
 Deviations from spherical geometry may affect this result, particularly in the case of a narrow bicone normal to the plane of the sky and aligned with the observer, which directly traces the density distribution in emission and absorption.
-Figure 3: Radial density profiles included in OutLines and given by Equations 18-23.
-Pulse-like density profiles, including the normal (‘Normal’, green, Eqn 21), log normal (‘LogNormal’, yellow, Eqn 22), and fast-rise exponential decay (‘FRED’, red, Eqn 23) distributions, represent physical scenarios such as blast waves, expanding shells, and single burst events.
-For visualization, the pulse-like profiles are arbitrarily set to maximize density at $r=3R_{0}$.
-Continuous-like density profiles, including power law (‘PowerLaw’, purple, Eqn 18), exponential (‘Exponential’, orange, Eqn 19), and double power law (‘DoublePowerLaw’, blue, Eqn 20) distributions, represent an integral over many pulses and will always maximize density at the base of the outflow where $r=R_{0}$.
-Figure 4: Nebular emission (top) and absorption (bottom) line profiles predicted by OutLines for fixed density, column density, and velocity field for different density distributions assuming a spherical geometry.
+Figure 3: Radial density profiles included in OutLines and given by Equations 19-28.
+Top: Wind-like density profiles, including power law (‘PowerLaw’, purple, Eqn 19), double power law (‘DoublePowerLaw’, yellow, Eqn 21), and exponential (‘Exponential’, green, Eqn 20) distributions.
+Center: Pulse-like density profiles, including the normal (‘Normal’, blue, Eqn 22), log normal (‘LogNormal’, red, Eqn 23), fast-rise exponential decay (‘FRED’, orange, Eqn 24), and derivative of the logistic (‘DLogic’, indigo, Eqn 25) distributions, represent physical scenarios such as blast waves, expanding shells, and single burst events.
+Bottom: Ensembles of pulse-like profiles without (black solid, Eqn 27) and with (black dashed, Eqn 28) damping.
+These ensembles represent the resolved elements of a radially clumpy wind or a series of outflowing episodes.
+Figure 4: Nebular emission (top) and absorption (bottom) line profiles predicted by OutLines for fixed flux, column density, and velocity field for different density distributions assuming a spherical geometry.
+Colors matched to the corresponding profiles in Figure 3.
 […]
@@
 We show the effects of a series of pulse-like events or damped pulses on the emergent line profiles in Figure 5 for shells resolved and unresolved in velocity.
+Figure 5 illustrates how infinitely thin shells with infinitely small separations ($\sigma_{x},x_{k}\to 0$) or relatively thick shells ($\sigma_{k}\gtrsim x_{k}$) produce a wind-like profile.
 One might invoke the central limit theorem to argue that a system of many shells or bubbles will result in a Gaussian line profile; however, we find that neither the damped nor undamped pulses behave in this manner.
-Indeed, TenorioTagle1996 find that only a Gaussian velocity field will produce an intrinsically Gaussian line profile for a sum of many shells, a scenario which does not appear to be consistent with the bulk motion of gas predicted by theory.
+Indeed, Tenorio-Tagle et al.
+(1996) find that only a Gaussian velocity field will produce an intrinsically Gaussian line profile for a sum of many shells, a scenario which does not appear to be consistent with the bulk motion of gas predicted by theory.
 They further demonstrate that profiles might only appear Gaussian due to smearing by a line spread function.
-These results are consistent with comparisons of integrated and spatially resolved observations of the ensemble of expanding shells in 30 Doradus where the bulk velocity field produces a line profile with a Gaussian-like core but distinctly non-Gaussian wings (Chu1994).
-Figure 5: Nebular emission (top) and absorption (bottom) line profiles predicted by OutLines for a series of expanding bubbles, shells, or other pulse-like episodes in cases of unresolved (orange) or resolved (blue) shells with (dashed) and without (solid) damping to mimic the effects of dissipation.
+These results are consistent with comparisons of integrated and spatially resolved observations of the ensemble of expanding shells in 30 Doradus where the bulk velocity field produces a line profile with a Gaussian-like core but distinctly non-Gaussian wings (Chu and Kennicutt, 1994).
+Figure 5: Nebular emission (top) and absorption (bottom) line profiles predicted by OutLines for an ensemble of expanding bubbles, shells, or other pulse-like episodes in cases of thick (orange, $\sigma_{x}=0.5$) or thin (blue, $\sigma_{x}=0.1$) widely-spaced shells and thin closely-spaced shells (green, $x_{k}=0.2$) with (dashed) and without (solid) damping to mimic the effects of dissipation.
 In all cases, density, column density, velocity field, and terminal velocity are fixed, and a spherical geometry is assumed.
@@
 2.4 Geometry
-Following Flury2023, line emission or absorption is distributed evenly across the polar coordinate for each deprojected velocity (see also, e.g., Beals1931; Zheng1990; TenorioTagle1996; 2015ApJ...801...43S; Xrism2025_PDS456) allowing us to rewrite the differential solid angle as
-${\rm d}\Omega={2\pi}{\rm d}w$ (27)
+For any given radial velocity $w$, line emission or absorption is distributed evenly across a given projection angle $\varrho$ (the polar coordinate), which we illustrate here under full special relativistic treatment.
+The special relativistic velocity projection of $w$ in an arbitrary polar direction $\varrho^{\prime}$ in the observer’s frame incorporates both longitudinal and transverse effects such that, following Einstein (1905, their §7),
+$\lambda=\lambda_{0}\gamma\left(1+\omega\cos\varrho^{\prime}\right)$ (29)
+where $\omega=w\>v_{\infty}/c$ and $\gamma^{-1}=\sqrt{1-\omega^{2}}$ is the Lorenzt factor.
+Additionally from §7 of Einstein (1905), the angle $\varrho$ in the source frame requires a relativistic aberration correction to the observer’s frame $\varrho^{\prime}$, with a differential which reduces to
+$\sin\varrho^{\prime}{\rm d}\varrho^{\prime}=\left(\frac{\lambda_{0}}{\lambda}\right)^{2}\sin\varrho{\rm d}\varrho.$ (30)
+Combining these identities and respective differentials with Equation 1 provides a convenient substitution for the $\sin\varrho{\rm d}\varrho$ term in ${\rm d}\Omega$ such that
+${\rm d}\Omega^{\prime}=2\pi\frac{\sqrt{1-\omega^{2}}}{\omega(1-\omega)^{2}}\sqrt{\frac{1-\omega}{1+\omega}}{\rm d}\omega$ (31)
 […]
-To illustrate the effects of assumed velocity field and related cosine projection onto the line of sight, we illustrate contours of constant observed velocity (isocontours) in Figure 6 (see also Castor1970, their Figure 1, 1987ApJ...314..726L, their Figure 2).
-Sometimes referred to as velocity surfaces (as in Castor1970), these isocontours represent the velocities which are cosine projected to a single observed velocity, assuming a particular velocity field and varying the power law index.
+To illustrate the effects of assumed velocity field and related $\cos\varrho$ projection onto the line of sight, we illustrate contours of constant observed velocity (isocontours) in Figure 6 (see also Castor 1970, their Figure 1, Lucy, 1971, their Figure 2).
+Sometimes referred to as velocity surfaces (as in Castor, 1970), these isocontours represent the velocities which are cosine projected to a single observed velocity, assuming a particular velocity field and varying the power law index.
 Figure 6: Velocity isocontours for the different velocity fields available in OutLines: the $\beta$ law approximation to CAK theory (top), an acceleration power law (middle), and a velocity power law (bottom).
@@
 2.4.1 Velocity Limits
-The equal contributions of each deprojected velocity to the absorption or emission allows us to reduce the geometry to the velocity limits $w_{\ell}$ and $w_{u}$ of the integrals in Equations 5 and 10.
+The weighted contributions of each deprojected velocity to the absorption or emission allows us to reduce the geometry to the velocity limits $w_{\ell}$ and $w_{u}$ of the integrals in Equations 5 and 10.
 Boundary conditions for velocity in the case of a uniform spherical outflow (illustrated in Figure 7) are given by some minimum and maximum projected velocities $w_{\ell},\ w_{u}$, respectively.
@@
 From Figure 7, the line-emitting gas posterior to the source (redshifted in the case of an outflow) cannot be seen by the observer, meaning that $w_{\ell}$ must account for obstruction by the launching source at $x=1$.
-Solutions for the source obstruction velocity limit are given in Flury2023, which we summarize here.
+Solutions for the source obstruction velocity limit are given in Flury et al.
+(2023), which we summarize here.
 From the Pythagorean theorem, the minimum velocity $w_{\ell}$ must satisfy
@@
 The limit of $w_{a}$ most dramatically affects the projected velocities close to the transverse axis.
-As a result, aperture clipping will predominantly truncate the core of the profile close to the central wavelength (see also 2015ApJ...801...43S); i.e., the broadest extent of the wings remains relatively unaffected.
+As a result, aperture clipping will predominantly truncate the core of the profile close to the central wavelength (see also Scarlata and Panagia, 2015); i.e., the broadest extent of the wings remains relatively unaffected.
 Moreover, absorption lines arise from gas anterior to the launch site that is backlit by the source, resulting in a “pencil beam” effect where the maximum velocity contributing to the profile is $w_{0}$ (hatches in Figure 7).
@@
 Colored circles indicate different deprojected velocities $w$ which can contribute to a single observed velocity surface $u$ with values of $u/w=0.15,0.5,0.85$ corresponding to yellow/green, orange/blue, and red/purple, respectively.
-Figure 8: Effect of aperture clipping on the observed line profile for nebular emission (left panel) and for a P Cygni line consisting of absorption and resonant emission with unity filling factor and $p_{res}=0.5$ (right panel).
+Figure 8: Effect of aperture clipping on the observed line profile for a P Cygni line consisting of absorption and resonant emission with filling factor $\epsilon=0.2$ and escape fraction $p_{res}=1$ (left panel) and for nebular emission (right panel).
 For visualization, lines are normalized to the case of no aperture clipping ($w_{a}=1$).
-Note that the P Cygni line simplifies to pure absorption in the case where $w_{a}=0$ ($w_{1}=w_{0}$) as in 2015ApJ...801...43S and that the effect of radiative trapping on the line core decreases as $w_{a}$ decreases.
+Dotted black line in the left panel indicates pure absorption, demonstrating that infilling can still occur in the fully aperture-clipped limit but that only an absorption profile is observed (cf.
+Scarlata and Panagia, 2015).
+Note that the effect of radiative trapping on the resonant line core decreases as $w_{a}$ decreases.
 Figure 9: Integral velocity limits for the emission line profile $\phi_{\lambda}$.
@@
 2.4.2 Directed Outflows
-Many outflows are not spherically uniform but rather are observed as bi-directional cones on a wide variety of scales (e.g., Fukui1986; Clegg1987; Heckman1990; Colina1991; Hirano1992; Marlow1995; Cameron2021).
-We consider a bi-directional conical geometry with a half opening angle $\theta_{o}$ and an inclination angle $i$ with respect to the observer, which we depict in Figure 10 (see also, e.g., Figure 5 in Hjelm1996).
+Many outflows are not spherically uniform but rather are observed as bi-directional cones on a wide variety of scales (e.g., Fukui et al., 1986; Clegg et al., 1987; Heckman et al., 1990; Colina et al., 1991; Hirano et al., 1992; Marlowe et al., 1995; Cameron et al., 2021).
+We consider a bi-directional conical geometry with a half opening angle $\theta_{o}$ and an inclination angle $i$ with respect to the observer, which we depict in Figure 10 (see also, e.g., Figure 5 in Hjelm and Lindblad 1996).
 OutLines includes a tool PlotGeometry for producing such figures as a visual aid.
@@
 OutLines computes the conical projection from $i,\theta_{o}$ and then calculates $\ell_{\rm w}$ by determining the angular length(s) of each deprojected velocity inscribed by the projected cone.
-For details on this method, we refer the reader to Appendix A (see, e.g., Zheng1990; Carr2018; Luminari2018, for earlier approaches to solving this problem).
+For details on this method, we refer the reader to Appendix A (see, e.g., Zheng et al., 1990; Carr et al., 2018; Luminari et al., 2018, for earlier approaches to solving this problem).
 We illustrate the effects of changing the opening angle $\theta_{o}$ and inclination angle $i$ in Figure 11.
@@
 For visualization, line profiles are all normalized to the spherical case ($i=0^{\circ}$, $\theta=90^{\circ}$).
+The effects of inclination on the emergent line profile shown in the upper left are similar to the 3D Monte Carlo simulation predictions for X-ray fluorescent lines from the skirt code (Vander Meulen et al., 2024; Bianchi et al., 2026), with face-on inclination yielding broader and flatter profiles and edge-on inclination yielding narrower and peakier profiles.
 2.4.3 Directed Outflows with Cavities
-An additional geometry observed in young stellar objects (e.g., Padgett1999; Louvet2018; deValon2020), stellar winds (e.g.
-Sirianni1998), starbursts (e.g., Axon1978; Marlow1995; Heckman1995; Rich2010; Yoshida2010), and AGN (e.g., Phillips1983; Hjelm1996; Crenshaw2000a; Crenshaw2000b; Muller2011; Venturi2018) is that of hollow cones of outflowing material.
-We include this geometry, assuming some cavity angle $\theta_{c}$ and opening angle $\theta_{o}$ (see also models by Das2006; Muller2011; Bae2016; Luminari2024A&A).
+An additional geometry observed in young stellar objects (e.g., Padgett et al., 1999; Louvet et al., 2018; de Valon et al., 2020), stellar winds (e.g.
+Sirianni et al., 1998), starbursts (e.g., Axon and Taylor, 1978; Marlowe et al., 1995; Heckman et al., 1995; Rich et al., 2010; Yoshida et al., 2011), and AGN (e.g., Phillips et al., 1983; Hjelm and Lindblad, 1996; Crenshaw and Kraemer, 2000; Crenshaw et al., 2000; Müller-Sánchez et al., 2011; Venturi et al., 2018) is that of hollow cones of outflowing material.
+We include this geometry, assuming some cavity angle $\theta_{c}$ and opening angle $\theta_{o}$ (see also models by Das et al., 2006; Müller-Sánchez et al., 2011; Bae and Woo, 2016; Luminari et al., 2024).
 OutLines implements this geometry by subtracting the fractional angular length $\ell_{\rm cavity}$ of each deprojected velocity inscribed by the cavity from fractional the angular length inscribed by the total cone (see Appendix A).
@@
 2.4.4 An Obstructing Disk
-The outflow may be launched from a region embedded in a disk of optically thick material, a phenomenon observed in a variety of astrophysical scenarios: expelled debris in planetary nebulae (e.g., Icke1981), protoplanetary material around young stellar objects (e.g., Padgett1999; Ilee2016), the dusty material concentrated around an AGN (e.g., Veilleux2001), or even the plane of a galaxy (e.g., Hjelm1996; Cronin2025).
+The outflow may be launched from a region embedded in a disk of optically thick material, a phenomenon observed in a variety of astrophysical scenarios: expelled debris in planetary nebulae (e.g., Icke, 1981), protoplanetary material around young stellar objects (e.g., Padgett et al., 1999; Ilee et al., 2016), the dusty material concentrated around an AGN (e.g., Veilleux et al., 2001), or even the plane of a galaxy (e.g., Hjelm and Lindblad, 1996; Cronin et al., 2025).
 In such cases, the disk will obstruct some of the observer’s view of the outflow.
@@
 We elaborate on this formalism in Appendix A.
-Because a dense disk of material will likely suppress any outwardly moving material within its plane, OutLines only implements a disk if cones are present.
+OutLines implements a disk either in tandem with the cone geometries or alone as a hemisphere geometry.
+Figure 13: Same as Figure 11 but for the effects of an obstructing disk as a function of inclination with respect to the observer in (left) a thin single-shell hemisphere geometry where all material posterior to the disk is obstructed (cf.
+Xrism Collaboration et al., 2025, their Fig 12) and (right) a biconical wind with $\theta_{0}=30^{\circ}$ and $x_{\rm disk}=5$.
+Colors indicate changes to the inclination angle $i$, illustrating the redistribution of flux from entirely blue-shifted ($i=0^{\circ}$) to evenly blue- and red-shifted ($i=90^{\circ}$).
 Figure 14: Same as Figure 11 but for bidirectional cones with fixed inclination and opening angle and an obscuring disk with varying radii, where a radius of 0 indicates a disk-free line profile.
@@
 For visualization, line profiles are normalized to the disk-free case in each panel.
-We illustrate the effects of including an obstructing disk in Figure 13.
+We illustrate the effects of including an obstructing disk for the hemisphere and bidirectional cone cases in Figure 13 and for the bidirectional cones in Figure 14.
 By definition, the disk will only obstruct the posterior cone, so if $i+\theta_{o}<90^{\circ}$, the disk will only obstruct the receding outflow, solely obstructing the redshifted component (left of Figure 14).
@@
 2.5 Outflow Properties
-Following Flury2023, the characteristic outflow radius $x_{out}$ corresponds to the maximum momentum density $\dot{\rho}_{p}=\mu m_{\rm H}n_{0}v_{\infty}n[x]w[x]$, i.e., where the wind effects are strongest.
-In the case of pulse-like density distributions, the center of the gas distribution $x_{1}$ will be close to, but not always the same as, $x_{out}$.The radius $x_{out}$ of maximum momentum density scales all other outflow properties, first and foremost the characteristic outflow velocity $v_{out}$ by determining the velocity at $x_{out}$ using the velocity field $w$ such that
-$v_{out}=v_{\infty}w[x_{out}].$ (32)
-The effective column density $N_{out}$ at the characteristic radius can be determined by integrating over the density profile from the base of the outflow ($x=1$) to the peak momentum density $x=x_{out}$ (e.g., Flury2023, their Equation 13 for the case of a CAK velocity field and power law density profile).
-Note that $N_{out}$ is not to be confused with $N$ from the absorption line profile, which is the total column density along the line of sight.
-$N_{out}$ necessarily depends on the space density $n_{0}$ and the radius at the base of the outflow $R_{0}$, which are unconstrained parameters in OutLines.
+Following Flury et al.
+(2023, see their Appendix B), the characteristic outflow radius $x_{\rm out}$ corresponds to the maximum momentum density $\dot{\rho}_{p}=\mu m_{\rm H}n_{0}v_{\infty}n[x]w[x]$, i.e., where the wind effects are strongest.
+In the case of pulse-like density distributions, the center of the gas distribution $x_{1}$ will be close to, but not always the same as, $x_{\rm out}$.The radius $x_{\rm out}$ of maximum momentum density scales all other outflow properties, first and foremost the characteristic outflow velocity $v_{\rm out}$ by determining the velocity at $x_{\rm out}$ using the velocity field $w$ such that
+$v_{\rm out}=v_{\infty}w[x_{\rm out}].$ (36)
+The effective column density $N_{\rm out}$ at the characteristic radius can be determined by integrating over the density profile from the base of the outflow ($x=1$) to the peak momentum density $x=x_{\rm out}$ (e.g., Flury et al., 2023, their Equation 13 for the case of a CAK velocity field and power law density profile).
+Note that $N_{\rm out}$ is not to be confused with $N$ from the absorption line profile, which is the total column density along the line of sight.
+$N_{\rm out}$ necessarily depends on the space density $n_{0}$ and the radius at the base of the outflow $R_{0}$, which are unconstrained parameters in OutLines.
 As such, OutLines computes the relative column density $\mathcal{R}$ where
-$\frac{N_{out}}{n_{0}R_{0}}=\mathcal{R}=\int\limits_{1}^{x_{out}}n[x]{\rm d}x.$ (33)
+$\frac{N_{\rm out}}{n_{0}R_{0}}=\mathcal{R}=\int\limits_{1}^{x_{\rm out}}n[x]{\rm d}x.$ (37)
 This quantity $\mathcal{R}$ becomes important for computing the strength of the outflow expressed via $\dot{M}$, $\dot{p}$, and $\dot{E}$.
-The mass outflow rate $\dot{M}$ is given by the momentum flux $\mu m_{\rm H}n_{0}v_{out}$ with a geometric projection $4\pi R_{0}^{2}(\cos\theta_{c}-\cos\theta_{o})$ such that
-$\dot{M}=4\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}x_{out}\mathcal{R}\mu m_{\rm H}v_{out},$ (34)
-a more general case of Equation 12 from Flury2023.
+The mass outflow rate $\dot{M}$ is given by the momentum flux $\mu m_{\rm H}n_{0}v_{\rm out}$ with a geometric projection $4\pi R_{0}^{2}(\cos\theta_{c}-\cos\theta_{o})$ such that
+$\dot{M}=4\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}v_{\rm out},$ (38)
+a more general case of Equation 12 from Flury et al.
+(2023).
 Such formalism is frequent in the literature with differences in implementation of the geometric projection (e.g.
-2005ApJS..160...87R; 2005ApJS..160..115R; Crenshaw2012; Maiolino2012; Fiore2017; Luminari2018; Luminari2024A&A; Flury2023).
+Rupke et al., 2005a, b; Crenshaw and Kraemer, 2012; Maiolino et al., 2012; Fiore et al., 2017; Luminari et al., 2018, 2024; Flury et al., 2023).
 The outflow momentum rate $\dot{p}$ expresses the rate at which momentum is injected into the outflow.
-Typically, $\dot{p}$ is defined as the product of outflow velocity and mass outflow rate (e.g., 2005ApJS..160..115R, their Equation 9) such that
-$\dot{p}_{out}=v_{out}\dot{M}_{out}=4\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}v_{out}^{2}.$ (35)
+Typically, $\dot{p}$ is defined as the product of outflow velocity and mass outflow rate (e.g., Rupke et al., 2005b, their Equation 9) such that
+$\dot{p}=v_{\rm out}\dot{M}=4\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}v_{\rm out}^{2}.$ (39)
 The outflow mechanical luminosity $\dot{E}$ expresses the injection of kinetic energy into the outflow.
-Typically, $\dot{E}$ is defined as the product of outflow velocity and mass outflow rate (e.g., 2005ApJS..160..115R, their Equation 11) such that
-$\dot{E}_{out}\begin{split}=&\frac{1}{2}\dot{M}\left(v_{out}^{2}+3\sigma_{v}^{2}\right)\\ =&2\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}\left(v_{out}^{3}+3v_{out}\sigma_{v}^{2}\right).\end{split}$ (36)
-These values can be readily computed using constraints on $x_{out}$ and $v_{\infty}$ and possible geometric terms $\theta_{o}$ and $\theta_{c}$, all obtained from the line profile.
+Typically, $\dot{E}$ is defined as the product of outflow velocity and mass outflow rate (e.g., Rupke et al., 2005b, their Equation 11) such that
+$\dot{E}_{out}\begin{split}=&\frac{1}{2}\dot{M}\left(v_{\rm out}^{2}+3\sigma_{v}^{2}\right)\\ =&2\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}\left(v_{\rm out}^{3}+3v_{\rm out}\sigma_{v}^{2}\right).\end{split}$ (40)
+These values can be readily computed using constraints on $x_{\rm out}$ and $v_{\infty}$ and possible geometric terms $\theta_{o}$ and $\theta_{c}$, all obtained from the line profile.
 The only two unknowns are $n_{0}$ and $R_{0}$, which cannot be directly informed by the line profile and require additional measurements (spatial for $R_{0}$, spectral for $n_{0}$) or assumptions based on prior knowledge.
-For instance, $n_{0}=10\rm~cm^{-3}$ and $R_{0}=0.5$ kpc may be reasonable for galactic outflows (e.g., 1985Natur.317...44C; Flury2023).
+For instance, $n_{0}=10\rm~cm^{-3}$ and $R_{0}=0.5$ kpc may be reasonable for galactic outflows (e.g., Chevalier and Clegg 1985; Flury et al.
+2023).
 As implemented here, we compute $\dot{M}$, $\dot{p}$, and $\dot{E}$ normalized to $n_{0}R_{0}^{2}$ with $n_{0}$ in cm-3 and $R_{0}$ in kpc, meaning the characteristics of the outwardly moving gas are written as
-$\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=0.191\mathcal{R}[\cos\theta_{c}-\cos\theta_{o}]\left[\frac{v_{out}}{\rm km~s^{-1}}\right]$ (37)
-$\frac{\dot{p}}{\rm dyne}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=10^{-3.2}\left[\frac{v_{out}}{\rm km~s^{-1}}\right]\left[\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\right]$ (38)
-$\frac{\dot{E}}{\rm erg~s^{-1}}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=10^{-6.5}\left[\frac{v_{out}}{\rm km~s^{-1}}\right]^{2}\left[\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\right]$ (39)
-Typically, $n_{0}R_{0}^{2}$ is of order unity for the case studies in §4, with values ranging from about 1 to 5 cm-3 kpc2, although for Mrk 462, Flury2023 find $n_{0}R_{0}^{2}=12$ cm-3 kpc2.
+$\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=0.191\mathcal{R}[\cos\theta_{c}-\cos\theta_{o}]\left[\frac{v_{\rm out}}{\rm km~s^{-1}}\right]$ (41)
+$\frac{\dot{p}}{\rm dyne}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=10^{-3.2}\left[\frac{v_{\rm out}}{\rm km~s^{-1}}\right]\left[\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\right]$ (42)
+$\frac{\dot{E}}{\rm erg~s^{-1}}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=10^{-6.5}\left[\frac{v_{\rm out}}{\rm km~s^{-1}}\right]^{2}\left[\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\right]$ (43)
+Typically, $n_{0}R_{0}^{2}$ is of order unity for the case studies in §4, with values ranging from about 1 to 5 cm-3 kpc2, although for Mrk 462, Flury et al. (2023) find $n_{0}R_{0}^{2}=12$ cm-3 kpc2.
 3 The OutLines Code
@@
 Below, we assess OutLines in the contexts of standard techniques, including empirical percentiles and multi-Gaussian fitting.
-Details of the architecture and use of the OutLines code can be found in Appendix C.
+Details of the architecture and use of the OutLines code can be found in Appendix B with additional examples available in Jupyter notebooks online at https://github.com/sflury/OutLines.
 3.1 Empirical Percentiles
 […]
-Figure 14: Velocity percentiles (black dashed lines) for an arbitrary emission (top) and absorption (bottom) line (purple) generated by OutLines for a directed, filled cone geometry (Geometry=’FilledCones’), default parameters, and otherwise default settings.
+Figure 15: Velocity percentiles (black dashed lines) for an arbitrary emission (top) and absorption (bottom) line (purple) generated by OutLines for a directed, filled cone geometry (Geometry=’FilledCones’) for $i=45^{\circ}$ with $v_{0}=0.1v_{\infty}$ (FromRest=False), default parameters, and otherwise default settings.
 The cumulative distribution function (fraction of integrated flux) is shown in red.
-For reference, we show the characteristic outflow velocity $v_{out}$ (Equation 32) in blue.
+For reference, we show the characteristic outflow velocity $v_{\rm out}$ (Equation 36) in blue.
 3.2 Modeling Line Profiles
@@
 Users are able to update any or all of these bound methods as needed.
-An example of modeling with OutLines is provided in Appendix E.
+An example of modeling with OutLines is provided in Appendix D.
 While we encourage simultaneous fitting of multiple features from the same or very similar species, we caution against modeling lines from different species which may not necessarily share, say, the same density profile parameters due to ionization structure within a multiphase outflow.
@@
 The double Gaussian model over-predicts the outflow flux by a factor of two and under-predicts the static gas flux by a factor of three while characterizing the outflow as red-shifted relative to the systemic velocity of the static gas.
-The double Gaussian model results lead to vastly different physical interpretations of the observations than those from OutLines: one would conclude the presence of a more massive outflow that is launched from a location traveling away from the static gas relative to the observer at $100$ $\rm km~s^{-1}$ but at characteristic velocities of $v_{out}=100$ $\rm km~s^{-1}$.
+The double Gaussian model results lead to vastly different physical interpretations of the observations than those from OutLines: one would conclude the presence of a more massive outflow that is systematically traveling away from the static gas at $15$ $\rm km~s^{-1}$, perhaps merger-like activity.
 One might even be inclined to interpret the double-Gaussian result as evidence for an inflow rather than an outflow, which is not the case.
-The OutLines best-fit model indicates an outflow launched from the same systemic velocity as the static gas with a bidirectional cone geometry and with characteristic velocities of $v_{out}=220$ $\rm km~s^{-1}$.
-Following Equation 34, accounting for the geometry (a factor of $1/(1-\cos\pi/6)\approx 7.5$), differences in line flux ($n_{0}\propto\sqrt{F_{0}}$ gives a factor $\sqrt{2}$), and difference in $v_{out}$ (factor of 0.5), the double Gaussian model yields a mass loss rate higher than that of OutLines by a factor of five.
+The OutLines best-fit model indicates an outflow launched from the same systemic velocity as the static gas with a bidirectional cone geometry and with characteristic velocities of $v_{\rm out}=220$ $\rm km~s^{-1}$.
+Following Equation 38, accounting for the geometry (a factor of $1/(1-\cos\pi/6)\approx 7.5$), differences in line flux ($n_{0}\propto\sqrt{F_{0}}$ gives a factor $\sqrt{2}$), and difference in $v_{\rm out}$ (factor of 0.5), the double Gaussian model yields a mass loss rate higher than that of OutLines by a factor of five.
 Thus, while the velocity inferred from the double Gaussian model is half that from OutLines, the double Gaussian model implies a substantially higher impact of the outflow on its environment than the OutLines model.
@@
 Absorption lines are more sensitive to the line-of-sight, as evidenced in the line profiles (Figures 2 and 4) and the narrow range in velocities contributing to each wavelength bin due to the “down-the-barrel” effect (Figures 7 and 9), making them good probes of the density profile and velocity field but poor tracers of the geometry.
-Emission lines trace the global gas distribution since velocities from across the majority of the gas contribute to each velocity bin (Figure 9), making them good probes of the geometry (Figures 11, 12, 13) but, as a result, lose some constraining power on the density profile and velocity field.
+Emission lines trace the global gas distribution since velocities from across the majority of the gas contribute to each velocity bin (Figure 9), making them good probes of the geometry (Figures 11, 12, 14) but, as a result, lose some constraining power on the density profile and velocity field.
 As such, combined modeling of both emission and absorption provides optimal constraints on the outflow properties.
@@
 The prior is assumed to be uniform bounded unless otherwise stated.
-We show posterior sampling for each case in Appendix F.
+We show posterior sampling for each case in Appendix E.
 Table 1: Best-fit parameters and properties of test-case objects: Knot A of giant extragalactic H ii region NGC 5471, super star cluster in the center of Green Pea J1044+3053, star-forming galaxy Mrk 1486, and Seyfert 2 NGC 2992.
-As discussed in §2.5, reported properties of $\dot{M}$, $\dot{p}$, and $\dot{E}$ are normalized to $n_{0}R_{0}^{2}$ with $n_{0}$ in units of cm-3 and $R_{0}$ in kpc (see Equations 37-39).
+As discussed in §2.5, reported properties of $\dot{M}$, $\dot{p}$, and $\dot{E}$ are normalized to $n_{0}R_{0}^{2}$ with $n_{0}$ in units of cm-3 and $R_{0}$ in kpc (see Equations 41-43).
 Object NGC 5471 Knot A J1044+0353 SSC Mrk 1486 NGC 2992 Sy2 nucl.
@@
 […]
-OutLines model BetaCAK, PowerLaw, Sphere BetaCAK, PowerLaw, Sphere BetaCAK, DampedPulses, Sphere AccPlaw, PowerLaw, HollowCones, Disk
+OutLines model BetaCAK, PowerLaw, Sphere BetaCAK, PowerLaw, Sphere, Launched BetaCAK, DampedPulses, Sphere AccPlaw, PowerLaw, HollowCones, Disk
 Parameters from MCMC Posterior Samples
-$\alpha$ $0.98$  $\pm$ $0.31$ $1.39$  $\pm$ $0.76$ $1.42$  $\pm$ $0.37$
-$\beta$ $4.07$  $\pm$ $1.24$ $2.27$  $\pm$ $1.10$ $0.71$  $\pm$ $0.25$ $2.48$  $\pm$ $0.42$
-$v_{\infty}$ $\rm km~s^{-1}$] $396$  $\pm$ $75$ $213$  $\pm$ $26$ $624$  $\pm$ $38$ $1034$  $\pm$ $108$
-$b$ or $\sigma_{v}$ $\rm km~s^{-1}$] $28$  $\pm$ 11 $34$  $\pm$ 2020 $50$  $\pm$ 2020 $77$  $\pm$ 11
+$\alpha$ $0.23$  $\pm$ $0.05$ $0.58$  $\pm$ $0.50$ $0.03$  $\pm$ $0.06$
+$\beta$ 77  $\pm$ 2 $1.87$  $\pm$ $1.09$ $0.75$  $\pm$ $0.27$ $2.02$  $\pm$ $0.17$
+$v_{\infty}$ $\rm km~s^{-1}$] $309$  $\pm$ $14$ $195$  $\pm$ $17$ $567$  $\pm$ $42$ $925$  $\pm$ $65$
+$b$ or $\sigma_{v}$ $\rm km~s^{-1}$] $30$  $\pm$ $0.5$ $5.0$  $\pm$ $2.4$ $64$  $\pm$ 2020 $82$  $\pm$ 11
 Properties from MCMC Posterior Samples
-$W_{50}$ (“FWHF”) $\rm km~s^{-1}$] $49$  $\pm$ 11 $87$  $\pm$ 44 $181$  $\pm$ $27$ $127$  $\pm$ 44
-$W_{80}$ $\rm km~s^{-1}$] $109$  $\pm$ 33 $147$  $\pm$ 66 $305$  $\pm$ $77$ $289$  $\pm$ $12$
-$x_{out}$ $5.16$  $\pm$ $2.04$ $2.63$  $\pm$ $^{1.78}_{1.09}$ $1.24$  $\pm$ $0.17$ $1.35$  $\pm$ $0.02$
-$v_{out}$ $165$  $\pm$ $15$ $72$  $\pm$ $17$ $188$  $\pm$ $78$ $597$  $\pm$ $267$
-$\dot{M}/n_{0}R_{0}^{2}$ $\rm M_{\odot}~yr^{-1}$] $105$  $\pm$ $59$ $22.1$  $\pm$ $^{33.6}_{14.2}$ $9.8$  $\pm$ $3.2$ $9.5$  $\pm$ $0.4$
-$\dot{p}/n_{0}R_{0}^{2}$ $\rm 10^{34}~dyne$] $10.9$  $\pm$ $7.2$ $1.00$  $\pm$ $^{2.21}_{0.71}$ $0.95$  $\pm$ $0.54$ $3.59$  $\pm$ $0.25$
-$\dot{E}/n_{0}R_{0}^{2}$ $\rm 10^{42}~erg~s^{-1}$] $0.90$  $\pm$ $0.69$ $0.04$  $\pm$ $^{0.11}_{0.03}$ $0.09$  $\pm$ $0.07$ $1.07$  $\pm$ $0.11$
+$W_{50}$ (“FWHF”) $\rm km~s^{-1}$] $47$  $\pm$ 11 $88$  $\pm$ 44 $184$  $\pm$ $27$ $235$  $\pm$ $28$
+$W_{80}$ $\rm km~s^{-1}$] $105$  $\pm$ 22 $150$  $\pm$ 66 $318$  $\pm$ $77$ $541$  $\pm$ $35$
+$x_{\rm out}$ $43$  $\pm$ $23$ $4.22$  $\pm$ $^{11.83}_{~~2.52}$ $1.24$  $\pm$ $0.17$ 2020  $\pm$ $0.02$
+$v_{\rm out}$ $246$  $\pm$ 66 $118$  $\pm$ $32$ $188$  $\pm$ $78$ $883$  $\pm$ $46$
+$\dot{M}/n_{0}R_{0}^{2}$ $\rm M_{\odot}~yr^{-1}$] $777$  $\pm$ $295$ $44.5$  $\pm$ $^{268.0}_{~~32.7}$ $4.8$  $\pm$ $3.2$ $721$  $\pm$ $321$
+$\dot{p}/n_{0}R_{0}^{2}$ $\rm 10^{34}~dyne$] $120$  $\pm$ $47$ $3.21$  $\pm$ $^{27.00}_{~~2.67}$ $0.57$  $\pm$ $0.54$ $430$  $\pm$ $168$
+$\dot{E}/n_{0}R_{0}^{2}$ $\rm 10^{42}~erg~s^{-1}$] $15$  $\pm$ 66 $0.19$  $\pm$ $^{0.16}_{2.09}$ $0.05$  $\pm$ $0.05$ $195$  $\pm$ $70$
 4.1 Evolution of H ii Regions
@@
 Winds launched from stars carry out recently formed metals dredged up to the surface and subsequently sweep up surrounding material.
-Radiation heats gas, often through photoionization, and imparts momentum to the ISM, which can result in additional launching or driving of winds, bubbles, and outflowing gas (e.g., Castor1975b; Weaver1977; Dyson1977; Dyson1979; Oey1997; Dopita2005).
-While stellar population models often contain predictions for these feedback effects (e.g., 1999ApJS..123....3L; starburst99; Eldrige2017; Byrne2022; Hawcroft2025), empirically constraining the kinematics and related mechanical feedback in H ii regions remains an active area of research (e.g., models and simulations by Geen2020; Geen2023; Lancaster2025a; Lancaster2025b) with implications for understanding chemical enrichment (e.g., Esteban1995; LopezSanchez2007; Stock2011) and dispersal timescales (e.g., Dopita2005; Menon2024).
+Radiation heats gas, often through photoionization, and imparts momentum to the ISM, which can result in additional launching or driving of winds, bubbles, and outflowing gas (e.g., Castor et al., 1975b; Weaver et al., 1977; Dyson, 1977, 1979; Oey and Clarke, 1997; Dopita et al., 2005).
+While stellar population models often contain predictions for these feedback effects (e.g., Leitherer et al., 1999, 2014; Eldridge et al., 2017; Byrne et al., 2022; Hawcroft et al., 2025), empirically constraining the kinematics and related mechanical feedback in H ii regions remains an active area of research (e.g., models and simulations by Geen et al., 2020, 2023; Lancaster et al., 2025b, a) with implications for understanding chemical enrichment (e.g., Esteban and Peimbert, 1995; López-Sánchez et al., 2007; Stock et al., 2011) and dispersal timescales (e.g., Dopita et al., 2005; Menon et al., 2024).
 One such kinematically rich and complex system is the giant extragalactic H ii region NGC 5471 located in M101.
-Many knots within this region exhibit broad and sometimes asymmetric nebular lines which cannot be described by a single Gaussian (e.g., Chu1986; Bresolin2020).
-To investigate the evolution of H ii regions in this case-study, we apply OutLines to the [S ii] $\lambda\lambda 6716,6731$ doublet, which traces the density and kinematics of low ionization gas, in Knot A of NGC 5471, the brightest knot identified by Skillman1985, as observed with the $R=6000$ MEGARA IFU spectrograph on the Gran Telescopia Canarias (Arellano-Cordova in prep, PI Esteban).
+Many knots within this region exhibit broad and sometimes asymmetric nebular lines which cannot be described by a single Gaussian (e.g., Chu and Kennicutt, 1986; Bresolin et al., 2020).
+To investigate the evolution of H ii regions in this case-study, we apply OutLines to the [S ii] $\lambda\lambda 6716,6731$ doublet, which traces the density and kinematics of low ionization gas, in Knot A of NGC 5471, the brightest knot identified by Skillman (1985), as observed with the $R=6000$ MEGARA IFU spectrograph on the Gran Telescopia Canarias (Arellano-Cordova in prep, PI Esteban).
 For our modeling, we use a 2.5′′ aperture extraction of the Knot A spectrum from the IFU cube.
-We obtain atomic data from NIST (NIST_ASD) for wavelengths (Kaufman1993) and spontaneous emission coefficients (transition probabilities) (Podobedova2009).
-We use collision strengths from Tayal2010.
+We obtain atomic data from NIST (Kramida et al., 2024) for wavelengths (Kaufman and Martin, 1993) and spontaneous emission coefficients (transition probabilities) (Podobedova et al., 2009).
+We use collision strengths from Tayal and Zatsarinny (2010).
 First, we fit the adjacent featureless continuum with a third-order polynomial, an appropriate approximation in the case of extended H ii regions which do not suffer appreciably from contamination such as stellar absorption except near permitted lines from, e.g., H i and He i.
@@
 Bottom: residual difference between the best-fit OutLines profile and the observed spectrum.
-From the best-fit line fluxes, we determine the electron density of both the nebula and outflow components from the [S ii] 6716/6731{6716}/{6731} flux ratio using PyNeb (2015A&A...573A..42L).
+From the best-fit line fluxes, we determine the electron density of both the nebula and outflow components from the [S ii] 6716/6731{6716}/{6731} flux ratio using PyNeb (Luridiana et al., 2015).
 The static gas flux ratio of 1.459 is in the low-density limit, requiring $n_{e}<10\rm\ cm^{-3}$.
-For the outflowing gas component, the doublet ratio is 0.831, which is where [S ii] is most sensitive simultaneous to density and to temperature.
+For the outflowing gas component, the doublet ratio is 0.831, which is where [S ii] is most sensitive simultaneously to density and to temperature.
 For a range in $T_{e}$ from 10 to 15 $\rm kK$, the inferred outflow density is anywhere from $970$ to $1140\rm~cm^{-3}$, with density values increasing with the assumed temperature.
 The implication from these densities is that the outflow is roughly two orders of magnitude more dense than the gas within the static nebular gas.
-While this increase in density may be due to a sweeping up of material in the wind, this dramatic change in density coupled with the supersonic flow could instead arise from compression by shocked gas as it rapidly cools via radiation (e.g., Flury2025b).
-The slope of the velocity field, $\beta=4$, corresponds to gradual acceleration likely caused by optically thick radiation pressure from early type stars (e.g., 1977ApJ...213..737B).
-While Bresolin2020 argue that the non-thermal radio SEDs of Sramek1986 are clear-cut evidence for supernova-driven bubbles, the line profile shape indicates a continuous wind rather than shells or bubbles and, from modeling with OutLines, disfavors the supernova scenario.
-Indeed, the OutLines results are very much in agreement with the fact that Knot A features the a very young ($3\rm~Myr$) cluster–in fact the youngest in the entirety of NGC 5471 (e.g., GarciaBenito2011, based on HST photometric color-magnitude diagrams).
-Knot A is also very massive ($10^{4}\rm~M_{\odot}$) and requires a significant population of O stars to produce the $Q=4\times 10^{51}\rm~phot~s^{-1}$ associated with the star cluster (GarciaBenito2011).
+While this increase in density may be due to a sweeping up of material in the wind, this dramatic change in density coupled with the supersonic flow implied by $v_{\rm out}$ could instead arise from compression by shocked gas as it rapidly cools via radiation (e.g., Flury et al., 2025a).
+Though the posterior on $\beta$ is not well constrained, the slope of the velocity field strongly favors $\beta>4$, indicating gradual acceleration likely caused by optically thick radiation pressure from early type stars (e.g., Barlow and Cohen, 1977).
+While Bresolin et al.
+(2020) argue that the non-thermal radio SEDs of Sramek and Weedman (1986) are clear-cut evidence for supernova-driven bubbles, the line profile shape indicates a continuous wind rather than shells or bubbles and, from modeling with OutLines, disfavors the supernova scenario.
+Indeed, the OutLines results are very much in agreement with the fact that Knot A features the a very young ($3\rm~Myr$) cluster–in fact the youngest in the entirety of NGC 5471 (e.g., García-Benito et al., 2011, based on HST photometric color-magnitude diagrams).
+Knot A is also very massive ($10^{4}\rm~M_{\odot}$) and requires a significant population of O stars to produce the $Q=4\times 10^{51}\rm~phot~s^{-1}$ associated with the star cluster (García-Benito et al., 2011).
 The age and mass of the stellar population indicate sufficient radiation available in the UV to radiatively drive a supersonic wind.
 Shocks triggered by this wind could explain the non-thermal emission in the radio in addition to the dense gas.
-Inhomogeneities in the gas distribution can readily dissipate the outward force due to pressure from the hot shocked gas (e.g., Dyson1977), allowing radiation pressure to dominate the acceleration of material out from the H ii region.
-Our results indicating escape of hot gas are consistent with the HarperClark2009 model for the Carina Nebula and the Lopez2011 multiwavelength study of the massive extragalactic H ii region 30 Doradus, both of which found strong evidence for radiation pressure driving and a deficit of X-ray flux associated with shocked gas.
+Inhomogeneities in the gas distribution can readily dissipate the outward force due to pressure from the hot shocked gas (e.g., Dyson, 1977), allowing radiation pressure to dominate the acceleration of material out from the H ii region.
+Our results indicating escape of hot gas are consistent with the Harper-Clark and Murray (2009) model for the Carina Nebula and the Lopez et al.
+(2011) multiwavelength study of the massive extragalactic H ii region 30 Doradus, both of which found strong evidence for radiation pressure driving and a deficit of X-ray flux associated with shocked gas.
 With the velocity field and line fluxes inferred by modeling nebular lines with OutLines, we have obtained key insight into the nature of the outflowing material: (i) that it is a wind, not a bubble, (ii) that the wind exhibits gas densities far higher than that of the static nebula, and (iii) that the wind is driven by substantial radiation pressure from O stars.
@@
 4.2 Lyman Continuum Escape from a Super Star Cluster
-Starburst galaxy J1044+0353 is a low mass ($M_{\star}\approx 10^{6.8}\rm~M_{\odot}$) Lyman continuum emitter candidate with $f_{esc}^{\rm Lyc}$ as high as 60% (Parker2025) and, as a nearby ($z=0.01287$) Green Pea, is a local analog for star-forming galaxies at cosmic dawn that are actively involved in re-ionizing the Universe (Martin2024).
-Recent IFU observations of J1044+0353 reveal expanding shells of gas at the periphery of the nebular gas flux distribution (Martin2024), indicative of blowout due to SNe from an aging stellar population 10-20 Myr after a burst of star formation (Peng2023).
-However, the central core of the galaxy hosts a super star cluster with several young (<2<2 Myr) stellar populations which may be responsible for driving additional feedback on shorter timescales (Olivier2022), providing a unique laboratory to explore the impact of young stars on their immediate environment in the context of stochastic, or bursty, star formation using OutLines.
-To undertake this investigation, we draw on archival HST/COS G130M FUV spectroscopy (R=10  [^0] R=10\ 000) of the central starburst of J1044+0353 from the COS LegAcy Spectroscopic SurveY (CLASSY, CLASSY; CLASSY2, GO 15840), which allows for detailed modeling of outflow properties along the line of sight without significant effects from the LSF even after downsampling to the 0.073 Å resolution (see CLASSY; CLASSY2; Parker2024).
+Starburst galaxy J1044+0353 is a low mass ($M_{\star}\approx 10^{6.8}\rm~M_{\odot}$) Lyman continuum emitter candidate with $f_{esc}^{\rm Lyc}$ as high as 60% (Parker et al., 2025) and, as a nearby ($z=0.01287$) Green Pea, is a local analog for star-forming galaxies at cosmic dawn that are actively involved in re-ionizing the Universe (Martin et al., 2024).
+Recent IFU observations of J1044+0353 reveal expanding shells of gas at the periphery of the nebular gas flux distribution (Martin et al., 2024), indicative of blowout due to SNe from an aging stellar population 10-20 Myr after a burst of star formation (Peng et al., 2023).
+However, the central core of the galaxy hosts a super star cluster with several young (<2<2 Myr) stellar populations which may be responsible for driving additional feedback on shorter timescales (Olivier et al., 2022), providing a unique laboratory to explore the impact of young stars on their immediate environment in the context of stochastic, or bursty, star formation using OutLines.
+To undertake this investigation, we draw on archival HST/COS G130M FUV spectroscopy (R=10  [^0] R=10\ 000) of the central starburst of J1044+0353 from the COS LegAcy Spectroscopic SurveY (CLASSY, Berg et al., 2022; James et al., 2022, GO 15840), which allows for detailed modeling of outflow properties along the line of sight without significant effects from the LSF even after downsampling to the 0.073 Å resolution (see Berg et al., 2022; James et al., 2022; Parker et al., 2024).
 The C ii $\lambda 1334$ absorption line is one of the strongest low ionization lines in the FUV due to the transition oscillator strength and comparatively high ionic abundance, making it an excellent tracer of outflows.
-Visual inspection of the C ii line profile suggests a classic blue-shifted continuous outflow plus a static ISM component (see, e.g., Figure 1 of 2015ApJ...809..147H).
-To obtain the normalized flux, we fit the starlight continuum within 3 Å of the feature with a third-order polynomial and divide the local spectrum by this fit (see, e.g., the VoigtFit software, VoigtFit, which takes a similar approach using Chebyshev polynomials).
-For modeling the C ii line with OutLines, we use atomic data from Morton2003.
+Visual inspection of the C ii line profile suggests a classic blue-shifted continuous outflow plus a static ISM component (see, e.g., Figure 1 of Heckman et al., 2015).
+To obtain the normalized flux, we fit the starlight continuum within 3 Å of the feature with a third-order polynomial and divide the local spectrum by this fit (see, e.g., the VoigtFit software, Krogager, 2018, which takes a similar approach using Chebyshev polynomials).
+For modeling the C ii line with OutLines, we use atomic data from Morton (2003).
 No infilling is apparent in the observed line profile, suggesting either maximal aperture clipping where pure absorption is an appropriate description of the emergent feature (see Figure 8) or an optically thick wind.
-For simplicity, we assume the OutLines defaults, although we note that applying more sophisticated geometry or density profiles negligibly impacts the quality of the fit.
+For simplicity, we assume the OutLines default geometry and density profile, although we note that applying more sophisticated geometry or density profiles negligibly impacts the quality of the fit.
 Indeed, a filled cone model yields best-fit inclinations of $i\to 0^{\circ}$ and opening angles of $\theta_{o}\to 90^{\circ}$, consistent with the filled sphere geometry.
 Additional consideration of pulse-like density profiles consistently favors a peak location at the base of the outflow with wide dispersion, which combined are consistent with the power law approximation.
+However, the sudden drop in flux immediately blueward of the systemic velocity of the system suggests a non-zero launch velocity.
+We therefore allow for non-zero launch velocities.
+The OutLines default top hat (uniform-bounded) prior is assumed.
 Best-fit parameters are in Table 1.
 […]
-Figure 17: MCMC fit of an OutLines.Absorption model to the C ii $\lambda 1334$ absorption line observed by HST/COS G130M (black, CLASSY).
+Figure 18: MCMC fit of an OutLines.Absorption model to the C ii $\lambda 1334$ absorption line observed by HST/COS G130M (black, Berg et al.
+2022).
 Line styles and colors as in Figure 17.
 Blue line indicates the intrinsic profile convolved with the HST/COS G130M line spread function (LSF).
-The best-fit parameters indicate that the starbursting core of J1044+0353 contains a spherically expanding, isothermal wind with a velocity field characteristic of optically thick, radiatively-driven winds (e.g., 1977ApJ...213..737B; Flury2023).
-However, the relatively low wind velocity ($v_{\infty}=215$ $\rm km~s^{-1}$, comparable to Voigt modeling of the silicon lines, Parker2024, and the “broad” component of nebular lines, Peng2023) and weak $\dot{p}$ and $\dot{E}$ (marginally non-zero) suggests that mechanical feedback from massive stars is relatively weak or even suppressed.
-Such suppression can occur in cases where catastrophic cooling of swept up dense material prevents acceleration up to more “traditional” velocities closer to 500 $\rm km~s^{-1}$ or more, is known to occur in Green Pea galaxies with young stellar populations (2017ApJ...851L...9J), and could be associated with Lyman continuum (LyC) escape (e.g., 2017ApJ...851L...9J; Jaskot2019).
-Combined with evidence for recent SNe, the weak feedback inferred from OutLines models supports the bursty star formation scenario for LyC escape suggested by Flury2022b in their assessment of a large sample of star-forming galaxies with LyC detections and later demonstrated by Flury2025a with measurement of stellar populations and ISM conditions in stacks of UV spectra of LyC emitters.
+The best-fit parameters indicate that the starbursting core of J1044+0353 contains a spherical expanding wind with a velocity field characteristic of radiatively-driven winds most likely due to dusty radiation pressure (e.g., Flury et al., 2023), although line-driving (Castor et al., 1975a) and optically thick (Barlow and Cohen, 1977) mechanisms are consistent with the range of $\beta$.
+Critically, the posterior on $\beta$ strongly disfavors ram pressure driving, suggesting that supernovae do not cause the current feedback.
+However, the relatively low wind velocity ($v_{\infty}=215$ $\rm km~s^{-1}$, comparable to Voigt modeling of the silicon lines, Parker et al.
+2024, and the “broad” component of nebular lines, Peng et al.
+2023) and weak $\dot{p}$ and $\dot{E}$ (marginally non-zero) suggests that mechanical feedback from massive stars is relatively weak or even suppressed.
+Such suppression can occur in cases where catastrophic cooling of swept up dense material prevents acceleration up to more “traditional” velocities closer to 500 $\rm km~s^{-1}$ or more, is known to occur in Green Pea galaxies with young stellar populations (Jaskot et al., 2017), and could be associated with Lyman continuum (LyC) escape (e.g., Jaskot et al., 2017, 2019).
+Combined with evidence for recent SNe, the weak feedback inferred from OutLines models supports the bursty star formation scenario for LyC escape suggested by Flury et al.
+(2022) in their assessment of a large sample of star-forming galaxies with LyC detections and later demonstrated by Flury et al.
+(2025b) with measurement of stellar populations and ISM conditions in stacks of UV spectra of LyC emitters.
 In their framework, SNe clear out channels in the ISM via mechanical feedback and a subsequent generation of stars 8-10 Myr later provides ionizing–but not mechanical–feedback to establish the density-bounded conditions necessary for Lyman continuum photons to escape anisotropically.
@@
 4.3 Outflow Formation and Blowout in a Starburst Galaxy
-Mrk 1486 is a starburst galaxy exhibiting a tentative He ii $\lambda 4686$ Wolf-Rayet feature (e.g., Izotov1997), metal rich extraplanar biconical outflows beyond the galactic plane (e.g., Duval2016; Cameron2021), and >14>14 Myr stellar populations (Parker2025, although see Peng2025, who find from the same UV spectra that these populations are even older, $\approx 100$ Myr).
-The archival HST/COS spectrum from RiveraThorsen2015 exhibits the classic blue-shifted asymmetric feature associated with outflows, albeit with at least two troughs present in the outflow feature across several low ionization lines (Si ii $\lambda\lambda 1191,1193,1304$, O i $\lambda 1302$, and Al ii $\lambda 1670$, see RiveraThorsen2015 their Figure 3, Parker2024, their Figure 7).
-A multi-troughed absorption feature can indicate a ensemble of shells produced by a series of episodes or a system of many star clusters (see Figure 5, also RiveraThorsen2015).
-For our analysis, we use the RiveraThorsen2015 spectrum reprocessed by CLASSY; CLASSY2, down-sampled to 0.073 Å resolution and normalized following the polynomial approach in §4.2.
+Mrk 1486 is a starburst galaxy exhibiting a tentative He ii $\lambda 4686$ Wolf-Rayet feature (e.g., Izotov et al., 1997), metal rich extraplanar biconical outflows beyond the galactic plane (e.g., Duval et al., 2016; Cameron et al., 2021), and >14>14 Myr stellar populations (Parker et al.
+2025, although see Peng et al.
+2025, who find from the same UV spectra that these populations are even older, $\approx 100$ Myr).
+The archival HST/COS spectrum from Rivera-Thorsen et al.
+(2015, GO 12583, PI Hayes) exhibits the classic blue-shifted asymmetric feature associated with outflows, albeit with at least two troughs present in the outflow feature across several low ionization lines (Si ii $\lambda\lambda 1191,1193,1304$, O i $\lambda 1302$, and Al ii $\lambda 1670$, see Rivera-Thorsen et al.
+2015 their Figure 3, Parker et al., 2024, their Figure 7).
+A multi-troughed absorption feature can indicate a ensemble of shells produced by a series of episodes or a system of many star clusters (see Figure 5, also Rivera-Thorsen et al.
+2015).
+For our analysis, we use the Rivera-Thorsen et al.
+(2015) spectrum reprocessed by Berg et al.
+(2022); James et al.
+(2022), down-sampled to 0.073 Å resolution and normalized following the polynomial approach in §4.2.
 To determine the origin of this multi-episode outflow, we can use OutLines to model the absorption lines.
@@
 We mask these features in order to include the fluorescent lines in our fit.
-Using atomic data from Kelleher2008 obtained from NIST (NIST_ASD) and the appropriate COS G130M LSF, we model the Si ii doublet complex of $\lambda\lambda 1191,1193$ resonant absorption and emission and $\lambda\lambda 1194,1197$ fluorescent emission using OutLines assuming a damped-pulses density profile to account for the multiple absorption troughs and otherwise adopting the default options of a spherical geometry and the $\beta$ CAK velocity field.
+Using atomic data from Kelleher and Podobedova (2008) obtained from NIST (Kramida et al., 2024) and the appropriate COS G130M LSF, we model the Si ii doublet complex of $\lambda\lambda 1191,1193$ resonant absorption and emission and $\lambda\lambda 1194,1197$ fluorescent emission using OutLines assuming a damped-pulses density profile to account for the multiple absorption troughs and otherwise adopting the default options of a spherical geometry and the $\beta$ CAK velocity field.
 The choice of spherical geometry is supported by the presence of relatively symmetric fluorescent Si ii $\lambda 1194,1197$ lines, although the COS LSF may complicate any inference of the geometry.
@@
 $\frac{F_{\lambda}}{F_{\lambda 0}}=1-C_{f}\left[1-\exp(-\tau_{0}\phi_{\lambda})\right]$ (44)
-where $C_{f}$ expresses the fraction of background light which is not obstructed by the foreground gas (e.g., Netzer1982; Arav1999; Arav2005; RiveraThorsen2015).
-The COS LSF can affect the partial covering fractions of saturated absorption lines (see, e.g., assessment by Flury2025a, their Appendix D, also Figure 3 of Jennings2025).
-Incorporating both $C_{f}$ and $\epsilon$ as free parameters into the initial fit, we find that while the COS LSF and infilling by the $1194$ fluorescent line can account for some of the line profile depths, $C_{f}=0.82\pm 0.03$ is necessary to provide an optimal description to the absorption features (comparable to the $C_{f}=0.8$ obtained by RiveraThorsen2015).
+where $C_{f}$ expresses the fraction of background light which is not obstructed by the foreground gas (e.g., Netzer, 1982; Arav et al., 1999, 2005; Rivera-Thorsen et al., 2015).
+The COS LSF can affect the partial covering fractions of saturated absorption lines (see, e.g., assessment by Flury et al., 2025b, their Appendix D, also Figure 3 of Jennings et al.
+2025).
+Incorporating both $C_{f}$ and $\epsilon$ as free parameters into the initial fit, we find that while the COS LSF and infilling by the $1194$ fluorescent line can account for some of the line profile depths, $C_{f}=0.82\pm 0.03$ is necessary to provide an optimal description to the absorption features (comparable to the $C_{f}=0.8$ obtained by Rivera-Thorsen et al., 2015).
 We also find $\epsilon=0.21\pm 0.03$ necessary to describe the fluorescent line profile.
-We adjust the prior to require the column densities of each line component to agree within 0.05 dex standard deviation of the column densities of the same components in the other line, effectively enforcing the atomic data.
-Based on visual assessment of the line profiles, we further impose a Gaussian prior requiring the terminal velocity to agree with the apparent velocity at which the blue profile wing reaches the continuum ($\approx 400$ $\rm km~s^{-1}$, accounting for the LSF) and the central velocity of the first pulse to agree with the velocity of the deepest trough ($\approx 130$ $\rm km~s^{-1}$), both to within a standard deviation of four resolution elements.
-We also impose that each shell is relatively thin by requiring the width of each shell to be $15\pm$ 15% of the launch radius.
 We show the best-fit line profile in Figure 19 and list the results in Table 1.
 As a test of our assumption of spherical geometry, we predict the emergent fluorescent Si ii $\lambda\lambda 1194,1197$ lines using OutLines.Fluorescent and the best-fit parameters from the posterior, which we show in Figure 19.
-Figure 18: MCMC fit of an OutLines model to the Si ii $\lambda\lambda 1191,1193,1194,1997$ doublet complex observed by HST/COS G130M (black, RiveraThorsen2015, GO 12583, PI Hayes).
+Figure 19: MCMC fit of an OutLines model to the Si ii $\lambda\lambda 1191,1193,1194,1997$ doublet complex observed by HST/COS G130M (black, Rivera-Thorsen et al.
+2015, GO 12583, PI Hayes).
 Shaded regions indicate suspected Milkyway contamination from Galactic Si iii, Mn $i$, and and O iv.
@@
 Notably, the less pronounced absorption troughs in C ii $\lambda 1334$, which traces a higher ionization state than Si ii, O i, and Al ii, suggests that the ionized phase of the outflow is established more rapidly than the neutral phase and may indicate that the expanding bubbles are entrained in or even driven by a hotter phase of the ISM.
-The accumulation of superbubbles close to the terminal velocity suggests that an outflow forms from these successive events with sufficient speed to exceed the escape velocity ($v_{esc}\approx$100 $\rm km~s^{-1}$as implied by the $M_{\star}=10^{9.3}\rm~M_{\odot}$ from 2018MNRAS.481.1690C).
-The recurrance of many such expanding shells could produce the biconical outflow observed in emission once the superbubbles escape the gravitational potential and blow out from the galactic disk (see spatially resolved cases in, e.g., Martin1998; McQuinn2019, and Duval2016 specifically regarding Mrk 1486).
-Moreover, the successive events are removing gas at mass outflow rates of $\dot{M}=7.9\rm~M_{\odot}~yr^{-1}$, which may temporarily rob the galaxy of baryonic material necessary to sustain the production of stars and could explain the 30-40 Myr lapse between bursts of star formation.
-Given that the outflows in Mrk 1486 are oxygen rich relative to their host galaxy (Cameron2021), it is useful to consider whether the outflow component of Si ii is metal-enhanced since, like oxygen, silicon is an $\alpha$ element and is predominantly produced by core-collapse supernovae (CCSNe, e.g., Kobayashi2020).
+The accumulation of superbubbles close to the terminal velocity suggests that an outflow forms from these successive events with sufficient speed to exceed the escape velocity ($v_{esc}\approx$100 $\rm km~s^{-1}$as implied by the $M_{\star}=10^{9.3}\rm~M_{\odot}$ from Chisholm et al.
+2018).
+The recurrance of many such expanding shells could produce the biconical outflow observed in emission once the superbubbles escape the gravitational potential and blow out from the galactic disk (see spatially resolved cases in, e.g., Martin, 1998; McQuinn et al., 2019, and Duval et al.
+2016 specifically regarding Mrk 1486).
+Moreover, the successive events are removing gas at mass outflow rates of $\dot{M}=4.8\rm~M_{\odot}~yr^{-1}$, which may temporarily rob the galaxy of baryonic material necessary to sustain the production of stars and could explain the 30-40 Myr lapse between bursts of star formation.
+Given that the outflows in Mrk 1486 are oxygen rich relative to their host galaxy (Cameron et al., 2021), it is useful to consider whether the outflow component of Si ii is metal-enhanced since, like oxygen, silicon is an $\alpha$ element and is predominantly produced by core-collapse supernovae (CCSNe, e.g., Kobayashi et al., 2020).
 From the best-fit line profiles from OutLines, the outflow $N_{o\ \rm Si^{+}}=10^{15.5}\rm~cm^{-2}$ while the static $N_{s\ \rm Si^{+}}=10^{13.5}\rm~cm^{-2}$.
-Ionization structure may account for some of the difference in Si+; nevertheless, the 2 dex increase in the outflow likely suggests substantial metal loading of the outflow (see, e.g., 2018MNRAS.481.1690C) due to enrichment by recent CCSNe.
-It may be tempting to conclude that hot gas or cosmic rays from these supernovae are driving the shells’ expansion, especially since, for the lowest velocity shell at 169 $\rm km~s^{-1}$, the dynamical time to traverse the inner 500 pc is $\la 4$ Myr, well after the onset of CCSNe for a burst at least 10 Myr old.
-That being said, the age reported by Parker2025 from fits to the UV continuum that Mrk 1486 corresponds to a combination of a substantial young ($\leq 5$ Myr) stellar population and a far older (40 Myr) population.
+Ionization structure may account for some of the difference in Si+; nevertheless, the 2 dex increase in the outflow likely suggests substantial metal loading of the outflow (see, e.g., Chisholm et al., 2018) due to enrichment by recent CCSNe.
+It may be tempting to conclude that hot gas or cosmic rays from these supernovae are driving the shells’ expansion, especially since, for the lowest velocity shell at 188 $\rm km~s^{-1}$, the dynamical time to traverse the inner 500 pc is $\la 2.5$ Myr, well after the onset of CCSNe for a burst at least 10 Myr old.
+That being said, the age reported by Parker et al.
+(2025) from fits to the UV continuum that Mrk 1486 corresponds to a combination of a substantial young ($\leq 5$ Myr) stellar population and a far older (40 Myr) population.
 In light of the more detailed star formation history, while CCSNe may be the progenitors of the material in these expanding shells, they may not necessarily be responsible for their acceleration.
 Here, the shape of the velocity field from OutLines can provide insight into the driving mechanism.
-Indeed, the velocity index of $\beta=0.7\pm 0.2$ disfavors (but cannot fully not reject) the rapid acceleration associated with ram pressure driving ($\beta\la 0.5$, e.g., 2005ApJ...618..569M; 2016MNRAS.463..541C).
-However, these $\alpha$-rich shells are more likely continuously accelerated by optically thin line-driven radiation pressure associated with $\beta=0.8$ (e.g., 1986A&A...164...86P).
-Moreover, the total energy injection rate of $10^{40}\rm~erg~s^{-1}$ for the series of shells is roughly consistent with stellar winds, which have $L_{mech}\approx 10^{39}\rm~erg~s^{-1}$ for a given stellar population (see, e.g., Starburst99 or BPASS, starburst99; Byrne2022, respectively).
+Indeed, the velocity index of $\beta=0.7\pm 0.2$ disfavors (but cannot fully not reject) the rapid acceleration associated with ram pressure driving ($\beta\la 0.5$, e.g., Murray et al., 2005; Chisholm et al., 2016).
+However, these $\alpha$-rich shells are more likely continuously accelerated by optically thin line-driven radiation pressure associated with $\beta=0.8$ (e.g., Pauldrach et al., 1986).
+Moreover, the total energy injection rate of $10^{40}\rm~erg~s^{-1}$ for the series of shells is roughly consistent with stellar winds, which have $L_{mech}\approx 10^{39}\rm~erg~s^{-1}$ for a given stellar population (see, e.g., Starburst99 or BPASS, Leitherer et al., 2014; Byrne et al., 2022, respectively).
 Radiative driving by young stellar populations may also account for the less pronounced C ii troughs via ionizing feedback.
-Furthermore, given that the travel time across the central kpc is $\la 4$ Myr, the >40>40 Myr stellar populations are unlikely to be responsible for these superbubbles.
+Furthermore, given that the travel time from the nucleus to a 1 kpc radius (roughly the COS aperture at the 139.1 Mpc distance of Mrk 1486) is $\la 5$ Myr, the >40>40 Myr stellar populations are unlikely to be responsible for these superbubbles.
 Corroborating evidence for optically thick radiation pressure is the presence of the fluorescent $\lambda 1194,1197$ lines but the absence of resonant line infilling.
-Such a scenario occurs in the optically thick line limit where photon escape primarily occurs through the fluorescent channel (e.g., 2015ApJ...801...43S, their Figures 5-6).
+Such a scenario occurs in the optically thick line limit where photon escape primarily occurs through the fluorescent channel (e.g., Scarlata and Panagia, 2015, their Figures 5-6).
 The bursy star formation in Mrk 1486 could optimize the production of $\alpha$ elements and energetic expanding shells via CCSNe from earlier generations of stars while the most recent stellar populations provide the momentum injection via radiation pressure to form and accelerate optically thick enriched superbubbles, driving them beyond the galactic disk to form an outflow.
@@
 4.4 Thermal Winds in AGN Radio Lobes
-The Seyfert 2 galaxy NGC 2992 has long been known to exhibit conical outflows (Allen1999; Veilleux2001).
-Using archival observations from the Siding Spring Seyfert Spectroscopic Snapshot Survey (S7, S7DR2) made using WiFES (WiFES), we have confirmed broad, asymmetric [O iii] $\lambda\lambda 4959,5007$ lines after summing over all the spaxels to boost signal to noise in the line profile and to mitigate any aperture effects.
+The Seyfert 2 galaxy NGC 2992 has long been known to exhibit conical outflows (Allen et al., 1999; Veilleux et al., 2001).
+Using archival observations from the Siding Spring Seyfert Spectroscopic Snapshot Survey (S7, Thomas et al., 2017) made using WiFES (Dopita et al., 2010), we have confirmed broad, asymmetric [O iii] $\lambda\lambda 4959,5007$ lines after summing over all the spaxels to boost signal to noise in the line profile and to mitigate any aperture effects.
 The low resolution ($R\approx$ 3 000 or FWHM of 105 $\rm km~s^{-1}$) requires corrections for the LSF by convolving with a Gaussian kernel with $\sigma_{\rm LSF}=$45 $\rm km~s^{-1}$; however, we note that the LSF predominantly impacts the narrow Gaussian core.
@@
 The bright intermediate portion of the wings suggests the acceleration power law for the velocity field.
-Visual assessment of the highly asymmetric line profile in the spectrum of NGC 2992 and comparison with models in Figures 11, 12, and 13 also suggests a hollow cone geometry with an obstructing disk.
+Visual assessment of the highly asymmetric line profile in the spectrum of NGC 2992 and comparison with models in Figures 11, 12, and 14 also suggests a hollow cone geometry with an obstructing disk.
 As such, we subsequently model the [O iii] doublet using OutLines assuming an acceleration power law, an open cone geometry, and an obstructing disk.
-We impose the following into the prior: $v_{\infty}$ within a 300 $\rm km~s^{-1}$ standard deviation of the apparent maximum velocity of 1100 $\rm km~s^{-1}$, a static velocity dispersion within 50 $\rm km~s^{-1}$ standard deviation of the apparent dispersion of 110 $\rm km~s^{-1}$, accounting for the LSF, line fluxes within 5% of the [O iii] 5007/4959 flux ratio of 2.98 given by atomic data (Storey2000), and angles within 10 standard deviation of the best guess values from the trust-reflective fit.
-With line profiles predicted by OutLines added to the continuum and subsequently convolved with the WiFES LSF (WiFES), we use emcee to sample the posterior on the outflow parameters, requiring $\theta_{o}-\theta_{c}>5^{\circ}$.
+We impose the following into the prior: line fluxes within 5% of the [O iii] 5007/4959 flux ratio of 2.98 given by atomic data (Storey and Zeippen, 2000) and $\theta_{o}-\theta_{c}>5^{\circ}$.
+With line profiles predicted by OutLines added to the continuum and subsequently convolved with the WiFES LSF (Dopita et al., 2010), we use emcee to sample the posterior on the outflow parameters.
 Due to the large number of parameters, potential degeneracies, and sensitivity to initial conditions, we run the sampler four times longer, for a total of 40 000 samples post burn-in.
-Implementation of our modeling is demonstrated for reference in Appendix D with the best-fit profile shown in Figure 19 and results listed in Table 1.
+Implementation of our modeling is demonstrated for reference in Appendix C with the best-fit profile shown in Figure 20 and results listed in Table 1.
 We note that the posterior sampling ultimately rejects the priors on the angles as the best-guess parameters fell within a local minimum.
-Figure 19: MCMC fit of an OutLines.Nebular model to the [O iii] $\lambda\lambda 4959,5007$ doublet observed in the nucleus of Seyfert galaxy NGC 2992 as part of the Siding Spring Seyfert Spectroscopic Snapshot Survey (S7, S7DR2) DR2, assuming a hollow cone plus disk outflow geometry.
+Figure 20: MCMC fit of an OutLines.Nebular model to the [O iii] $\lambda\lambda 4959,5007$ doublet observed in the nucleus of Seyfert galaxy NGC 2992 as part of the Siding Spring Seyfert Spectroscopic Snapshot Survey (S7, Thomas et al., 2017) DR2, assuming a hollow cone plus disk outflow geometry.
 Line styles and colors as in Figure 17.
@@
 […]
-The slope $\beta=2.48$ indicates, for the acceleration power law velocity field, rapid acceleration to the terminal velocity (see Figure 2), which can be interpreted as thermal (ram pressure) or cosmic ray (magnetic pressure) driving as each can produce similar rapid-acceleration velocity fields based on the shape of the velocity field (see Figure 1, also 2005ApJ...618..569M; 2016MNRAS.463..541C; Flury2023).
-Thermal driving occurs when “cool" ($10^{4}\rm~K$) clouds are embedded in an adiabatically expanding hot ($10^{6}\rm~K$) gas (2005ApJ...618..569M).
-Cosmic ray occurs driving where clouds are accelerated by the diffusion (Breitschwerdt1993; Quataert2022a) or streaming (Ipavich1975; Breitschwerdt1991; Quataert2022b) of relativistic charged particles–launched by hot ionized media such as shocks or supernovae (Colgate1979; Blandford1980)–as they interact with magnetic fields in the ISM, resulting in a radially outward pressure gradient.
-NGC 2992 is known to contain $\approx 0.5$ kpc radio lobes forming an hourglass or “figure eight” shape, indicative of bi-directional bubbles of hot gas driven by nuclear activity (Wehrle1988; Allen1999).
-Combining optical and radio measurements indicates the presence of hot cavities which are heated (likely through shocks) by jets or accretion disk winds and within which are entrained clouds of cooler gas (Allen1999; Veilleux2001).
-The presence of such lobes and the associated hollow cavities suggests either or both (a plausible scenario, e.g., Breitschwerdt2002; Everett2008) mechanisms are involved in launching the outflow in NGC 2992.
-Spatial analysis of the outflow in Veilleux2001 found an inclination of $i=68\pm 3^{\circ}$ and an opening angle of $\theta_{o}=62.5-67.5^{\circ}$.
+The slope $\beta=2.02$ indicates, for the acceleration power law velocity field, rapid acceleration to the terminal velocity (see Figure 2), which can be interpreted as thermal (ram pressure) or cosmic ray (magnetic pressure) driving as each can produce similar rapid-acceleration velocity fields based on the shape of the velocity field (see Figure 1, also Murray et al., 2005; Chisholm et al., 2016; Flury et al., 2023).
+Thermal driving occurs when “cool" ($10^{4}\rm~K$) clouds are embedded in an adiabatically expanding hot ($10^{6}\rm~K$) gas (Murray et al., 2005).
+Cosmic ray occurs driving where clouds are accelerated by the diffusion (Breitschwerdt et al., 1993; Quataert et al., 2022b) or streaming (Ipavich, 1975; Breitschwerdt et al., 1991; Quataert et al., 2022a) of relativistic charged particles–launched by hot ionized media such as shocks or supernovae (Colgate and Petschek, 1979; Blandford and Ostriker, 1980)–as they interact with magnetic fields in the ISM, resulting in a radially outward pressure gradient.
+NGC 2992 is known to contain $\approx 0.5$ kpc radio lobes forming an hourglass or “figure eight” shape, indicative of bi-directional bubbles of hot gas driven by nuclear activity (Wehrle and Morris, 1988; Allen et al., 1999).
+Combining optical and radio measurements indicates the presence of hot cavities which are heated (likely through shocks) by jets or accretion disk winds and within which are entrained clouds of cooler gas (Allen et al., 1999; Veilleux et al., 2001).
+The presence of such lobes and the associated hollow cavities suggests either or both (a plausible scenario, e.g., Breitschwerdt et al., 2002; Everett et al., 2008) mechanisms are involved in launching the outflow in NGC 2992.
+Spatial analysis of the outflow in Veilleux et al.
+(2001) found an inclination of $i=68\pm 3^{\circ}$ and an opening angle of $\theta_{o}=62.5-67.5^{\circ}$.
 OutLines finds $i=53.19\pm 1.32^{\circ}$ and $\theta_{o}=68.60\pm 1.26^{\circ}$.
@@
 Not only have we accurately recovered the spatial flux distribution with OutLines, but we have also inferred the presence of a cavity comprising the inner two thirds of the outflow cone ($\theta_{c}=48.64\pm 1.82^{\circ}$), suggesting that cooler line-emitting clouds can only survive towards the edges of the hot cavity where the thermal wind may be somewhat cooler as it expands and contacts the surrounding ISM.
-Figure 20: Greyscale map across the central kiloparsec of NGC 2992 of the continuum-subtracted, integrated [O iii] doublet flux in the S7 DR2 (S7DR2).
+Figure 21: Greyscale map across the central kiloparsec of NGC 2992 of the continuum-subtracted, integrated [O iii] doublet flux in the S7 DR2 (Thomas et al., 2017).
 Drawn cones indicate the geometry inferred from the best-fit OutLines profile to the integrated emission line profile.
@@
 With the best-fit parameters, OutLines computes high mass loss rates of at least 40 $\rm M_{\odot}~yr^{-1}$.
-From the $v_{0}\sim 300$ $\rm km~s^{-1}$ rotational velocity of the galaxy (Veilleux2001), the outflow must be capable of ejecting material from the host galaxy with the characteristic velocity exceeding the escape velocity by a factor of 3.
+From the $v_{0}\sim 300$ $\rm km~s^{-1}$ rotational velocity of the galaxy (Veilleux et al., 2001), the outflow must be capable of ejecting material from the host galaxy with the characteristic velocity exceeding the escape velocity by a factor of 3.
 Together, these results suggest NGC 2992 is in a state of AGN-triggered inside-out quenching with the combined effects of hot cavities and the evacuation of cool gas away from the plane of the galaxy, both effects preventing future star formation.
 Interestingly, this quenching feedback is not sufficiently columnated to be driven by a jet but instead is most likely due to expanding gas heated by winds launched from the accretion disk.
-While the implications of quenching are clear for galaxy evolution, the association of hot winds and shocks also has implications for the measurement of chemical abundances in AGN (e.g., Dors2021).
+While the implications of quenching are clear for galaxy evolution, the association of hot winds and shocks also has implications for the measurement of chemical abundances in AGN (e.g., Dors et al., 2021).
 Here, inferring underlying physics, geometry, and outflow strength from a single-aperture spectrum demonstrates the power and effectiveness of OutLines in the context of AGN feedback.
@@
   • Young (<4<4 Myr) H ii regions undergoing expansion can exhibit two orders of magnitude in gas compression due to wind-triggered shocks even while optically thick radiation pressure is the main driving mechanism of the bulk motion of the gas.
-Thus, while shocks may dominate the radio continuum, it is radiation pressure from stars which primarily shapes the evolution of H ii regions.
+Thus, while shocks may dominate the radio continuum, it is radiation pressure from stars which primarily shapes the evolution of young H ii regions.
   • Suppression of optically thin radiative feedback occurs in young (<2<2 Myr) super star clusters in dwarf galaxies like the Green Peas indicates that the most recent stellar populations cannot mechanically evacuate optically thin channels for Lyman continuum photons to escape.
@@
 Acknowledgements
-We thank the referee for comments which improved this manuscript.
 We thank the beta testers, including K.
@@
 Arellano Cordova, K.
-Duncan, and T.
+Duncan, A.
+Hall, and T.
 Horn, whose efforts helped to optimize the use of this code.
@@
 We thank K.
-Parker for sharing results from the recently accepted Parker2025.
+Parker for sharing results from the recently accepted Parker et al.
+(2025).
 We thank K.
-Duncan for insight regarding organization of this manuscript.
+Duncan for insight and comments regarding this manuscript.
+We thank A.
+Rankine and A.
+Lawrence for suggestions regarding the relativistic corrections.
 We also thank M.
@@
 Facilities: HST, GTC, SSO
-Software, repositories, and other digital resources: corner (corner), emcee (emcee), matplotlib (matplotlib), MAST, NIST/ASD (NIST_ASD), numpy (numpy), OutLines (this work, see also 10.5281/zenodo.11238265), PyNeb (2015A&A...573A..42L), S7 DR2 (S7DR2), scipy (scipy), vygrboi
+Software, repositories, and other digital resources: corner (Foreman-Mackey, 2016), curlyBrace, emcee (Foreman-Mackey et al., 2013), matplotlib (Hunter, 2007), MAST, NIST/ASD (Kramida et al., 2024), numpy (van der Walt et al., 2011), OutLines (this work, see also 10.5281/zenodo.11238265), PyNeb (Luridiana et al., 2015), S7 DR2 (Thomas et al., 2017), scipy (P.
+Virtanen, 2020), vygrboi
 AI use: No language learning model or otherwise generative artificial intelligence was used in this work.
@@
 […]
-A.1 Spherical Cap Projection
-The ellipse is the projected spherical cap with axes $g$ and $h$ and is offset from the projected outflow source by some distance $S$.
+A.1 Cones As A Spherical Cap Projection
+An outflowing cone on the plane of the sky appears as a the projected spherical cap with axes $g$ and $h$ and is offset from the projected outflow source by some distance $S$.
 These quantities can be expressed in terms of the inclination and opening angle of the ouflow center such that
@@
 $h$ $=\cos{i}\sin{\theta}.$ (47)
+We illustrate a single case of such a projection in Figure 22.
 In the case where $i+\theta>90^{\circ}$, the projection is no longer an ellipse but rather an ellipse plus a lune to account for the projection of the “top” of the spherical cap and an additional lune to the opposing side to account for the projection of the second cone.
 We visualize such a scenario in the last two panels of Figure 23.
+Figure 22: A cone geometry in the case of $i=40^{\circ}$, $\theta=30^{\circ}$.
+Left panels illustrate the cones seen edge-on, i.e., transverse to the line of sight.
+Right panels illustrate the cones seen face-on, i.e., projected onto the plane of the sky.
+Solid ellipses indicates the cone projected anterior to source relative to the observe while dashed ellipses indicate the posterior projection.
+Top panels illustrate the geometric terms $S$, $g$, and $h$ which translate the opening angle $\theta$ and inclination $i$ of the cones into a projection onto the sky.
+Bottom panels illustrate how an observed velocity $u$ de-projects into an azimuthal band with radius $r_{w}$ with unique azimuthal coverage $\pi\ell$ within the cone defined by the intersection points $p,q$ of the velocity band with the projected cone.
+High $u/w$ (low $r_{w}$) corresponds to wings of a line profile far from the systemic velocity while low $u/w$ (high $r_{w}$) corresponds to the core of the line profile close to the systemic velocity.
 A.2 Cone Inscription of Deprojected Velocities
 The fractional arc length $\ell_{\rm w}$ of the deprojected velocity inscribed by the cone can be determined by solving the system of linear equations for a circle and the projected spherical cap area to determine where they intersect, which in turn determine the bounds on the integral(s) over the azimuthal coordinate of the solid angle.
-The azimuthal “ring” of gas with velocity $w$ is defined by the sine of the angle between the observed and deprojected velocities $u$ and $w$, respectively, such that its radius is
-$r_{\rm w}=\sqrt{1-(u/w)^{2}}.$ (44)
-Solving the system of equations of the projected cap and the deprojected velocity ring gives intersection points of ($p$, $q$) such that
-$p=\pm\sqrt{r_{\rm b}^{2}-q^{2}}$ (45)
+The azimuthal “ring” of gas with velocity $w$ has a radius $r_{\rm w}$ defined by the sine of the angle $\varrho$ between the observed and deprojected velocities $u$ and $w$, respectively.
+Relativistic aberration (i.e., beaming effect) is invariant across the azimuthal angle, ergo a classical treatment in the source frame is still accurate for deriving the geometry of the projected cone.
+The apparent deprojection radius $r_{\rm w}$ for each azimuthal ring is thus
+$r_{\rm w}=\sqrt{1-\left(\frac{u}{w}\right)^{2}},$ (48)
+which we show in Figure 22. Solving the system of equations of the projected cap and the deprojected velocity ring gives intersection points of ($p$, $q$) such that
+$p=\pm\sqrt{r_{\rm w}^{2}-q^{2}}$ (49)
 and, from substituting $p$ and employing the quadratic equation,
@@
 $B$ $=-2S\left(\frac{g}{h}\right)^{2}$ (52)
-$C$ $=r_{\rm b}^{2}+\left(\frac{gS}{h}\right)^{2}-g^{2}.$ (49)
+$C$ $=r_{\rm w}^{2}+\left(\frac{gS}{h}\right)^{2}-g^{2}.$ (53)
 To reduce computation time, $A$ and $B$ are calculated outside of the integral as only $C$ depends on the de-projected velocity $w$.
 The intersection points ($p$, $q$) given by $A$, $B$, and $C$ yield the arc of the enclosed deprojected velocity such that
-$\pi\ell_{\rm w}=\begin{cases}\arccos{\frac{q}{r_{\rm b}}}&q>0\\ \pi-\arccos{\frac{q}{r_{\rm b}}}&q<0\end{cases}$ (50)
+$\pi\ell_{\rm w}=\begin{cases}\arccos{\frac{q}{r_{\rm w}}}&q>0\\ \pi-\arccos{\frac{q}{r_{\rm w}}}&q<0\end{cases}$ (54)
 which notably avoids the additional step of calculating $p$.
+We illustrate such a solution for $\pi\ell$, including the velocity dependence, in Figure 22.
 If multiple real solutions for ($p$, $q$) exist due to the lune from the projection of the second cone into the plane, then OutLines simply sums the two arcs.
@@
 Calling OutLines.PlotConeProjection(i,thetaO) will draw the geometry and projection as shown in Figure 23 for any inclination and opening angle.
-Figure 21: Directed outflows (left of each panel) projected onto the plane of the sky (right of each panel) for various combinations of $i$ and $\theta_{o}$ (orange).
+Figure 23: Following Figure 14, we illustrate bidirectional cone geometries (left of each panel) projected onto the plane of the sky (right of each panel) for various combinations of $i$ and $\theta_{o}$ (orange).
 Circles indicate deprojected velocities of u/w=u/w=0.4 (green), 0.6 (blue), and 0.8 (purple).
@@
 Calling OutLines.PlotConeProjection(i,thetaO,thetaC) will illustrate projections similar to those of the filled cones but for the hollow cone geometry.
-Figure 22: Same as Figure 21 but for a directed outflow with an inner cavity.
+Figure 24: Same as Figure 23 but for directed cones with an inner cavity.
 A.4 Obstruction by a Disk
-In §2.4, we present the possible inclusion of a disk which encircles the source of the outflow and is inclined orthogonally to the cone axis.
-This disk obstructs the posterior cone but may affect either the red-shifted or both wings of the line profile.
+In §2.4, we present the possible inclusion of a disk which encircles the source of the outflow and is inclined orthogonally to a cone or hemisphere axis.
+This disk obstructs the posterior cone or hemisphere but may affect either the red-shifted or both wings of the line profile.
 OutLines projects the disk onto the line of sight and then determines the length of each deprojected velocity which is still observed after accounting for the disk effect.
-Since the disk is centered at the source of the outflow and assumed to be orthogonal to the outflow cone, the projected disk is simply an ellipse with semi-major and semi-minor axes of ($w[x_{d}|\beta],w[x_{d}|\beta]\sin i$).
+Since the disk is centered at the source of the outflow and assumed to be orthogonal to the outflow cone, the projected disk is simply an ellipse with semi-major and semi-minor axes of ($w[x_{d}|\beta],w[x_{d}|\beta]\cos i$).
 Thus, the intersection points ($p_{d}$$q_{d}$) between the deprojected velocity and the disk can be written as
-$p_{d}=\pm\sqrt{r_{\rm b}^{2}-q_{d}^{2}}$ (51)
+$p_{d}=\pm\sqrt{r_{\rm w}^{2}-q_{d}^{2}}$ (55)
 and
-$q_{d}=\pm\sqrt{\frac{r_{\rm b}^{2}-w^{2}[x_{d}|\beta]}{\csc^{2}i-1}}.$ (52)
+$q_{d}=\pm\sqrt{\frac{r_{\rm w}^{2}-w^{2}[x_{d}|\beta]}{\sec^{2}i-1}}.$ (56)
 Rather than compute $\ell_{\rm w}$ and $\ell_{\rm disk}$ separately, which would needlessly incur a computational overhead, we can instead simply compare $q$ and $q_{d}$ to determine the final $\ell_{\rm w}$.
-For the posterior cone, $\ell_{\rm w}$ is given by
-$\pi\ell_{\rm w}=\begin{cases}\arccos{\frac{q}{r_{\rm b}}}&q>q_{d}\\ \arccos{\frac{q_{d}}{r_{\rm b}}}&-q_{d}<q<q_{d}\\ 2\arccos{\frac{q_{d}}{r_{\rm b}}}-\arccos{\frac{q}{r_{\rm b}}}&q<-q_{d}\\ \end{cases}.$ (53)
+For the posterior cone or hemisphere, $\ell_{\rm w}$ is given by
+$\pi\ell_{\rm w}=\begin{cases}\arccos{\frac{q}{r_{\rm w}}}&q>q_{d}\\ \arccos{\frac{q_{d}}{r_{\rm w}}}&-q_{d}<q<q_{d}\\ 2\arccos{\frac{q_{d}}{r_{\rm w}}}-\arccos{\frac{q}{r_{\rm w}}}&q<-q_{d}\\ \end{cases}.$ (57)
 If $q$ has multiple solutions due to the projected lune from the anterior cone, OutLines will add to $\ell_{\rm w}$ the first case evaluated using the second solution for $q$ with the imposition that $q>q_{d}$ since the anterior cone is not obstructed by the disk.
@@
 Under this formalism, $\ell_{\rm cavity}$ can still be subtracted to determine the hollow cone geometry as both $\ell_{\rm w}$ and $\ell_{\rm cavity}$ will account for any disk effects on the deprojected velocity apriori.
-Appendix B Non-zero Launch Velocities
-While the OutLines default assumes that gas is launched from rest, we have allowed for the possibility that the gas starts from some launch velocity $v_{0}$, which we include as an optional free parameter.
-The velocity limits are accordingly updated so that the minimum velocity $w_{\ell}$ which contributes to an observed velocity $|u|$ is never below $v_{0}/v_{\infty}$.
-The velocity fields in Table 4 and Equations 15-17 are modified by a factor of $1-v_{0}/v_{\infty}$ with the addition of a constant $v_{0}/v_{\infty}$ such that the relative velocity not from rest $w_{\rm nfr}$ is
-$w_{\rm nfr}=w[x|\beta]\left(1-\frac{v_{0}}{v_{\infty}}\right)+\frac{v_{0}}{v_{\infty}}.$ (55)
-Emergent profiles are similar to those presented in TenorioTagle1996, which included distributions of non-zero shell ages, and to those presented in 2015ApJ...801...43S, which account for non-zero launch velocities.
-We show example line profiles with varying launch velocities in Figure 23.
-Future work will revisit this possibility in greater detail.
-Figure 23: Effect of varying the launch velocity from rest ($v_{0}=0$) to 30% of the terminal velocity on nebular emission (left) and absorption (right) line profiles, here normalized to the case of launching from rest for visualization.
+Appendix B Implementation of OutLines
 Table 2: Left: the Profile object classes in OutLines with which users can generate model profiles for different spectral line types.
 Right: keyword settings for the Profile classes to customize models.
+Disk not implemented for spherical geometries.
+Pulse only applicable ensemble density profiles.
 class use atomic data keywords options default
@@
 Aperture
+Pulse
 Table 3
@@
 boolean
+Table 5
+Bubbles/Shells
 Spherical
@@
 False
+‘Normal’
 Resonant resonant emission lines $\lambda_{0},\ f,\ p_{r}$
@@
 Absorption resonant absorption lines $\lambda_{0},\ f$
-Appendix C Implementation of OutLines
-C.1 Installation
+B.1 Installation
 Installing OutLines can be done in the terminal command line using pip as shown below.
@@
 Other dependencies, including numpy, scipy, and matplotlib will be automatically managed by pip, including within conda environments.
-C.2 The Profile Classes
+B.2 The Profile Classes
 The primary front-end user interaction with OutLines is through the various Profile object classes, which we list in Table 2.
@@
 The terms and related keyword arguments for the profile model are listed in Table 3 for the geometry, Table 4 for the velocity field, and Table 5 for the density profile, with the OutLines defaults indicated by a .
-Table 3: Geometry keyword options and related required arguments for OutLines. indicates the OutLines default. indicates geometry only applicable to emission lines.
+In some cases, users may want to consider a velocity field or density profile not included in OutLines.
+In such cases, if a user passes a function to either keyword instead of a string, OutLines will implement the user-defined field or profile.
+Table 3: Geometry keyword options and related required arguments for OutLines. indicates the OutLines default.
 Geometry terms assumptions
 ‘Spherical’  none $i=0$, $\theta_{c}=0^{\circ}$, $\theta_{o}=90^{\circ}$
+‘Hemisphere’ $i$ $\theta_{c}=0^{\circ}$, $\theta_{o}=90^{\circ}$
 ‘FilledCones’ $i$, $\theta_{o}$ $\theta_{c}=0^{\circ}$
@@
 VelocityField equation references
-‘BetaCAK’  15 1975ApJ...195..157C; Castor1979
-‘AccPlaw’ 16 2010ApJ...717..289S
-‘VelPlaw’ 17 FaucherGiguere2012
+‘BetaCAK’  15 Castor et al. (1975a); Castor and Lamers (1979)
+‘AccPlaw’ 16 Steidel et al. (2010)
+‘VelPlaw’ 17 Faucher-Giguère and Quataert (2012)
 Table 5: Density profile keyword options and related required arguments for OutLines. indicates the OutLines default.
 DensityProfile terms equation
-‘PowerLaw’  $\alpha$ 18
-‘Exponential’ $\gamma$ 19
-‘Normal’ $x_{1}$, $\sigma_{x}$ 21
-‘LogNormal’ $\log x_{1}$, $\log\sigma_{x}$ 22
-‘Shell’ $x_{1}$, $\sigma_{x}$ 24
-‘FRED’ $x_{1}$, $\tau_{1}$, $\tau_{2}$ 23
-‘DoublePowerLaw’ $\alpha_{1}$, $\alpha_{2}$, $x_{1}$ 20
-‘Pulses’ $x_{1}$, $\sigma_{x}$, $x_{k}$ 25
-‘DampedPulses’ $\gamma$, $x_{1}$, $\sigma_{x}$, $x_{k}$ 26
+Winds
+‘PowerLaw’  $\alpha$ 19
+‘Exponential’ $\gamma$ 20
+‘DoublePowerLaw’ $\alpha_{1}$, $\alpha_{2}$, $x_{1}$ 21
+Bubbles / Shells
+‘Normal’ $x_{1}$, $\sigma_{x}$ 22
+‘LogNormal’ $\log x_{1}$, $\log\sigma_{x}$ 23
+‘DLogic‘ $\log x_{1}$, $k$ 25
+‘Shell’ $x_{1}$, $\sigma_{x}$ 26
+‘FRED’ $x_{1}$, $\tau_{1}$, $\tau_{2}$ 24
+Ensembles
+‘Pulses’ $x_{1}$, $\sigma_{x}$, $x_{k}$ 27
+‘DampedPulses’ $\gamma$, $x_{1}$, $\sigma_{x}$, $x_{k}$ 28
 Users will primarily use OutLines by instantiating one of the several Profile classes: Nebular, Resonant, Fluorescent, and Absorption.
 The object-bound methods of the Profile classes can be used to obtain or update parameter values and to generate line profiles.
-We provide examples of using these classes to generate line profiles in Appendix D.
+We provide examples of using these classes to generate line profiles in Appendix C.
 To obtain line profiles computed by OutLines, users can call the get_profile() bound method for either class, which requires only an array of wavelengths and computes a line profile for the parameters stored in the object.
@@
 All Profile classes can each support multiple lines, something we encourage users to consider.
-C.3 The Properties Class
+B.3 The Properties Class
 After modeling a wind or outflow from observed spectral lines, users will want to derive important diagnostic properties to determine what drives the wind, how, and to what extent the outflow impacts its environment.
@@
 x.out $x_{out}$ characteristic radius
-v.out 32 $v_{out}$ characteristic radius
-Rcal 33 $\mathcal{R}$ relative column density
-Mdot 37 $\dot{M}/n_{0}R_{0}^{2}$ mass outflow rate
-pdot 38 $\dot{p}/n_{0}R_{0}^{2}$ momentum flux
-Edot 39 $\dot{E}/n_{0}R_{0}^{2}$ mechanical luminosity
-C.4 Visualization
+v.out 36 $v_{out}$ characteristic radius
+Rcal 37 $\mathcal{R}$ relative column density
+Mdot 41 $\dot{M}/n_{0}R_{0}^{2}$ mass outflow rate
+pdot 42 $\dot{p}/n_{0}R_{0}^{2}$ momentum flux
+Edot 43 $\dot{E}/n_{0}R_{0}^{2}$ mechanical luminosity
+B.4 Visualization
 To assist with interpretation of outflow properties and geometry, OutLines contains several functions to generate figures like the ones presented here.
@@
 To produce the edge-on geometric cartoon as in Figure 10, we provide the PlotGeometry function.
-To produce the velocity projections for filled and hollow cones as in Figures 21 and 22, we provide the PlotConeProjection function.
-C.5 Numerical Methods
-Some equations require root-finding where analytic solutions are not possible, including the “source-effect” and aperture velocity limits (Equations 28 and 30, respectively) and the characteristic outflow radius $x_{out}$ where the momentum density is maximized.
-As in Flury2023, we employ Brent’s method implemented by scipy.optimize to solve for the root in each case.
-The initial profile model of Flury2023 solved the integral for the emission line profile (Equation 10) using the adaptive Romberg method, which included Richardson extrapolation under the Bulirsch-Stoer method.
+To produce the cone and velocity projection geometry for filled cones as in Figure 22, we provide the PlotProjGeom function.
+To produce the velocity projections for filled and hollow cones as in Figures 23 and 24, we provide the PlotConeProj function.
+B.5 Numerical Methods
+Some equations require root-finding where analytic solutions are not possible, including the “source-effect” and aperture velocity limits (Equations 32 and 34, respectively) and the characteristic outflow radius $x_{out}$ where the momentum density is maximized.
+As in Flury et al.
+(2023), we employ Brent’s method implemented by scipy.optimize to solve for the root in each case.
+The initial profile model of Flury et al.
+(2023) solved the integral for the emission line profile (Equation 10) using the adaptive Romberg method, which included Richardson extrapolation under the Bulirsch-Stoer method.
 As implemented by the now-deprecated scipy.integrate.romberg, this method can be costly and in later scipy versions has loss of accuracy in non-spherical geometries due to poor sampling of the integrand.
-For improvements to both speed and accuracy, we now adopt Gaussian quadrature with a fixed number of nodes (as implemented by scipy.integrate.fixed_quad) to solve each of the integrals for $\phi_{\lambda}$.
-We find at least 16 nodes necessary to obtain reliable profile solutions in many cases but conservatively assume 256 nodes for optimal accuracy across the profile parameter spaces.
-For a single core on a conventional laptop computer, benchmarking indicates that OutLines can calculate an emission line profile at $R=5,000$ in $\approx$20 ms for any assumed geometry.
-Given reasonable guess parameters, iterative non-linear least squares fitting routines like the trust-reflective algorithm (see scipy.optimize.least_squares) perform rapidly, typically converging on a solution within 10 to 30 s depending on the data and the assumed geometry.
-For 25 walkers, a typical MCMC sampler (here benchmarked using emcee, emcee) takes anywhere from $0.1$ to $2$ s per step to run on a conventional laptop computer when fitting an OutLines model to an observed line profile, even when accounting for complex geometries and LSF effects.
+For improvements to both speed and accuracy, we now adopt Gauss-Legendre quadrature with a fixed number of nodes to solve the integrals for $\phi_{\lambda}$.
+We implement quadrature integration following the scipy.integrate.fixed_quad; however, for speed, we have pre-computed the nodes and weights using scipy.special.roots_legendre, which takes $\approx 1\rm\>ms$ for each call to scipy.integrate.fixed_quad.
+We find at least 16 nodes necessary to obtain reliable profile solutions in many cases but conservatively assume 96 nodes for optimal accuracy across the profile parameter spaces without significant computational expense.
+For bubbles, we restrict velocity bounds where possible to improve sampling of the density profile.
+Future work may include adopting Gauss-Kronrod integration for improved accuracy and precision, e.g., following Laurie (1997).
+To normalize the line profile, we use a linear integration method to effectively handle cusps which may occur at peaks and troughs, typically near the line core.
+We implement this integration following the shoelace algorithm case of Green’s theorem.
+Benchmarking against the scipy.integrate.trapezoid implementation of trapezoid rule integration, we find comparable results with a marginally faster (by $1\rm\>ms$) compute time.
+For a single core on a conventional laptop computer, benchmarking indicates that OutLines can calculate an emission line profile at $R=5,000$ in $6\rm\>ms$ for the most complex geometry (disk plus hollow cones), and that the compute time increases almost linearly with $R$ such that run times are $2.5\rm\>ms$ at $R=2,000$, $4\rm\>ms$ at $R=3,000$, $13\rm\>ms$ at $R=10,000$, $25\rm\>ms$ at $R=20,000$, and $65\rm\>ms$ at $R=50,000$.
+For the simplest geometry (spherical), the compute times are much faster ($1\rm\>ms$ for $R\leq 30,000$).
+Given reasonable guess parameters, iterative non-linear least squares fitting routines like the trust-reflective algorithm (see scipy.optimize.least_squares) perform rapidly, typically converging on a solution within 10 to 30 s depending on the data and the assumed model.
+For 25 walkers, a typical MCMC sampler (here benchmarked using emcee, Foreman-Mackey et al.
+2013) takes anywhere from $0.1$ to $2$ s per step to run on a conventional laptop computer when fitting an OutLines model to an observed line profile, even when accounting for complex geometries and LSF effects.
 The addition of convolution with a line spread function does increase the computation time; however, the extent to which this occurs depends on the resolution, size of the data set, and the implementation of the convolution.
-Appendix D Example Line Profiles with OutLines
+Appendix C Example Line Profiles with OutLines
 Below, we provide example usages of the Absorption, Nebular, and Fluorescent line profile object classes in OutLines.
@@
 We also demonstrate how to compute outflow properties from the Properties class and print the results to the command line.
-D.1 C ii $\lambda 1334$
-Example instantiations of Absorption, Resonant, and Fluorescent objects from the Profile class using atomic data from Morton2003.
+C.1 C ii $\lambda 1334$
+Example instantiations of Absorption, Resonant, and Fluorescent objects from the Profile class using atomic data from Morton (2003).
 Here, the object classes are instantiated, documentation, setting, and parameters are printed to the terminal, and the column density and terminal velocity are updated.
@@
 21 model.update_params('TerminalVelocity', 500.00)
-D.2 [O iii] $\lambda 4959,5007$
-Example insantiation of the Nebular object from the Profile class in OutLines using atomic data from NIST (NIST_ASD) and Storey2000.
+C.2 [O iii] $\lambda 4959,5007$
+Example insantiation of the Nebular object from the Profile class in OutLines using atomic data from NIST (Kramida et al., 2024) and Storey and Zeippen (2000).
 Here, we illustrate how to customize the model with keyword arguments passed to the object class on instantiation, updating parameters, printing setting and parameters, computing and printing outflow properties, and visualizing a line profile as in Figure 25.
@@
 25plt.show()
-Figure 24: Example [O iii] $\lambda\lambda 4959,5007$ line doublet profile produced using OutLines as shown in Appendix D.2.
-Appendix E Example Modeling
+Figure 25: Example [O iii] $\lambda\lambda 4959,5007$ line doublet profile produced using OutLines as shown in Appendix C.2.
+Appendix D Example Modeling
 Below, we provide examples for modeling absorption and nebular emission line features using line profiles computed by OutLines.
 We indicate suggested methods for obtaining best-guess parameters and demonstrate how to use OutLines in tandem with MCMC samplers using the initial value generator and logarithmic posterior probability contained in the object instance of the profile class.
-E.1 Si ii $\lambda 1190,1193$ Absorption Doublet
+D.1 Si ii $\lambda 1190,1193$ Absorption Doublet
 1# import the OutLines package
@@
 21model.print_params()
-E.2 [S ii] $\lambda 6716,6731$ Emission Doublet
+D.2 [S ii] $\lambda 6716,6731$ Emission Doublet
 1# import the OutLines package
@@
 31props.print_props()
-E.3 [O iii] $\lambda 4959,5007$ Emission Doublet for the Seyfert 2 in NGC 2992
+D.3 [O iii] $\lambda 4959,5007$ Emission Doublet for the Seyfert 2 in NGC 2992
 1# import the OutLines package
@@
 6from scipy.optimize import curve_fit
-7# MC sampler for posterior
-8import emcee
-9# log prior -- uniform bounded
-10def log_prior(self,theta):
+7# scipy convolution function for speed
+8from scipy.signal import convolve
+9# MC sampler for posterior
+10import emcee
+11# log prior -- uniform bounded
+12def log_prior(self,theta):
 13 tlower,tupper = self.get_bounds()
@@
 16 # suggest atomic data to within 5%
-15 # suggest that the best-guess angles
-16 # are accurate within 10 degrees
 17 if all(tbound) and theta[8]-theta[9] > 0.087:
-18 return -0.5*((2.99e5*theta[1]-1100)/300)**2 \
-19 -0.5*((2.99e5*theta[0]- 75)/50)**2 \
-20 -0.5*((theta[7]-1.25)/0.2)**2 \
-21 -0.5*((theta[8]-0.95)/0.2)**2 \
-22 -0.5*((theta[9]-0.40)/0.2)**2 \
-23 -0.5*((theta[4]/theta[3]-2.98)/0.05)**2 \
+18 return -0.5*((theta[4]/theta[3]-2.98)/0.05)**2 \
 19 -0.5*((theta[6]/theta[5]-2.98)/0.05)**2
@@
 21 return -inf
-27# instantiate the nebular line
-28# profile object using
-29# rest wavelengths from NIST
-30kwargs = dict(Geometry='HollowCones',AddStatic=True,\ VelocityField='AccPlaw',Disk=True)
-31model = OL.Nebular([4958.911,5006.843],**kwargs)
-32# update some parameters
-33model.update_params(['TerminalVelocity'],[500,30,15,45])
-34model.update_params(['OpeningAngle','CavityAngle',\ 'Inclination'],[60,30,60])
-35model.update_params(['FluxOutflow1','FluxStatic1'],\ [1/2.98,1/2.98]) # Storey & Zeppen 2000
-36# simple trust-reflective fit for initial guess
-37tr_kwargs = dict(p0=model.get_params(),\ bounds=model.get_bounds(),sigma=flux_error)
-38p0,c0 = curve_fit(model.Profile,wave,flux,**tr_kwargs)
-39model.update_params(model.get_param_names(),p0)
-40# initialize walkers and sampler and run the sampler
-41init,nwalk = model.init_params()
-42# use the custom prior
-43model.log_prior = MethodType(log_prior,model)
-44# run the MCMC sampler
-45sampler = emcee.EnsembleSampler(nwalk,model.npar,\ model.log_probability,args=(wave,flux,flux_error))
-46sampler.run_mcmc(init,1200, progress=True)
-47samples = sampler.get_chain(discard=200,flat=True)
-48# extract parameters from posterior sample
-49for i,param in enumerate(model.get_param_names()):
+22# log likelihood for chi squared
+23# including convolution of line profile with LSF
+24# assuming the LSF kernel `kern` is defined
+25def log_likelihood(self,theta,x,y,yerr):
+26 ybar = convolve(model.Profile(x,*theta)+cont[ind],kern,'same')
+27 ln_prb = -((y-ybar)/yerr)**2 - 2*log(yerr)
+28 return nansum(ln_prb)
+29# instantiate the nebular line
+30# profile object using
+31# rest wavelengths from NIST
+32kwargs = dict(Geometry='HollowCones',AddStatic=True,\ VelocityField='AccPlaw',Disk=True)
+33model = OL.Nebular([4958.911,5006.843],**kwargs)
+34# update some parameters
+35model.update_params(['TerminalVelocity'],[500,30,15,45])
+36model.update_params(['OpeningAngle','CavityAngle',\ 'Inclination'],[60,30,60])
+37model.update_params(['FluxOutflow1','FluxStatic1'],\ [1/2.98,1/2.98]) # Storey & Zeppen 2000
+38# simple trust-reflective fit for initial guess
+39tr_kwargs = dict(p0=model.get_params(),\ bounds=model.get_bounds(),sigma=flux_error)
+40p0,c0 = curve_fit(model.Profile,wave,flux,**tr_kwargs)
+41model.update_params(model.get_param_names(),p0)
+42# initialize walkers and sampler and run the sampler
+43init,nwalk = model.init_params()
+44# use the custom prior and likelihood
+45model.log_prior = MethodType(log_prior,model)
+46model.log_likelihood = MethodType(log_likelihood,model)
+47# run the MCMC sampler
+48sampler = emcee.EnsembleSampler(nwalk,model.npar,\ model.log_probability,args=(wave,flux,flux_error))
+49sampler.run_mcmc(init,1200, progress=True)
+50samples = sampler.get_chain(discard=200,flat=True)
+51# extract parameters from posterior sample
+52for i,param in enumerate(model.get_param_names()):
 53 value = nanquantile(samples[:,i],0.5)
 54 model.update_params(param,value)
-52# print line profile parameters to terminal
-53model.print_params()
-54# obtain outflow properties
-55props = OL.Properties(model)
-56# print outflow properties to terminal
-57props.print_props()
-Appendix F Posterior Sampling
-All posterior samples for each of the OutLines modeling cases were obtained using emcee (emcee) and visualized with figures produced by corner (corner) and subsequently adapted.
-Results are shown for NGC 5471 Knot A in Figure 25, the super star cluster of Green Pea J1044+0353 in Figure 26, Mrk 1486 in 27, and the Seyfert 2 nucleus of NGC 2992 in 28.
+55# print line profile parameters to terminal
+56model.print_params()
+57# obtain outflow properties
+58props = OL.Properties(model)
+59# print outflow properties to terminal
+60props.print_props()
+Appendix E Posterior Sampling
+All posterior samples for each of the OutLines modeling cases were obtained using emcee (Foreman-Mackey et al., 2013) and visualized with figures produced by corner (Foreman-Mackey, 2016) and subsequently adapted.
+Results are shown for NGC 5471 Knot A in Figure 26, the super star cluster of Green Pea J1044+0353 in Figure 27, Mrk 1486 in 28, and the Seyfert 2 nucleus of NGC 2992 in 29.
 […]